### v2

# From graph to Riesz continuity

Abstract
We show that every graph continuous family of unbounded operators in a Hilbert space becomes Riesz continuous after multiplication by an appropriate family of unitary operators. This result provides a simple definition of the index for graph continuous families of Fredholm operators and leads to two corollaries for operators with compact resolvents: (1) the identity map between the spaces of such operators equipped with the Riesz and the graph topology is a homotopy equivalence; (2) every graph continuous family of such operators acting between fibers of Hilbert bundles becomes Riesz continuous in appropriate trivializations of the bundles.
For self-adjoint operators, multiplication by unitaries should be replaced by conjugation. In general, a graph continuous family of self-adjoint operators with compact resolvents cannot be made Riesz continuous by an appropriate conjugation. We obtain a partial analogue of the “trivialization” result above for self-adjoint operators and describe obstructions to existence of such a trivialization arising in the general case. This part of results is closely related to the recent work [$I_{2}$] of N. Ivanov.
footnotetext: Department of Mathematics, Technion – Israel Institute of Technology.footnotetext: This work was partially supported by ISF grants no. 431/20 and 844/19.
1 Introduction
Let $H$ and $H^{\prime}$ be separable complex Hilbert spaces of infinite dimension. We denote by $\mathcal{B}(H,H^{\prime})$ the space of bounded linear operators $H\to H^{\prime}$ with the norm topology; by $\mathcal{K}(H,H^{\prime})$ the subspace of $\mathcal{B}(H,H^{\prime})$ consisting of compact operators, by $\mathcal{U}(H)$ and $\mathcal{P}(H)$ the subspaces of $\mathcal{B}(H)=\mathcal{B}(H,H)$ consisting of unitary operators and projections respectively (by projections we always mean orthogonal projections).
Regular operators.   An unbounded operator $A$ from $H$ to $H^{\prime}$ is a linear operator defined on a subspace $\operatorname{dom}(A)$ of $H$ and taking values in $H^{\prime}$. Such an operator $A$ is called closed if its graph is closed in $H\oplus H^{\prime}$, and densely defined if its domain $\operatorname{dom}(A)$ is dense in $H$. It is called regular if it is closed and densely defined. Let $\mathcal{R}(H,H^{\prime})$ denote the set of all regular operators $H\to H^{\prime}$ and $\mathcal{R}^{\mathrm{sa}}(H)\subset\mathcal{R}(H)=\mathcal{R}(H,H)$ denote the subset of self-adjoint operators.
Two topologies on regular operators.   The set $\mathcal{R}(H,H^{\prime})$ of regular operators has two important natural topologies. The Riesz topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
$\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow\mathcal{B}(H,H^{\prime}),\quad\bm{\chi}(A)=A(1+A^{*}A)^{-1/2}$
from the norm topology on the space of bounded operators. Here $\bm{\chi}$ is the so called bounded transform map. The graph topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
$\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow\mathcal{P}(H\oplus H^{\prime})$
from the norm topology on the space of projections, where $\mathbf{p}$ is the map taking a regular operator to the orthogonal projection onto its graph. Let ${{}^{r}\mathcal{R}}(H,H^{\prime})$, resp. ${{}^{g}\mathcal{R}}(H,H^{\prime})$ denote the space of regular operators $H\to H^{\prime}$ equipped with the Riesz, resp. graph topology.
The Riesz topology is strictly stronger than the graph topology. In other words, the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is continuous, while the identity map ${{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ is not. On the subset $\mathcal{B}(H,H^{\prime})$ of bounded operators these two topologies coincide.
From graph to Riesz continuity.   Riesz continuous families of operators are much easier to deal with than mere graph continuous families. This motivates the approach that we are developing in this paper.
Suppose that one is given a graph continuous family of regular operators. One can try to modify it, in an appropriate way, so that the resulting family becomes Riesz continuous, then investigate the modified family and transfer results of this investigation to the original family. By an “appropriate way” we mean that such a transfer of results is possible, that is, some important properties of the original family may be recovered from the corresponding properties of the modified family. Of course, this description is quite vague; we will make it more precise below.
Let us illustrate our approach by an example of a family consisting of invertible operators. Recall that an operator $A\in\mathcal{R}(H)$ is called invertible if it is bijective as a map $\operatorname{dom}(A)\to H$. Such operator has a unique polar decomposition $A=u\cdot|A|$, where $u\in\mathcal{U}(H)$ is a unitary operator and $|A|=\sqrt{A^{*}A}$ is a positive operator. If $\mathcal{A}\colon X\to\mathcal{R}(H)$ is graph continuous, then the map $x\mapsto|\mathcal{A}_{x}|$ is Riesz continuous. If additionally all operators $\mathcal{A}_{x}$ are invertible, then the map $x\mapsto u_{x}=\mathcal{A}_{x}|\mathcal{A}_{x}|^{-1}$ from $X$ to ${{}^{s}{\mathcal{U}}}(H)$, the unitary group equipped with the strong operator topology, is continuous. Taking $v_{x}=u_{x}^{*}$, we obtain a continuous map $v\colon X\to{{}^{s}{\mathcal{U}}}(H)$ such that the map $x\mapsto v_{x}\mathcal{A}_{x}=|\mathcal{A}_{x}|$ is Riesz continuous. The original family $\mathcal{A}$ can be recovered from the modified family $|\mathcal{A}|$ by the left multiplication by $u_{x}$.
This construction cannot be extended continuously to general non-invertible families because the polar decomposition of a general non-invertible operator involves a partial isometry which is not unitary. Moreover, this partial isometry does not depend continuously on the operator. However, one can still modify a family to a Riesz continuous one using the multiplication by an appropriate family of unitary operators, as we show in Theorem A below (we will multiply on the right, but it really does not matter: one can always pass to adjoint operators, replacing the left multiplication with the right one). In contrast with the construction above, the modified family no longer consists of positive operators, nor is it determined by the original family in any canonical way. But the modified family still carries an important information about the original one.
Theorem A.
There exists a map $v\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H)$, which is both Riesz-to-norm and graph-to-strong continuous, such that the induced map
$\Phi\colon\mathcal{R}(H,H^{\prime})\to\mathcal{R}(H,H^{\prime}),\quad A\mapsto A\cdot v(A)$
is graph-to-Riesz continuous.
[…]
$K^{0}$ index for graph continuous families.   Let $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}$ be a graph continuous family and $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem A. Then $\bm{\chi}\circ\Phi\circ\mathcal{A}=\varphi\circ\mathcal{A}$ is a norm continuous family of bounded operators. Suppose that all $\mathcal{A}_{x}$ are Fredholm. Then the operators $\varphi(\mathcal{A}_{x})$ are also Fredholm, and one can define the index of $\mathcal{A}$ as the index of this norm continuous family of bounded Fredholm operators:
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\varphi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X).$
This definition of the index is discussed in the last part of Section 3; it is justified by Theorem B below.
Equivalently, the index of $\mathcal{A}$ may be defined as the index of the family of projections $\Gamma_{x}\to H^{\prime}$ after a trivialization of the Hilbert bundle $\Gamma$ formed by the graphs $\Gamma_{x}$ of operators $\mathcal{A}_{x}$. (This trivialization should be chosen in the category of bundles with the structure group $\mathcal{U}$, not ${{}^{s}{\mathcal{U}}}$.) Note that such a trivialization is not canonical; it is defined only up to a continuous map $X\to\mathcal{U}(H)$. However, change of a trivialization does not affect the homotopy type of the corresponding family of bounded Fredholm operators, since the unitary group $\mathcal{U}(H)$ is contractible.
Searching for such a definition of the index was part of the motivation behind Theorem A. An alternative approach to the index of graph continuous families of Fredholm operators is provided by the previous paper [$P_{3}$] of the author. Another approach, for compact base spaces and in a more general framework of Hilbert modules over C*-algebras, is provided by results of M. Joachim [J]. Recently N. Ivanov [$I_{1}$, $I_{2}$] developed a new approach to the index of families, which works under much weaker continuity assumptions than graph continuity. We show in the end of Section 3 that for graph continuous families of Fredholm operators our definition of the index given above coincides with Ivanov’s index.
Homotopy equivalences.   A regular operator $A$ is said to have compact resolvents***If the resolvent set $\operatorname{Res}(A)$ of $A$ is non-empty (for example, if $A$ is self-adjoint), then this definition agrees with the usual sense of the words “compact resolvent”: $(A-\lambda)^{-1}$ is compact for some (and then every) $\lambda\in\operatorname{Res}(A)$. However, a regular operator may have an empty resolvent set; the definition above covers such operators as well. if both $(1+A^{*}A)^{-1}$ and $(1+AA^{*})^{-1}$ are compact operators. Let $\mathcal{R}_{F}(H)$ and $\mathcal{R}_{K}(H)$ be the subsets of $\mathcal{R}(H)$ consisting of Fredholm operators and operators with compact resolvents, respectively. These subsets are invariant under the multiplication by unitary operators, so $\Phi(\mathcal{R}_{K})\subset\mathcal{R}_{K}$ and $\Phi(\mathcal{R}_{F})\subset\mathcal{R}_{F}$.
It was shown in the previous paper of the author that the identity maps
${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}\;\;\text{and}\;\;{{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$
are homotopy equivalences. See [$P_{3}$, Theorem A]. However, the proof in [$P_{3}$] provides no intuition on how a homotopy inverse map might look like. The following theorem reveals a possible form of homotopy inverse maps and provides an alternative proof of homotopy equivalence of the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$.
Theorem B.
Let $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem A. Then the restriction of $\Phi$ to the subspace of operators with compact resolvents is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$. Similarly, the restriction of $\Phi$ to the subspace of Fredholm operators is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$.
Let $J$ be a subset of $\mathcal{K}(H)$ invariant under the conjugation by unitary operators. Then one can replace $\mathcal{R}_{K}$ in the statement of Theorem B with the set
$\mathcal{R}_{J}(H)=\left\{A\in\mathcal{R}(H)\,|\,(1+A^{*}A)^{-1},(1+AA^{*})^{-1}\in J\right\}.$
(In particular, $J$ can be the subset of trace class operators, or the subset of Hilbert–Schmidt operators.) Such a set $\mathcal{R}_{J}$ is invariant under the multiplication by unitary operators, so both $\Phi$ and homotopies used in the proof of Theorem B preserve $\mathcal{R}_{J}$. This provides the following generalization of the first part of Theorem B.
Theorem B’. Let $J$ be a subset of $\mathcal{K}(H)$ invariant under the conjugation by unitary operators. Then the identity map ${{}^{r}\mathcal{R}}_{J}(H)\to{{}^{g}\mathcal{R}}_{J}(H)$ is a homotopy equivalence, with a homotopy inverse map given by $\Phi\colon{{}^{g}\mathcal{R}}_{J}\to{{}^{r}\mathcal{R}}_{J}$.
Hilbert bundles.   Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be locally trivial Hilbert bundles over a topological space $X$ with a fiber $H$. The structure group of a general Hilbert bundle is the unitary group ${{}^{s}{\mathcal{U}}}(H)$ equipped with the strong operator topology. As was shown by Dixmier and Douady, every such bundle over a paracompact base space $X$ is trivial, that is, isomorphic to the trivial Hilbert bundle $X\times H\to X$. See [DD], Theorem 1. More generally, every numerable Hilbert bundle is trivial. However, automorphisms of the trivial Hilbert bundle $X\times H\to X$ are given by continuous maps $X\to{{}^{s}{\mathcal{U}}}(H)$ and preserve neither norm nor Riesz continuity of operator families, since the action of ${{}^{s}{\mathcal{U}}}(H)$ on $\mathcal{B}(H)$ by conjugation, $(u,a)\mapsto uau^{-1}$, is not continuous. Therefore, the notions of a norm continuous family of bounded operators or of a Riesz or graph continuous family of regular operators are not defined for a general Hilbert bundle.
Nevertheless, the actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on the space $\mathcal{K}(H,H^{\prime})$ of compact operators are continuous, and thus the notion of a norm continuous family of compact operators $k_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ is well defined. By the same reason, the notion of a graph continuous family of regular operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ with compact resolvents is well defined. Applying Theorem A to such a family, we obtain the following result.
Theorem C.
Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be numerable Hilbert bundles over a topological space $X$. Let $\mathcal{A}$ be a graph continuous family of regular operators $\mathcal{A}_{x}\in\mathcal{R}_{K}(\mathcal{H}_{x},\mathcal{H}^{\prime}_{x})$ with compact resolvents. Suppose that a global trivialization of one of these Hilbert bundles, $\mathcal{H}\cong X\times H$ or $\mathcal{H}^{\prime}\cong X\times H^{\prime}$, is fixed. Then there is a global trivialization of the second bundle taking $\mathcal{A}$ to a Riesz continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$.
We prove Theorems B and C in Section 3.
Self-adjoint operators.   For self-adjoint operators, multiplication $A\mapsto Au$ or $A\mapsto uA$ should be replaced by the conjugation $A\mapsto uAu^{-1}$. Instead of independent trivializations of two Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, as in Theorem C, one should look for a one trivialization of a given Hilbert bundle $\mathcal{H}$. This narrows down available options drastically. In particular, Theorem A has no direct analogue for self-adjoint operators, even invertible ones. A graph continuous family of self-adjoint operators with compact resolvents may fail to be Riesz continuous in any trivialization of $\mathcal{H}$, as we will explain below. The next theorem shows that the obstruction to the existence of such a trivialization has purely local nature.
From local to global trivializations.   Let $\mathcal{A}$ be a family of regular operators $\mathcal{A}_{x}\in\mathcal{R}(\mathcal{H}_{x})$ acting on fibers of a Hilbert bundle $\mathcal{H}$. Following Ivanov [$I_{2}$], we say that a trivialization $\tau\colon\left.\mathcal{H}\right|_{Y}\to Y\times H$ of $\mathcal{H}$ over a subset $Y\subset X$ is adapted to $\mathcal{A}$ if it takes $\mathcal{A}$ to a Riesz continuous map $\tau_{*}\mathcal{A}\colon Y\to\mathcal{R}(H)$. We say that $\mathcal{A}$ is Riesz quasi-continuous if, for every $x\in X$, it admits an adapted trivialization over a neighborhood of $x$.
Theorem D.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$, and let $\mathcal{A}$ be a family of self-adjoint operators with compact resolvents acting on fibers of $\mathcal{H}$. Suppose that $\mathcal{A}$ is Riesz quasi-continuous (that is, admits local adapted trivializations). Then there exists a global trivialization $\tau\colon\mathcal{H}\to X\times H$ adapted to $\mathcal{A}$. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
Applying the last statement of the theorem to the lifting of $\mathcal{A}$ to $X\times[0,1]$, we see that every two adapted trivializations are homotopic and thus the homotopy class of a Riesz continuous map $\tau_{*}\mathcal{A}\colon X\to\mathcal{R}_{K}^{\mathrm{sa}}$ does not depend on the choice of an adapted trivialization $\tau$. For a family $\mathcal{A}$ satisfying assumptions of the theorem, this allows to define the $K^{1}$ index of $\mathcal{A}$ as the index of the norm continuous family $\bm{\chi}(\tau_{*}\mathcal{A})$ of bounded Fredholm self-adjoint operators.
The main case of Theorem D is the case of families consisting of operators which are neither essentially positive nor essentially negative. In this case a similar trivialization theorem was recently proved by different methods and under much weaker continuity assumptions by N. Ivanov. See [$I_{2}$], Theorems 4.6 and 4.7. In combination with Lemma 5.1 from his paper [$I_{3}$], his trivialization theorem implies Theorem D for families consisting of operators which are neither essentially positive nor essentially negative.
Reading Ivanov’s paper [$I_{2}$] led the author to realizing that a part of the original proof of Theorem D can be separated from the rest and has an independent interest. In more details, Ivanov mentioned in the introduction to the first version of [$I_{2}$] that there is a notion of a polarization of a Hilbert bundle standing behind some proofs in [$I_{2}$] (he worked out the details in the current version of [$I_{2}$]). In Section 4 we introduce the notion of a polarization of a Hilbert bundle related to Ivanov’s polarizations in the same way as compact operators are related to finite rank ones, and prove an analogue of his trivialization theorem. See Theorem 4.1.
From graph to Riesz continuity.   Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators with compact resolvents acting on fibers of a Hilbert bundle $\mathcal{H}$, as above. The first obstruction to existence of an adapted trivialization is caused by the fact that $\mathcal{R}^{\mathrm{sa}}_{K}$ is path connected in the graph topology, but has three connected components in the Riesz topology. Two points from different components of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$ (for example, an essentially positive and an essentially negative operator) may be connected by a graph continuous path in $\mathcal{R}_{K}^{\mathrm{sa}}(H)$, but every trivialization of the trivial Hilbert bundle $[0,1]\times H$ will leave this path Riesz discontinuous.
When all the operators $\mathcal{A}_{x}$ are contained in a fixed connected component of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$, there is another kind of local obstruction which we now describe. Let $\lambda\in\mathbb{R}$ be chosen outside of the spectrum of $\mathcal{A}_{x_{0}}$, $x_{0}\in X$. Then the spectral projections $p_{x}=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})$ (here $\operatorname{\mathbb{1}}_{S}$ denotes the characteristic function of a subset $S\subset\mathbb{R}$) give rise to the Hilbert bundles $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$, with fibers $\mathcal{H}^{+}_{x}=\operatorname{Ran}(p_{x})$ and $\mathcal{H}^{-}_{x}=\operatorname{Ker}(p_{x})$, over some neighborhood of $x_{0}$. A local adapted trivialization of $\mathcal{H}$ near $x_{0}$ exists if and only if both $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ are locally trivial near $x_{0}$. See Proposition 5.3.
Suppose first that all the operators $\mathcal{A}_{x}$ are essentially negative. Since the norm and strong topologies coincide on the space of projections of a fixed finite rank, an obstruction to existence of local adapted trivializations near $x_{0}$ arises when the rank of $p_{x}$ is not locally constant near $x_{0}$. In other words, there are eigenvalues of $\mathcal{A}_{x}$ going to +$\infty$+ ∞ when $x\to x_{0}$. If this is not the case and the operators $\mathcal{A}_{x}$ are uniformly bounded from above in a neighborhood of each point $x_{0}\in X$, then $\mathcal{A}$ is Riesz continuous in every trivialization of $\mathcal{H}$. The situation with essentially positive operators is completely the same, but +$\infty$+ ∞ and bounds from above should be replaced by $\infty$- ∞ and bounds from below. See, e.g., Rellich’s example [Ka, Example V-4.14].
Suppose now that all $\mathcal{A}_{x}$ are contained in the third connected component of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}(H)$, ${{}^{r}\mathcal{R}}_{K}^{\star}(H)$, which consists of operators that are neither essentially positive nor essentially negative. Then both Hilbert bundles $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ have infinite rank. By a result of Dixmier and Douady, every separable Hilbert bundle of infinite rank over a paracompact base space of finite covering dimension is trivial. Recall that $X$ is said to have finite covering dimension if there is an $n\in\mathbb{N}$ such that every finite open covering of $X$ has a finite open refinement of order $\leqslant n$.
Let us say that a space has locally finite covering dimension if each point has a neighborhood of finite covering dimension. Using the aforementioned result of Dixmier and Douady, we prove a partial analogue of Theorem C for self-adjoint operators.
Theorem E.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a metrizable space $X$ of locally finite covering dimension. Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators $\mathcal{A}_{x}\in\mathcal{R}^{\star}_{K}(\mathcal{H}_{x})$, with compact resolvents, that are neither essentially positive nor essentially negative. Then $\mathcal{A}$ becomes Riesz continuous in some global trivialization of $\mathcal{H}$. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
Since the group ${{}^{s}{\mathcal{U}}}(H)$ is contractible, Theorem E implies that the identity map ${{}^{r}\mathcal{R}}^{\star}_{K}\to{{}^{g}\mathcal{R}}^{\star}_{K}$ is a weak homotopy equivalence. The stronger result, of a homotopy equivalence of this map, is proven in the previous paper of the author by a different method. See [$P_{3}$], Theorems B and C.
Theorem E does not hold without additional assumptions for general metrizable base spaces, even compact, as we demonstrate by Example 5.5.
Elliptic operators on closed manifolds.   Let us describe an important class of situations where Theorems C and D may be useful. Let $\mathcal{M}\to X$ be a smooth fibration whose fibers $\mathcal{M}_{x}$ are smooth closed manifolds, and let $\mathcal{E}$, $\mathcal{E}^{\prime}$ be smooth vector bundles over $\mathcal{M}$. Let $\mathcal{D}_{x}$ be a continuous (in an appropriate sense) family of elliptic pseudodifferential operators on $\mathcal{M}_{x}$, of positive order, acting from fibers of $\mathcal{E}$ to fibers of $\mathcal{E}^{\prime}$. A local trivialization of $\mathcal{M}$ turns $\mathcal{D}$ to a family of elliptic operators acting between fibers of fixed vector bundles $E$ and $E^{\prime}$ over a fixed closed manifold $M$. The bounded transform takes this new family to a continuous family of zeroth order elliptic operators, which is norm continuous as a family of bounded operators $L^{2}(M;E)\to L^{2}(M;E^{\prime})$. Hence the family of unbounded operators $\mathcal{D}_{x}$ is Riesz continuous in this local trivialization. However, the fibration $\mathcal{M}\to X$ is in general non-trivial, and thus no such “geometric” global trivialization exists in the general case. The structure group of the corresponding Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, where $\mathcal{H}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$ and $\mathcal{H}^{\prime}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}^{\prime}_{x})$, has the topology coarser than the norm topology, so Kuiper’s theorem cannot be used in order to trivialize $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Dixmier-Douady theorem, on the other hand, can be used, but trivializations it provides do not take $\mathcal{D}$ to a Riesz continuous family of operators.
Theorem C closes this gap, showing that such a family $\mathcal{D}$ can be made Riesz continuous by appropriate global trivializations of the Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, though at the cost of losing the differential and geometric structure. Theorem D solves the same problem for families of self-adjoint elliptic operators acting on fibers of $\mathcal{E}$.
The author arrived at Theorem D in an attempt to understand a step in the proof of a result of Melrose and Piazza. See [$MP_{1}$], Proposition 1. Melrose and Piazza deal with a family of self-adjoint differential operators of first order over a smooth fibration $\mathcal{M}\to X$, as above, and show that if the $K^{1}$-index of the family vanishes, then the family admits a spectral section. In their proof, they claim that the $L^{2}$ spaces of the fibers, $L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$, can be identified so that all operators act on a fixed Hilbert space $H$. Unfortunately, they did not explain how such an identification can be obtained. The author was not able to find a simple explanation of this step. Our Theorem D, as well as recent results [$I_{2}$, Theorems 4.5 or 4.6] or [$I_{3}$] of N. Ivanov, allows to overcome this difficulty. Similarly, Theorem C allows to pass to a fixed Hilbert space in the proof of [$MP_{2}$, Proposition 2], an odd analog of [$MP_{1}$, Proposition 1].
Elliptic operators on manifolds with boundary.   When the fibers $\mathcal{M}_{x}$ are compact manifolds with boundary, the situation becomes more complicated. In this case $\mathcal{D}_{x}$ is not even a closed operator usually. One needs to add an elliptic boundary condition $\mathcal{B}_{x}$ to $\mathcal{D}_{x}$, and then such a pair $(\mathcal{D}_{x},\mathcal{B}_{x})$ determines a regular operator with compact resolvents. Taking a local trivialization of $\mathcal{M}$ as above, one obtains a family of boundary value problems $(\mathcal{D}_{x},\mathcal{B}_{x})$ on a fixed manifold $M$. If both $\mathcal{D}_{x}$ and $\mathcal{B}_{x}$ depend continuously on $x$ in an appropriate sense, then the corresponding family of regular operators with compact resolvents is graph continuous. See [$P_{1}$, Appendix A.5] or [BZ, Corollary A.6.3]. Theorem C shows that such a family $\mathcal{D}$ can be made Riesz continuous by appropriate global trivializations of the corresponding Hilbert bundles.
Suppose now that we have a family of self-adjoint elliptic boundary value problems. As before, a local trivialization of $\mathcal{M}$ gives rise to a family of boundary value problems on a fixed manifold $M$. Unfortunately, there are no criteria of Riesz continuity for families of unbounded operators arising from boundary value problems, except for several special cases. Moreover, there are simple examples of such families that are not Riesz quasi-continuous, see e.g. Rellich’s example. Theorem E, however, requires only graph continuity of a family of unbounded operators, so it can be applied to families of elliptic self-adjoint boundary value problems, provided that they are neither essentially positive nor essentially negative and the space of parameters is good enough.
Acknowledgment.   I am grateful to N. V. Ivanov for attracting my attention to the work of Dixmier and Douady [DD] and for helpful remarks.
2 Preliminaries
Grassmanian.   The set $\operatorname{Gr}(H)$ of closed subspaces of $H$ is called the Grassmanian. An orthogonal projection may be identified with its range, which provides a bijection between $\operatorname{Gr}(H)$ and $\mathcal{P}(H)$. The Grassmanian is equipped with the topology (and even metric) induced by this bijection from the norm on the space $\mathcal{P}(H)$ of projections. We will use closed subspaces and projections by turn, depending on what is more convenient at the moment.
Canonical decomposition.   Recall a standard construction that will be used throughout the paper. The trivial Hilbert bundle over $\operatorname{Gr}(H)$ with the fiber $H$ is canonically decomposed into the direct sum
$\operatorname{Gr}(H)\times H=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$
of two Hilbert bundles, whose fibers over $L\in\operatorname{Gr}$ are $\mathcal{H}^{+}_{L}=L$ and $\mathcal{H}^{-}_{L}=L^{\bot}$, the orthogonal complement of $L$ in $H$. This decomposition is locally trivial in the following sense: for every $L_{0}\in\operatorname{Gr}(H)$ there is a continuous map $g\colon B\to\mathcal{U}(H)$, where $B$ is the open ball of radius $1$ around $L_{0}$, such that $L=g_{L}L_{0}$ for every $L\in B$. In terms of projections, this is written as $p=g_{p}p_{0}g_{p}^{*}$ for $\left\|p-p_{0}\right\|<1$, where $p_{0}$ is the projection onto $L_{0}$. See the proof of [Ka, Theorem I.6.32] for the construction of such a map $g$.
An important subspace of $\operatorname{Gr}(H)$ is
$\operatorname{Gr}^{\star}(H)=\left\{L\in\operatorname{Gr}(H)\,|\,\dim L=\infty=\operatorname{codim}L\right\}.$
Let $L_{0}\in\operatorname{Gr}^{\star}(H)$ be a fixed subspace. The map $\mathcal{U}(H)\to\operatorname{Gr}^{\star}(H)$, $u\mapsto uL_{0}$ is a locally trivial principal bundle with the contractible structure group $\mathcal{U}(L_{0})\times\mathcal{U}(L_{0}^{\bot})$ over a paracompact base space and therefore has a section
(2.1) $g\colon\operatorname{Gr}^{\star}(H)\to\mathcal{U}(H),\;\;\text{so that}\;\;L=g_{L}L_{0}\;\;\text{for all}\;\;L\in\operatorname{Gr}^{\star}(H).$
It follows that the restrictions of $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ to $\operatorname{Gr}^{\star}$ are trivial Hilbert bundles with the structure group $\mathcal{U}$. In addition, contractibility of the total space $\mathcal{U}(H)$ and the fiber $\mathcal{U}\times\mathcal{U}$ implies contractibility of the base space $\operatorname{Gr}^{\star}$. See [AS, proof of Lemma 3.6].
Bounded transform.   Let us recall the definition and the main properties of the bounded transform. See, e.g., [Sc], Theorem 7.5 for the detail.
For a regular operator $A\colon H\to H^{\prime}$, the operator $1+A^{*}A$ is regular, self-adjoint, and surjective; its inverse $(1+A^{*}A)^{-1}$ is a positive bounded operator. The bounded transform (or the Riesz map)
$\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{B}(H,H^{\prime}),\quad\bm{\chi}(A)=A(1+A^{*}A)^{-1/2},$
defines the inclusion of the set $\mathcal{R}(H,H^{\prime})$ of regular operators to the closed unit ball
$\mathcal{D}(H,H^{\prime})=\left\{a\in\mathcal{B}(H,H^{\prime})\,|\,\left\|a\right\|\leqslant 1\right\}.$
The image of $\bm{\chi}$ consists of strict contractions, that is, operators $a\in\mathcal{D}$ such that $1-a^{*}a$ is injective; this image is dense in $\mathcal{D}$. For $a=\bm{\chi}(A)$ the following identities hold:
(2.2) $(1+A^{*}A)^{-1}=1-a^{*}a\;\;\text{and}\;\;A=a(1-a^{*}a)^{-1/2}.$
In particular, a regular operator $A$ has compact resolvents if and only if $\bm{\chi}(A)$ is essentially unitary. If a regular operator $A$ is self-adjoint, then so is $\bm{\chi}(A)$; more generally, $\bm{\chi}(A^{*})=\bm{\chi}(A)^{*}$.
Graphs of operators.   Let $\Gamma_{A}\subset H\oplus H^{\prime}$ denote the graph of a regular operator $A\colon H\to H^{\prime}$. There is a canonical isometric isomorphism
$\Gamma_{A}\to H,\quad\xi\oplus A\xi\mapsto\sqrt{1+A^{*}A}\,\xi\;\text{ for }\xi\in H.$
The inverse isomorphism $H\to\Gamma_{A}$, which we consider as an isometry $\mathbf{w}_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$, is given by the formula
(2.3) $\mathbf{w}_{A}=(1+A^{*}A)^{-1/2}\oplus A(1+A^{*}A)^{-1/2}=(1+A^{*}A)^{-1/2}\oplus\bm{\chi}(A).$
The orthogonal projection onto $\Gamma_{A}$ is
(2.4) $\mathbf{p}_{A}=\mathbf{w}_{A}\mathbf{w}_{A}^{*}=\begin{pmatrix}(1+A^{*}A)^{-1}&(1+A^{*}A)^{-1}A^{*}\\ A(1+A^{*}A)^{-1}&1-(1+AA^{*})^{-1}\end{pmatrix}\in\mathcal{P}(H\oplus H^{\prime}).$
In terms of $a=\bm{\chi}(A)$ this is written as follows:
$\mathbf{w}_{A}=\tilde{\mathbf{w}}_{a}=\sqrt{1-a^{*}a}\oplus a,\quad\quad\mathbf{p}_{A}=\tilde{\mathbf{p}}_{a}=\tilde{\mathbf{w}}_{a}\tilde{\mathbf{w}}_{a}^{*}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix},$
so $\mathbf{p}$ factors through a continuous map $\tilde{\mathbf{p}}\colon\mathcal{D}(H,H^{\prime})\to\mathcal{P}(H\oplus H^{\prime})$. Restricting $\tilde{\mathbf{p}}$ to the subspace ${{}^{r}\mathcal{R}}$ of $\mathcal{D}$, we see that the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is included into the following commutative square and thus is continuous:
[…]
Fredholm operators.   Similar to the bounded case, a regular operator is called Fredholm if its range is closed and its kernel and cokernel are finite-dimensional. The bounded transform $\bm{\chi}(A)$ is equal to the composition of the isomorphism $\mathbf{w}_{A}\colon H\to\Gamma_{A}$ and the restriction $p^{\prime}_{A}\colon\Gamma_{A}\to H^{\prime}$ of the orthogonal projection $p^{\prime}\colon H\oplus H^{\prime}\to H^{\prime}$. The range of $p^{\prime}_{A}$ coincides with the range of $A$. The kernel of $p^{\prime}_{A}$ is isomorphic to the kernel of $A$. Therefore, a regular operator $A$ is Fredholm if and only if $\bm{\chi}(A)$ is Fredholm.
Graph-to-strong* continuity of the bounded transform.   By the definition of Riesz topology, the bounded transform $\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{D}(H,H^{\prime})$ is Riesz-to-norm continuous. In the proof of Theorem A we also need another type of continuity of $\bm{\chi}$.
Recall that a family of bounded operators $a_{x}\in\mathcal{B}(H,H^{\prime})$ is said to be continuous in the strong operator topology (or strongly continuous) if the family $a_{x}\xi$ of vectors in $H^{\prime}$ is continuous for every $\xi\in H$. A family $a_{x}$ is said to be continuous in the strong* operator topology if both families $a_{x}$ and $a_{x}^{*}$ are strongly continuous.
Proposition 2.1.
The bounded transform is graph-to-strong* continuous.
Proof. Since $\mathbf{p}_{A^{*}}=1-J\mathbf{p}_{A}J^{*}$, where $J=\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)$, the map $A\mapsto A^{*}$ is graph continuous. Since $\bm{\chi}(A)^{*}=\bm{\chi}(A^{*})$, it is enough to prove graph-to-strong continuity of $\bm{\chi}$.
For $A\in\mathcal{R}(H,H^{\prime})$ we denote
(2.5) $a=\bm{\chi}(A),\quad q=(1+A^{*}A)^{-1/2}=(1-a^{*}a)^{1/2},\quad b=aq.$
The projection $\mathbf{p}_{A}$ is given by the formula
(2.6) $\mathbf{p}_{A}=\tilde{\mathbf{p}}_{a}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&qa^{*}\\ aq&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&b^{*}\\ b&aa^{*}\end{pmatrix}.$
Since the map $A\mapsto b$ is given by the $(2,1)$-component of $\mathbf{p}_{A}$, it is graph-to-norm continuous. Similarly, the map $A\mapsto q$ is the square root of the $(1,1)$-component of $\mathbf{p}_{A}$ and thus is graph-to-norm continuous.
Choose $\xi\in H$ and $A_{0}\in\mathcal{R}(H,H^{\prime})$. Let $a_{0}$, $q_{0}$, $b_{0}$ be determined by $A_{0}$ as in (2.5). We need to show that the map ${{}^{g}\mathcal{R}}(H,H^{\prime})\to H^{\prime}$, $A\mapsto\bm{\chi}(A)\xi$ is continuous at $A=A_{0}$. Let $\varepsilon>0$. The operator $q_{0}$ is self-adjoint and injective and thus has dense range, so one can find $\eta\in H$ such that $\left\|q_{0}\eta-\xi\right\|<\varepsilon$. Let $V$ be a neighborhood of $A_{0}$ in ${{}^{g}\mathcal{R}}(H,H^{\prime})$ such that $\left\|(q-q_{0})\eta\right\|<\varepsilon$ and $\left\|(b-b_{0})\eta\right\|<\varepsilon$ for every $A\in V$. Then
$\left\|a\xi-a_{0}\xi\right\|\;<\;\left\|\left(a-a_{0}\right)q_{0}\eta\right\|+2\varepsilon\;\leqslant\;\left\|a\left(q_{0}-q\right)\eta\right\|+\left\|\left(aq-a_{0}q_{0}\right)\eta\right\|+2\varepsilon$
$\leqslant\;\left\|\left(q-q_{0}\right)\eta\right\|+\left\|\left(b-b_{0}\right)\eta\right\|+2\varepsilon<4\varepsilon$
for every $A\in V$. It follows that $\bm{\chi}$ is graph-to-strong continuous at an arbitrary point $A_{0}\in\mathcal{R}$ and thus on the whole $\mathcal{R}$. $\square$
Remark.
In fact, this reasoning provides a stronger result. It concerns another natural topology on $\mathcal{R}(H,H^{\prime})$, which is is strictly coarser then the graph topology, induced by the inclusion $\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow{{}^{s}{\mathcal{P}}}(H\oplus H^{\prime})$ from the strong operator topology on the space of projections. On the subset of self-adjoint operators it coincides with the strong resolvent topology, which is induced by the map $A\mapsto(A+i)^{-1}$, or equivalently by the map $A\mapsto(A-i)^{-1}$, from the strong operator topology on bounded operators. The reasoning in the proof of Proposition 2.1 shows that the bounded transform $\bm{\chi}\colon\mathcal{R}\to\mathcal{D}$ is continuous with respect to this coarser topology on $\mathcal{R}$ and the strong* topology on $\mathcal{D}$.
Proposition 2.1 follows from a similar result for self-adjoint operators and the restriction of $\bm{\chi}$ to $\mathcal{R}^{\mathrm{sa}}$, $A\mapsto A(1+A^{2})^{-1/2}$, by the standard trick of passing from $A$ to the self-adjoint operator $\left(\begin{smallmatrix}0&A^{*}\\ A&0\end{smallmatrix}\right)$. For self-adjoint operators, a similar reasoning provides continuity of an arbitrary function $f\in C_{b}(\mathbb{R})$ with respect to the strong resolvent topology on $\mathcal{R}^{\mathrm{sa}}$ and the strong topology on $\mathcal{D}$. A different proof of this fact based on the spectral theorem is given by [RS], Theorem VIII.20(b). A similar result for compact spaces of parameters is given by [W], Proposition 1.1 of Ch. Wahl.
Action of the unitary groups.   The unitary groups $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ act on $\mathcal{R}(H,H^{\prime})$ by the right and left multiplication respectively. These actions preserve the subsets $\mathcal{R}_{F}$ and $\mathcal{R}_{K}$. For $u\in\mathcal{U}(H^{\prime})$ and $v\in\mathcal{U}(H)$ we have
$\bm{\chi}(uA)=u\bm{\chi}(A),\quad\bm{\chi}(Av)=\bm{\chi}(A)v,\quad\Gamma_{uA}=(1\oplus u)\Gamma_{A},\quad\Gamma_{Av}=(v^{-1}\oplus 1)\Gamma_{A},$
so the actions of $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ (equipped with the norm topology) are continuous on both ${{}^{r}\mathcal{R}}$ and ${{}^{g}\mathcal{R}}$.
Recall that ${{}^{s}{\mathcal{U}}}(H)$, the unitary group equipped with the strong operator topology, is a topological group. Indeed, the multiplication map $\mathcal{B}(H)\times\mathcal{B}(H)\to\mathcal{B}(H)$, $(a,b)\mapsto ab$, is strongly continuous on $S\times\mathcal{B}(H)$ for every bounded subset $S$ of $\mathcal{B}(H)$. In particular, it is strongly continuous on $\mathcal{U}(H)\times\mathcal{U}(H)$. For every $u,u_{0}\in\mathcal{U}(H)$ and $\xi\in H$,
$\left\|(u^{-1}-u_{0}^{-1})\xi\right\|=\left\|u(u^{-1}-u_{0}^{-1})\xi\right\|=\left\|(u-u_{0})u_{0}^{-1}\xi\right\|,$
so the inverse map $\mathcal{U}(H)\to\mathcal{U}(H)$, $u\mapsto u^{-1}=u^{*}$, is also strongly continuous.
The actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{r}\mathcal{R}}$ and ${{}^{g}\mathcal{R}}$ are not continuous. However, continuity of their actions on $\mathcal{K}(H\oplus H^{\prime})$ provides the following result.
Proposition 2.2.
The actions of the groups ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$ are continuous
Proof. The graph of an operator with compact resolvents belongs to the restricted Grassmanian
(2.7) $\operatorname{Gr}_{K}(H\oplus H^{\prime})\cong\mathcal{P}_{K}(H\oplus H^{\prime})=\left\{p\in\mathcal{P}(H\oplus H^{\prime})\,|\,p-p_{\infty}\text{ is compact}\right\},$
where $p_{\infty}\in\mathcal{P}(H\oplus H^{\prime})$ is the projection onto $0\oplus H^{\prime}$. The restriction of the embedding $\mathbf{p}\colon{{}^{g}\mathcal{R}}\hookrightarrow\mathcal{P}$ to operators with compact resolvents provides the embedding $\mathbf{p}\colon{{}^{g}\mathcal{R}}_{K}\hookrightarrow\mathcal{P}_{K}$. The left action of ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{g}\mathcal{R}}(H,H^{\prime})$ lifts to the action on $\mathcal{P}(H\oplus H^{\prime})$:
$(u,p)\mapsto(1\oplus u)p(1\oplus u^{-1})=(1\oplus u)(p-p_{\infty})(1\oplus u)^{-1}+p_{\infty}.$
Since the action of ${{}^{s}{\mathcal{U}}}(H\oplus H^{\prime})$ on $\mathcal{K}(H\oplus H^{\prime})$ by conjugation is continuous, the action of ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on $\mathcal{P}_{K}(H\oplus H^{\prime})$ is also continuous, as well as its restriction to ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$. The right action of ${{}^{s}{\mathcal{U}}}(H)$ on ${{}^{g}\mathcal{R}}_{K}$ is continuous by the same argument. $\square$
3 Non-self-adjoint operators
[…]
we obtain a desired map $v\colon{{}^{g}\mathcal{R}}\to{{}^{s}{\mathcal{U}}}(H)$. If $\Gamma$ is considered as a Hilbert bundle over ${{}^{r}\mathcal{R}}$, then $\mathbf{w}$ becomes a trivialization in the category of bundles with the structure group $\mathcal{U}(H)$, so $v$ is continuous as a map ${{}^{r}\mathcal{R}}\to\mathcal{U}(H)$ as well.
Isometries.   Let
=$\mathcal{I}=\left\{u\in\mathcal{B}(H,H\oplus H^{\prime})\,|\,u^{*}u=\operatorname{Id}_{H}\right\}$
be the space of isometries (i.e., isometric embeddings) $H\to H\oplus H^{\prime}$ equipped with the usual norm topology. The canonical map
$\bm{\pi}\colon\mathcal{I}\to\operatorname{Gr}(H\oplus H^{\prime}),\quad u\mapsto\operatorname{Ran}(u).$
takes an isometry $u$ to its range; the corresponding projection is $uu^{*}$. The image $\bm{\pi}(\mathcal{I})$ of this map consists of infinite-dimensional subspaces of $H\oplus H^{\prime}$. We will need the following simple corollary of Kuiper’s theorem [Ku] about contractibility of $\mathcal{U}(H)$.
Lemma 3.1.
The map $\bm{\pi}\colon\mathcal{I}\to\bm{\pi}(\mathcal{I})$ is a principal $\mathcal{U}(H)$-bundle admitting a continuous section (and thus is trivial).
Proof. The group $\mathcal{U}(H)$ acts on $\mathcal{I}$ by the right multiplication. Every two isometries with the same range differ by an uniquely defined unitary operator $H\to H$, so this action turns $\bm{\pi}$ into a principal $\mathcal{U}(H)$-bundle. This bundle has a section by the same standard arguments as were used in the beginning of Section 2. Indeed, for every $u_{0}\in\mathcal{I}$ and $L_{0}=\bm{\pi}(u_{0})$ there is a continuous map $g$ from a neighborhood $B$ of $L_{0}$ to $\mathcal{U}(H\oplus H^{\prime})$ such that $L=g_{L}L_{0}$ for $L\in B$. The rule $L\mapsto g_{L}\cdot u_{0}$ determines a local section of $\bm{\pi}$ over $B$. Hence $\bm{\pi}$ is a locally trivial principal bundle over $\bm{\pi}(\mathcal{I})$ with the contractible structure group $\mathcal{U}(H)$. The base space $\bm{\pi}(\mathcal{I})$ is metric and thus paracompact. Therefore, $\bm{\pi}$ has a global section. $\square$
Proof of Theorem A.   The graphs of regular operators have infinite dimension, so ${{}^{g}\mathcal{R}}\subset\bm{\pi}(\mathcal{I})\subset\operatorname{Gr}(H\oplus H^{\prime})$. By Lemma 3.1, the restriction of $\bm{\pi}$ to ${{}^{g}\mathcal{R}}$ has a section. For the rest of the proof we fix such a section $u\colon{{}^{g}\mathcal{R}}\to\mathcal{I}$. In other words, $u$ is a graph-to-norm continuous map taking a regular operator $A$ to an isometry $u_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$.
Let $p^{\prime}\colon H\oplus H^{\prime}\to H^{\prime}$ be the projection onto the second summand. The composition
$\varphi_{A}=p^{\prime}\cdot u_{A}\colon H\to H^{\prime}$
is a contraction, and the corresponding map $\varphi\colon{{}^{g}\mathcal{R}}(H,H^{\prime})\to\mathcal{D}(H,H^{\prime})$ is continuous.
The maps defined above are not canonical: they depend on the choice of a section $u$. On the other hand, formula (2.3) provides a canonical isometry $\mathbf{w}_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$:
$\mathbf{w}_{A}=(1+A^{*}A)^{-1/2}\oplus\bm{\chi}(A)=\sqrt{1-a^{*}a}\oplus a,\;\;\text{where}\;\;a=\bm{\chi}(A).$
The corresponding contraction $p^{\prime}\cdot\mathbf{w}_{A}\colon H\to H^{\prime}$ coincides with the bounded transform $\bm{\chi}(A)$ of $A$. In contrast with $\varphi$, it is not continuous as a map ${{}^{g}\mathcal{R}}\to\mathcal{D}$.
Two isometries $u_{A},\mathbf{w}_{A}\colon H\to H\oplus H^{\prime}$ have the same range $\Gamma_{A}$ and thus differ by a unitary operator
$v_{A}=\mathbf{w}_{A}^{*}\cdot u_{A}\in\mathcal{U}(H),\;\;\text{so that}\;\;u_{A}=\mathbf{w}_{A}\cdot v_{A}.$
We claim that the map $v\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H)$ defined by this formula satisfies conclusions of the theorem. Indeed, for $\Phi_{A}=A\cdot v_{A}$ we have
$\bm{\chi}(\Phi_{A})=\bm{\chi}(A\cdot v_{A})=\bm{\chi}(A)\cdot v_{A}=p^{\prime}\cdot\mathbf{w}_{A}\cdot v_{A}=p^{\prime}\cdot u_{A}=\varphi_{A},$
so $\bm{\chi}\circ\Phi=\varphi\colon\mathcal{R}(H,H^{\prime})\to\mathcal{B}(H,H^{\prime})$ is graph-to-norm continuous and thus $\Phi$ is graph-to-Riesz continuous. By the construction, $u\colon\mathcal{R}\to\mathcal{I}$ is graph-to-norm (and thus also Riesz-to-norm) continuous. The coisometry $\mathbf{w}_{A}^{*}\colon H\oplus H^{\prime}\to H$ is given by the formula
$\mathbf{w}_{A}^{*}(\xi\oplus\eta)\;=\;(1+A^{*}A)^{-1/2}\;\xi+\bm{\chi}(A)^{*}\,\eta\;=\;\sqrt{1-a^{*}a}\;\xi+a^{*}\,\eta\;\;\;\text{for}\;\;\xi\in H,\;\eta\in H^{\prime}.$
Clearly, the maps $A\mapsto\mathbf{w}_{A}^{*}$ and $A\mapsto v_{A}=\mathbf{w}_{A}^{*}\cdot u_{A}$ are Riesz-to-norm continuous.
The first component of $\mathbf{w}^{*}$, $A\mapsto(1+A^{*}A)^{-1/2}$, is the square root of the $(1,1)$-component of $\mathbf{p}_{A}$ and thus is even graph-to-norm continuous. The second component $A\mapsto\bm{\chi}(A)^{*}$ is graph-to-strong continuous by Proposition 2.1. Therefore, the map $A\mapsto\mathbf{w}_{A}^{*}$ is graph-to-strong continuous. The multiplication map $(b,c)\mapsto b\cdot c$ is strongly continuous on uniformly bounded subsets of operators, and the norms of $\mathbf{w}_{A}^{*}$ and $u_{A}$ are bounded by $1$. It follows that the map $A\mapsto v_{A}=\mathbf{w}_{A}^{*}\cdot u_{A}$ is graph-to-strong continuous. This completes the proof of the theorem. $\square$
Remark.
Our construction of $v$ depends on the choice of a section $u$. Since $\bm{\pi}$ is a principal $\mathcal{U}(H)$-bundle, any other choice of a section differs from $u$ by the multiplication by a graph-to-norm continuous map ${{}^{g}\mathcal{R}}(H,H^{\prime})\to\mathcal{U}(H)$. Therefore, our construction determines $v$ up to the right multiplication by such a map.
Proof of Theorem B.   In order to prove the first part of the theorem we need to show that $\Phi$ is homotopic to the identity as a self-map of both ${{}^{r}\mathcal{R}}_{K}$ and ${{}^{g}\mathcal{R}}_{K}$. For the second part we need to prove a similar result for ${{}^{r}\mathcal{R}}_{F}$ and ${{}^{g}\mathcal{R}}_{F}$.
By Kuiper’s theorem, the group $\mathcal{U}(H)$ is contractible in the norm topology, so the map $v\colon\mathcal{U}\to\mathcal{U}$ is homotopic to the identity map $\mathcal{U}\to\mathcal{U}$. Since the action of $\mathcal{U}(H)$ on ${{}^{r}\mathcal{R}}(H,H^{\prime})$ is continuous, a homotopy between these two maps provides a homotopy between $\Phi\colon{{}^{r}\mathcal{R}}_{K}\to{{}^{r}\mathcal{R}}_{K}$ and $\operatorname{Id}\colon{{}^{r}\mathcal{R}}_{K}\to{{}^{r}\mathcal{R}}_{K}$, and a similar homotopy for ${{}^{r}\mathcal{R}}_{F}\to{{}^{r}\mathcal{R}}_{F}$.
By a result of Dixmier and Douady [DD, Lemma 3], the group ${{}^{s}{\mathcal{U}}}(H)$ is contractible, so the map $v\colon{{}^{s}{\mathcal{U}}}\to{{}^{s}{\mathcal{U}}}$ is homotopic to the identity map ${{}^{s}{\mathcal{U}}}\to{{}^{s}{\mathcal{U}}}$. The action of ${{}^{s}{\mathcal{U}}}(H)$ on ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$ is continuous (see Proposition 2.2), so a homotopy between these two maps provides a homotopy between the maps $\Phi\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$. Therefore, the restriction of $\Phi$ to the subspace of operators with compact resolvents is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$.
In order to finish the proof of the Fredholm part, we need to show that the maps $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ are homotopic. This part is more delicate because the action of ${{}^{s}{\mathcal{U}}}$ on ${{}^{g}\mathcal{R}}_{F}$ is not continuous, so we cannot directly apply a contraction of ${{}^{s}{\mathcal{U}}}$ for this purpose. However, we can use a part of [$P_{3}$, Theorem A] which states that the embedding ${{}^{g}\mathcal{R}}_{K}\hookrightarrow{{}^{g}\mathcal{R}}_{F}$ is a homotopy equivalence. We already proved that the maps $\Phi\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ are homotopic; it follows that $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ are also homotopic. Therefore, the restriction of $\Phi$ to the subspace of Fredholm operators is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$. $\square$
Proof of Theorem C.   Suppose that a global trivialization of $\mathcal{H}^{\prime}$ is fixed. Choose an arbitrary trivialization $\tau\colon\mathcal{H}\to X\times H$ of $\mathcal{H}$. In these trivializations $\mathcal{A}$ becomes a graph continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$. Let $v\colon{{}^{g}\mathcal{R}}(H,H^{\prime})\to{{}^{s}{\mathcal{U}}}(H)$ be a map provided by Theorem A. Then the map $X\to{{}^{s}{\mathcal{U}}}(H)$, $x\mapsto v(\mathcal{A}_{x})$ is continuous, so
$\tau_{v}\colon(x,\xi)\mapsto(x,v(\mathcal{A}_{x})\xi)$
is an automorphism of the trivial Hilbert bundle $X\times H$. The inverse automorphism $\tau_{v}^{-1}$ takes $\mathcal{A}$ to a Riesz continuous map
$\mathcal{A}^{\prime}\colon X\to\mathcal{R}_{K}(H,H^{\prime}),\quad\mathcal{A}^{\prime}_{x}=\mathcal{A}_{x}\cdot v(\mathcal{A}_{x}).$
It follows that the composition $\tau_{v}^{-1}\circ\tau$ satisfies conclusion of the theorem.
If instead a global trivialization of $\mathcal{H}$ is fixed, then one can interchange $\mathcal{H}$ and $\mathcal{H}^{\prime}$ by passing to the adjoint operators. Since the map $A\mapsto A^{*}$ is both graph and Riesz continuous, it follows from the first part of the proof that there is a global trivialization of $\mathcal{H}^{\prime}$ taking $\mathcal{A}$ to a Riesz continuous map. $\square$
$K^{0}$ index for graph continuous families.   By a classical result of Atiyah and Jänich, the space $\mathcal{B}_{F}(H)$ of bounded Fredholm operators is a classifying space for the functor $K^{0}$. See [A], Theorem A1. For an arbitrary parameter space $X$, $K^{0}(X)$ might be defined as the set $[X,\mathcal{B}_{F}]$ of homotopy classes of maps $X\to\mathcal{B}_{F}(H)$, with the group structure on $[X,\mathcal{B}_{F}]$ introduced in the standard way. The class of a continuous map $\alpha\colon X\to\mathcal{B}_{F}$ in $K^{0}(X)$ is called the index of $\alpha$:
$\operatorname{ind}(\alpha)=[\alpha]\in[X,\mathcal{B}_{F}]=K^{0}(X).$
It can be easily seen, using linear homotopies, that the bounded transform map $\bm{\chi}\colon{{}^{r}\mathcal{R}}_{F}\to\mathcal{B}_{F}$ is homotopy inverse to the embedding $\mathcal{B}_{F}\hookrightarrow{{}^{r}\mathcal{R}}_{F}$, so both these maps are homotopy equivalencies. Therefore, the index of a Riesz continuous family of Fredholm operators is naturally defined as the index of its bounded transform:
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{r}\mathcal{R}}_{F}.$
Let now $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ be a graph continuous family of Fredholm operators. The space ${{}^{g}\mathcal{R}}_{F}$ is also a classifying space for the functor $K^{0}$, see [J], Theorem 3.5(i) or [$P_{3}$], Theorem A. Hence the homotopy class of $\mathcal{A}$ determines an element of $K^{0}(X)$ which can be called the index of $\mathcal{A}$:
(3.1) $\operatorname{ind}(\mathcal{A})=[\mathcal{A}]\in[X,{{}^{g}\mathcal{R}}_{F}]=K^{0}(X).$
By the definition, two graph continuous families of Fredholm operators are homotopic if and only if their indices coincide.
Theorems A and B of the present paper allow to relate the index of $\mathcal{A}$ given by (3.1) to the Fredholm index of bounded operators as follows. The composition of $\mathcal{A}$ with a map $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ provided by Theorem A is Riesz continuous, so $\bm{\chi}\circ\Phi\circ\mathcal{A}$ is a norm continuous family of bounded Fredholm operators. Theorem B provides the equality
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\Phi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F},$
which can be used as the definition of the index of $\mathcal{A}$. Note that Theorem B implies that the homotopy class of $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{r}\mathcal{R}}_{F}$ does not depend on the choice of $v$, so $\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$ is also independent of this choice. Equivalently, the index can be defined in terms of graphs of operators as described in the Introduction.
For a Riesz continuous family $\mathcal{A}\colon X\to\mathcal{R}_{F}$ (in particular, for a norm continuous family of bounded Fredholm operators) the two definitions of the index coincide. Indeed, $\Phi(\mathcal{A}_{x})=\mathcal{A}_{x}\cdot v(\mathcal{A}_{x})$, so $\Phi\circ\mathcal{A}$ differs from $\mathcal{A}$ by the multiplication by a map $v\circ\mathcal{A}\colon X\to\mathcal{U}(H)$. If $\mathcal{A}$ is Riesz continuous, then $v\circ\mathcal{A}$ is norm continuous. Since $\mathcal{U}(H)$ is contractible, $\Phi\circ\mathcal{A}$ and $\mathcal{A}$ are homotopic as maps from $X$ to ${{}^{r}\mathcal{R}}_{F}$ and thus their indices coincide.
Recently N. Ivanov [$I_{1}$, $I_{2}$] developed a new approach to the analytical index of families. He introduced the notion of a “Fredholm family” and defined the $K^{0}(X)$-valued index, which we will denote by $\operatorname{Ind}$, for Fredholm families parametrized by points of a paracompact space $X$. See [$I_{2}$], Section 3. A Fredholm family is a family of Fredholm operators $\mathcal{A}_{x}$ such that the finite-dimensional subspaces $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$ and $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}^{*}_{x}|)$, as well as the restriction of $\mathcal{A}_{x}$ to $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$, depend continuously on $x$ for an appropriate locally chosen $\varepsilon>0$, without any additional continuity assumptions. Ivanov’s index $\operatorname{Ind}$ is defined both for operators acting on a fixed Hilbert space and on fibers of a Hilbert bundle; it is invariant under isomorphisms of Hilbert bundles and under passing to the bounded transform.
Let $\mathcal{A}$ be a graph continuous family of Fredholm operators in $H$ parametrized by points of a paracompact space $X$. Then $\mathcal{A}$ is a Fredholm family in Ivanov’s sense, and we claim that our definition of the index of $\mathcal{A}$ provides the same element of $K^{0}(X)$ as Ivanov’s definition. Indeed, the properties of $\operatorname{Ind}$ mentioned above imply the equality $\operatorname{Ind}(\mathcal{A})=\operatorname{Ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$. Consider the norm continuous family $\alpha=\bm{\chi}\circ\Phi\circ\mathcal{A}$ of bounded Fredholm operators in $H$. By Theorem 7.2 of [$I_{2}$], $\operatorname{Ind}\alpha$ coincides with the index of $\alpha$ defined by Segal in [Se]. The last index is equal to the homotopy class $[\alpha]$ of $\alpha$ in $[X,\mathcal{B}_{F}(H)]=K^{0}(X)$ by [Se], Theorem 5.1(b). Taking all this together, we obtain
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\alpha)=[\alpha]=\operatorname{Ind}(\alpha)=\operatorname{Ind}(\mathcal{A})\in K^{0}(X).$
4 Polarizations
Polarizations of a Hilbert space.   A decomposition of a Hilbert space $H$ into the orthogonal sum $H=H^{+}\oplus H^{-}$ of two closed infinite-dimensional subspaces is called a polarization of $H$. It is determined by the subspace $H^{+}$ or equivalently by the orthogonal projection $p\in\mathcal{P}^{\star}(H)$ onto $H^{+}$; we will identify polarizations with the corresponding subspaces/projections.
There are two natural equivalence relations for polarizations of $H$, which we call “finite” and “compact”. Two polarizations $p$ and $q$ of $H$ are said to be compactly compatible if $p-q$ is a compact operator; they are said to be finitely compatible if $p-q$ is a finite rank operator (that is, the corresponding subspaces are commensurable).
The compact restricted Grassmanian corresponding to the polarization given by a projection $p$ is defined as the space of all polarizations compatible with $p$:
$\operatorname{Gr}_{K}=\operatorname{Gr}_{K}(H;p)=\left\{q\in\mathcal{P}^{\star}(H)\,|\,q-p\in\mathcal{K}(H)\right\}.$
It is equipped with the norm topology induced from the natural inclusion $\operatorname{Gr}_{K}(H;p)\hookrightarrow\mathcal{P}(H)$, or equivalently the inclusion $\operatorname{Gr}_{K}(H;p)\hookrightarrow\mathcal{K}(H)$, $q\mapsto q-p$. The finite restricted Grassmanian $\operatorname{Gr}_{f}(H;p)$ is defined in a similar manner; it is a subspace of $\operatorname{Gr}_{K}(H;p)$. The restricted Grassmanian depends only on the corresponding equivalence class of the projection $p$.
Subbundles of a Hilbert bundle.   We will consider only locally trivial Hilbert bundles with a separable fiber $H$. However, we want to deal with their subbundles which are not necessarily locally trivial. Such subbundles arise, for example, as spectral sections of graph continuous families of Fredholm operators.
In general, a subbundle $\mathcal{E}$ of $\mathcal{H}$ is given by a family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ which is continuous in an appropriate sense. We will use the following notion of continuity: we say that $\mathcal{E}$ is a subbundle of $\mathcal{H}$ if $\mathcal{E}$ is the range of a bundle projection, that is, a continuous bundle map $p\colon\mathcal{H}\to\mathcal{H}$ such that $p_{x}\in\mathcal{P}(\mathcal{H}_{x})$ for every $x\in X$. In this case, $\mathcal{H}\ominus\mathcal{E}=\operatorname{Ran}(1-p)$ is also a subbundle. If both $\mathcal{E}$ and $\mathcal{H}\ominus\mathcal{E}$ are subbundles of $\mathcal{H}$ in the sense of [DD], then they are subbundles in our sense. See [DD], Proposition 16.
Equivalently, for every $x\in X$, some (and then every) local trivialization of $\mathcal{H}$ over a neighborhood $X^{\prime}$ of $x$ takes $p$ to a continuous map $X^{\prime}\to{{}^{s}{\mathcal{P}}}(H)$, where ${{}^{s}{\mathcal{P}}}(H)$ denotes the space $\mathcal{P}(H)$ equipped with the strong operator topology. In particular, subbundles of the trivial Hilbert bundle $H_{X}=X\times H$ are in one-to-one correspondence with continuous maps $X\to{{}^{s}{\mathcal{P}}}(H)$.
Two subbundles of $\mathcal{H}$ given by bundle projections $p$ and $q$ are said to be (compactly) compatible if $p-q$ is a norm continuous family of compact operators. They are said to be finitely compatible if additionally all operators $p_{x}-q_{x}$ have finite rank.
Polarizations of a Hilbert bundle.   We need only compact polarizations in this paper, so we will mostly use the term “polarization” without the adjective “compact”. The changes required to obtain the finite version are obvious.
Locally, a polarization of $\mathcal{H}$ is determined by a decomposition of $\mathcal{H}$, over some open subset $X^{\prime}\subset X$, into the orthogonal sum $\mathcal{H}=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$ of two subbundles of infinite rank. Equivalently, it is determined by the subbundle $\mathcal{H}^{+}$ of $\left.\mathcal{H}\right|_{X^{\prime}}$, or by the bundle projection with the range $\mathcal{H}^{+}$. By definition, two compatible subbundles of $\mathcal{H}$ over $X^{\prime}$ determine the same polarization.
Globally, a polarization of $\mathcal{H}$ is determined by an open covering $(X_{i})$ of $X$ and subbundles $\mathcal{H}^{+}_{i}$ of $\left.\mathcal{H}\right|_{X_{i}}$ such that $\mathcal{H}^{+}_{i}$ has infinite rank and corank and is compatible with $\mathcal{H}^{+}_{j}$ over $X_{i}\cap X_{j}$ for every $i,j$. Following Ivanov [$I_{2}$], we call such a collection $(X_{i},\mathcal{H}^{+}_{i})$ a polarization atlas. Two polarization atlases are said to be equivalent if their union is also a polarization atlas. A polarization of $\mathcal{H}$ is an equivalence class of polarization atlases.
A polarization $\Pi$ of $\mathcal{H}$ determines the restricted Grassmanian bundle $\operatorname{Gr}_{K}(\mathcal{H};\Pi)$ in an obvious way. This is one of the motivations behind the notion of a polarization.
Two polarizations of $\mathcal{H}$ are said to be homotopic if they are restrictions to $X\times\left\{0\right\}$ and $X\times\left\{1\right\}$ of some polarization of the Hilbert bundle ×$\mathcal{H}\times[0,1]\to X\times[0,1]$.
Adapted trivializations.   Following [$I_{2}$], we say that a trivialization $\mathcal{H}\to X\times H$ is adapted to a subbundle $\mathcal{E}$ of $\mathcal{H}$ if it takes the family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ to a continuous map $X\to\operatorname{Gr}(H)$. If a trivialization of $\mathcal{H}$ is adapted to $\mathcal{E}$, then it is adapted to every subbundle compatible with $\mathcal{E}$.
Let $\Pi$ be a polarization given by an atlas $(X_{i},\mathcal{E}_{i})$. We say that a trivialization of $\mathcal{H}$ is adapted to $\Pi$ if its restrictions to $X_{i}$ are adapted to $\mathcal{E}_{i}$. The notion of an adapted local trivialization is defined in an obvious way. We call a polarization $\Pi$ of $\mathcal{H}$ tame if it admits local adapted trivializations (that is, for every point $x\in X$ there is a local trivialization of $\mathcal{H}$ adapted to $\Pi$ over some neighborhood of $x$).
The notion of a polarization introduced by Ivanov in the appendix to the third version of his preprint [$I_{2}$] is a particular case of a tame finite polarization in our terminology.
Theorem 4.1.
Let $\Pi$ be a tame polarization of a locally trivial Hilbert bundle $\mathcal{H}$ over a paracompact space $X$. Then $\Pi$ admits a global adapted trivialization. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization. In particular, every two adapted trivializations are homotopic (as trivializations adapted to $\Pi$).
Proof. Let us describe the plan of the proof. For an arbitrary polarization $\Pi$ we define a fiber bundle $\mathcal{T}=\mathcal{T}_{\Pi}$ over $X$ whose (local or global) sections are exactly (local or global) trivializations of $\mathcal{H}$ adapted to $\Pi$. In particular, $\mathcal{T}$ admits local sections if and only if $\Pi$ is tame. We will show that in this case the bundle $\mathcal{T}$ is locally trivial and has a contractible fiber. Such a bundle, over a paracompact base space $X$, admits global sections, and every section over a closed subset $Y\subset X$ can be extended to a global section [DD, Lemma 4]. Therefore, $\Pi$ admits global adapted trivializations, and every trivialization of $\mathcal{H}$ over $Y$ adapted to $\Pi$ can be extended to a global adapted trivialization.
Let $H_{X}$ be the trivial Hilbert bundle over $X$ with the fiber $H$. Consider the fiber bundle $\mathcal{U}(\mathcal{H},H_{X})$ associated with $\mathcal{H}$ whose sections are Hilbert bundle isomorphisms $\mathcal{H}\to H_{X}$ (that is, trivializations of $\mathcal{H})$. Every two such isomorphisms differ by an automorphism of $H_{X}$, which is defined by a continuous map $X\to{{}^{s}{\mathcal{U}}}(H)$. Therefore, $\mathcal{U}(\mathcal{H},H_{X})$ is a locally trivial principal bundle over $X$ with the structure group ${{}^{s}{\mathcal{U}}}(H)$. Its fiber over $x$ is ${{}^{s}{\mathcal{U}}}(\mathcal{H}_{x},H)$, and ${{}^{s}{\mathcal{U}}}(H)$ acts on the fiber by the left multiplication. A trivialization $\tau\colon\mathcal{H}\to H_{X}$ takes $\mathcal{U}(\mathcal{H},H_{X})$ to the trivial bundle $\mathcal{U}(H_{X},H_{X})=X\times{{}^{s}{\mathcal{U}}}(H)$.
The key ingredient of our proof is the bundle $\mathcal{T}\to X$ which we will now define. The total space $\mathcal{T}$ coincides with $\mathcal{U}(\mathcal{H},H_{X})$ as a set, and the projection to $X$ is the same, but we equip $\mathcal{T}$ with a finer topology determined by $\Pi$. Hence the identity map
(4.1) $\iota\colon\mathcal{T}\to\mathcal{U}(\mathcal{H},H_{X})$
will be continuous and, in particular, the projection $\mathcal{T}\to X$ will be continuous.
Suppose first that the polarization $\Pi$ is defined by a subbundle $\mathcal{E}$ of $\mathcal{H}$, that is, the atlas of $\Pi$ has only one chart $(X,\mathcal{E})$. Then $\mathcal{E}$ determines the map
(4.2) $\iota_{\mathcal{E}}\colon\mathcal{T}\to\operatorname{Gr}(H)\cong\mathcal{P}(H),\quad(x,u)\mapsto u(\mathcal{E}_{x})\;\;\text{for}\;\;x\in X\;\;\text{and}\;\;u\in\mathcal{U}(\mathcal{H}_{x},H).$
We equip $\mathcal{T}=\mathcal{T}_{\mathcal{E}}$ with the weakest topology for which both maps (4.1) and (4.2) are continuous.
If $\mathcal{E}$ is replaced by a compatible subbundle $\mathcal{F}$, then $\mathcal{T}_{\mathcal{F}}=\mathcal{T}_{\mathcal{E}}$. Indeed, in this case the difference $k_{x}=p_{x}-q_{x}$ of projections onto $\mathcal{F}_{x}$ and $\mathcal{E}_{x}$ is a norm continuous family of compact operators, so $\iota_{\mathcal{F}}$ differs from $\iota_{\mathcal{E}}$ by a continuous map
$\mathcal{U}(\mathcal{H},H_{X})\to\mathcal{K}(H),\quad(x,u)\mapsto uk_{x}u^{*}.$
Therefore, $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ induce the same refinement of the original topology on $\mathcal{U}(\mathcal{H},H_{X})$ and thus $\mathcal{T}_{\mathcal{E}}$ depends only on the equivalence class of compatible subbundles $\mathcal{E}$. This allows to define the bundle $\mathcal{T}_{\Pi}$ for an arbitrary polarization $\Pi$ of $\mathcal{H}$.
More precisely, let $\Pi$ be given by an atlas $(X_{i},\mathcal{E}_{i})$, $i\in I$. Then the restrictions of $\mathcal{T}_{\mathcal{E}_{i}}$ and $\mathcal{T}_{\mathcal{E}_{j}}$ to $X_{i}\cap X_{j}$ coincide. We equip $\mathcal{T}_{\Pi}$ with the topology defined by the requirement that its restriction to $X_{i}$ coincides with $\mathcal{T}_{\mathcal{E}_{i}}$ for every $i\in I$.
By the definition of $\mathcal{T}_{\Pi}$, its (local) sections are exactly (local) trivializations of $\mathcal{H}$ adapted to $\Pi$. In particular, $\mathcal{T}_{\Pi}$ admits local sections if $\Pi$ is tame. The following lemma provides a stronger result.
Lemma 4.2.
If $\Pi$ is tame, then the bundle $\mathcal{T}_{\Pi}$ is locally trivial.
Proof. Since $\Pi$ is tame, each point $x_{0}\in X$ has a neighborhood $X^{\prime}$ such that some trivialization of $\mathcal{H}$ over $X^{\prime}$ is adapted to $\Pi$. We can suppose that $X^{\prime}$ is contained in some $X_{i}$; then this trivialization takes $\mathcal{E}_{i}$ to the subbundle $\mathcal{E}^{\prime}$ given by a continuous map $\mathcal{E}^{\prime}\colon X^{\prime}\to\operatorname{Gr}^{\star}(H)$. Let $g\colon\operatorname{Gr}^{\star}\to\mathcal{U}(H)$ be a map given by (2.1) and $\bar{g}_{L}=g_{L}^{-1}$. Then the composition of $\mathcal{E}^{\prime}$ with $\bar{g}\colon\operatorname{Gr}^{\star}\to\mathcal{U}(H)$ provides the automorphism $G$ of the trivial Hilbert bundle $H_{X^{\prime}}$ which takes $\mathcal{E}^{\prime}_{x}$ to the subspace $E$ of $H$ independent of $x$. We can then replace the original adapted trivialization by its composition with $G$. The new trivialization takes $\mathcal{E}_{i}$ to the subbundle $E_{X^{\prime}}$ of $H_{X^{\prime}}$ and thus provides a bundle isomorphism $\left.\mathcal{T}\right|_{X^{\prime}}\to\mathcal{T}^{\prime}$, where the bundle $\mathcal{T}^{\prime}$ is build (in the way described above) from the trivial bundle $H_{X^{\prime}}$ and its subbundle $E_{X^{\prime}}$. We will prove that the bundle $\mathcal{T}^{\prime}$ is trivial and find its fiber.
By the definition, the total space $\mathcal{T}^{\prime}$ is obtained from the product $X^{\prime}\times{{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that the map $\mathcal{T}^{\prime}\to\operatorname{Gr}^{\star}(H)$, $(x,u)\mapsto uE$ becomes continuous. This map does not actually depend on $x$. Therefore, $\mathcal{T}^{\prime}$ coincides with the trivial bundle $X^{\prime}\times T\to X^{\prime}$, where the fiber $T$ is obtained from ${{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that
(4.3) $\iota_{E}\colon T\to\operatorname{Gr}^{\star}(H),\quad u\mapsto uE$
becomes continuous. In other words, $T$ has the topology induced by the inclusion
$T\hookrightarrow{{}^{s}{\mathcal{U}}}(H)\times\operatorname{Gr}^{\star}(H),\quad u\mapsto(u,uE).$
Hence the restriction of $\mathcal{T}_{\Pi}$ to a neighborhood of an arbitrary point $x_{0}$ is a trivial bundle with the fiber $T$. $\square$
Lemma 4.3.
The space $T$ defined above is contractible.
Proof. The fiber of $\iota_{E}$ over $E$ is the subspace of ${{}^{s}{\mathcal{U}}}(H)$ consisting of unitary operators preserving $E$. Denote this subspace by ${{}^{s}{\mathcal{U}}}_{E}$. We will show that (4.3) is a trivial bundle over $\operatorname{Gr}^{\star}$ and construct its trivialization $\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}\to T$.
Let $g\colon\operatorname{Gr}^{\star}(H)\to\mathcal{U}(H)$ be a continuous map such that $L=g_{L}E$ for all $L\in\operatorname{Gr}^{\star}(H)$, as in (2.1). We define $\varphi$ by the formula
$\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}(H)\;\to\;T,\quad\varphi(v,L)=g_{L}\cdot v.$
Since
(4.4) $(g_{L}v)E=g_{L}E=L\;\;\text{for every}\;\;v\in{{}^{s}{\mathcal{U}}}_{E},$
the composition $\iota_{E}\circ\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}\to\operatorname{Gr}^{\star}$ is the projection of the product on the second factor; in particular, it is continuous. The composition of $\varphi$ with the identity map $T\to{{}^{s}{\mathcal{U}}}$ is given by $(v,L)\mapsto g_{L}v$ and is also continuous. Therefore, $\varphi$ itself is continuous.
Equality (4.4) allows to recover $L$ and $v$ from $u=g_{L}v$ by taking $L=uE$ and $v=g_{L}^{-1}u$. Since $vE=g_{L}^{-1}uE=g_{L}^{-1}L=E$, the unitary operator $v=g_{L}^{-1}u$ belongs to ${{}^{s}{\mathcal{U}}}_{E}$, and we obtain the continuous map
$\psi\colon T\to{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}(H),\quad\psi(u)=\left(g_{uE}^{-1}\cdot u,\,uE\right).$
By the construction, $\psi\circ\varphi(v,L)=(v,L)$ for every $(v,L)\in{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}$. Conversely,
$\varphi\circ\psi(u)=\varphi\left(g_{uE}^{-1}\cdot u,\,uE\right)=g_{uE}\cdot g_{uE}^{-1}\cdot u=u\;\;\text{for every}\;\;u\in T.$
Therefore, $\psi$ is inverse to $\varphi$ and thus $\varphi$ is a homeomorphism.
The domain of $\varphi$ is the product of two factors. The second factor $\operatorname{Gr}^{\star}$ is contractible, see [AS, proof of Lemma 3.6]. The first factor ${{}^{s}{\mathcal{U}}}_{E}={{}^{s}{\mathcal{U}}}(E)\times{{}^{s}{\mathcal{U}}}(E^{\bot})$ is the product of two copies of ${{}^{s}{\mathcal{U}}}$. By [DD, Lemma 3], ${{}^{s}{\mathcal{U}}}$ is contractible, so ${{}^{s}{\mathcal{U}}}_{E}$ is also contractible. It follows that the product ${{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}$ is contractible. This completes the proof of the lemma. $\square$
End of the proof of Theorem 4.1.   Lemmas 4.2 and 4.3 together with the reasoning in the beginning of the proof show that every adapted trivialization of $\mathcal{H}$ over a closed subspace $Y\subset X$ can be extended to a global adapted trivialization. If $\tau_{0}$ and $\tau_{1}$ are trivializations of $\mathcal{H}$ adapted to $\Pi$, then $\tau_{0}\sqcup\tau_{1}$ is a trivialization of $\tilde{\mathcal{H}}$ over $X\times\left\{0,1\right\}$ adapted to $\tilde{\Pi}$, where $\tilde{\mathcal{H}}$ and $\tilde{\Pi}$ are the liftings of $\mathcal{H}$ and $\Pi$ to $X\times[0,1]$. The polarization $\tilde{\Pi}$ is tame, so $\tau_{0}\sqcup\tau_{1}$ can be extended from $X\times\left\{0,1\right\}$ to a trivialization of $\tilde{\mathcal{H}}$ adapted to $\tilde{\Pi}$. This provides a homotopy between $\tau_{0}$ and $\tau_{1}$ in the class of trivializations adapted to $\Pi$ and completes the proof of the theorem. $\square$
Finite covering dimension.   The proof of Theorem E in the next section is based on the following result.
Theorem 4.4.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a metrizable space $X$ of locally finite covering dimension. Then every polarization $\Pi$ of $\mathcal{H}$ is tame and thus satisfies all the conclusions of Theorem 4.1.
Proof. Every subspace $X^{\prime}$ of $X$ is metrizable and thus paracompact, and has locally finite covering dimension by [E, Theorem 7.3.1]. By [DD, Theorem 5], every separable Hilbert bundle over $X^{\prime}$ of infinite rank is locally trivial. The bundle $\mathcal{H}$ is trivial by [DD, Theorem 1] and has separable fiber, so it is separable in the sense of [DD], and so are its subbundles over $X^{\prime}$. Therefore, every infinite rank subbundle of $\mathcal{H}$ over $X^{\prime}$ is locally trivial.
Let $(X_{i},\mathcal{E}_{i})$ be an atlas of $\Pi$. By the reasoning above, both $\mathcal{E}_{i}$ and $\mathcal{E}_{i}^{\bot}$ are locally trivial. The direct sum of their local trivializations provides a local trivialization of $\mathcal{H}$ which takes $\mathcal{E}_{i}$ to a constant subspace and thus is adapted to $\mathcal{E}_{i}$. Therefore, $\Pi$ is tame. Since $X$ is paracompact, $\Pi$ satisfies the assumptions of Theorem 4.1. $\square$
Remark.
For a general space $X$, such a result is no longer true. Indeed, Dixmier and Douady constructed a subbundle $\mathcal{E}$ of the trivial Hilbert bundle $\mathcal{H}$ over a compact space $X$, which is not locally trivial [DD, Theorem 6]. We recall their construction in Example 5.5 below. Every fiber of $\mathcal{E}$ has infinite rank and corank, so $\mathcal{E}$ determines a polarization of $\mathcal{H}$. However, $\mathcal{E}$ has no adapted trivializations, even locally.
5 Self-adjoint operators
Adapted trivializations.   We say that a (local or global) trivialization of $\mathcal{H}$ is adapted to a family $\mathcal{A}$ of regular operators $\mathcal{A}_{x}\in\mathcal{R}(\mathcal{H}_{x})$ if it takes $\mathcal{A}$ to a Riesz continuous family of operators in $H$. If operators $\mathcal{A}_{x}$ are bounded (say, they are projections), then an adapted trivialization takes $\mathcal{A}$ to a norm continuous family. So the definition from the previous section, of a trivialization adapted to a subbundle, agrees with the present definition.
Bounded operators.   Let $\mathcal{A}$ be a family of self-adjoint bounded operators acting on fibers of a Hilbert bundle $\mathcal{H}$. We want to know whether $\mathcal{A}$ admits a global adapted trivialization. Clearly, for that to be true, $\mathcal{A}$ has to be norm quasi-continuous (that is, admit local adapted trivializations). The following theorem shows that for essentially unitary operators this local condition is also sufficient. The author does not know whether this is true for arbitrary bounded self-adjoint operators.
Theorem 5.1.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$ and $\mathcal{A}$ be a family of self-adjoint essentially unitary operators. Suppose that $\mathcal{A}$ is norm quasi-continuous. Then there exists a global trivialization of $\mathcal{H}$ adapted to $\mathcal{A}$. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization. In particular, every two such trivializations are homotopic (in the space of trivializations adapted to $\mathcal{A}$).
Proof. Let $X^{+}$, resp. $X^{-}$ be the subset of points $x\in X$ for which $\mathcal{A}_{x}$ is essentially positive, resp. negative, and let $X^{\star}=X\setminus(X^{+}\cup X^{-})$. Since $\mathcal{A}$ is norm quasi-continuous, each of the three subspaces $X^{+}$, $X^{-}$, and $X^{\star}$ is open and thus closed in $X$, so it is enough to prove the theorem for each of these subspaces separately.
For $x\in X^{+}$, the operator $\mathcal{A}_{x}$ is essentially positive and essentially unitary, so $\mathcal{A}_{x}-1$ is compact. Over $X^{+}$ the family $x\mapsto\mathcal{A}_{x}-1$ is a norm quasi-continuous family of compact operators, so both $\mathcal{A}-1$ and $\mathcal{A}$ are norm continuous in every trivialization of $\mathcal{H}$ over $X^{+}$. The similar reasoning shows that $\mathcal{A}$ is norm continuous in every trivialization of $\mathcal{H}$ over $X^{-}$. Trivializations of $\mathcal{H}$ over $X^{+}\cup X^{-}$ are identified with maps $X^{+}\cup X^{-}\to{{}^{s}{\mathcal{U}}}(H)$, so every two trivializations are homotopic.
It remains to prove the theorem over $X^{\star}$. Without loss of generality we can suppose that $X^{\star}=X$. Consider the family of spectral projections
(5.1) $p_{\lambda}(x)=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})\;\;\text{for}\;\;x\in X_{\lambda}=\left\{x\in X\,|\,\lambda\in\operatorname{Res}(\mathcal{A}_{x})\right\}.$
The sets $X_{\lambda}$ with $\lambda$ running the interval $(-1,1)$ form an open covering of $X$. A local trivialization of $\mathcal{H}$ adapted to $\mathcal{A}$ takes $p_{\lambda}$ to a norm continuous family over $X_{\lambda}$. For $\lambda<\mu$ the difference
(5.2) $p_{\lambda}(x)-p_{\mu}(x)=\operatorname{\mathbb{1}}_{[\lambda,\mu)}(\mathcal{A}_{x})=\operatorname{\mathbb{1}}_{[\lambda,\mu]}(\mathcal{A}_{x})$
is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$. Hence the collection $(X_{\lambda},p_{\lambda})$, $\lambda\in(-1,1)$ is an atlas of a tame polarization $\Pi_{\mathcal{A}}$ of $\mathcal{H}$ over $X$. The difference $k_{\lambda}=\mathcal{A}-(2p_{\lambda}-1)$ is a norm continuous family of compact operators over $X_{\lambda}$ in a local trivialization adapted to $\mathcal{A}$, so it is norm continuous in every trivialization of $\mathcal{H}$ over $X_{\lambda}$. A trivialization of $\mathcal{H}$ over $X_{\lambda}$ is adapted to $\mathcal{A}=k_{\lambda}+(2p_{\lambda}-1)$ if and only if it is adapted to $p_{\lambda}$. Hence a trivialization is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$. It remains to apply Theorem 4.1. $\square$
Remark.
In fact, the main part of Theorem 5.1, about operators which are neither essentially positive nor essentially negative, is equivalent to Theorem 4.1. Indeed, let $(X_{i},p_{i})$ be an atlas of a polarization $\Pi$. We associate with it the family $\alpha$ of self-adjoint essentially unitary operators $\alpha(x)=\sum_{i}\rho_{i}(x)r_{i}(x)$, where $(\rho_{i})$ is a partition of unity subordinated to the covering $(X_{i})$ and $r_{i}=2p_{i}-1$. Then $\alpha-r_{i}=\sum_{j}\rho_{j}(r_{j}-r_{i})$ is a norm continuous family of compact operators over $X_{i}$. Therefore, a (local or global) trivialization of $\mathcal{H}$ is adapted to $\Pi$ if and only if it is adapted to $\alpha$.
Proof of Theorem D.   The bounded transform takes self-adjoint operators with compact resolvents to self-adjoint essentially unitary operators and Riesz (quasi-)continuous families to norm (quasi-)continuous families. Therefore, Theorem D follows immediately from Theorem 5.1.
Spectral projections.   For a family $\mathcal{A}_{x}$, $x\in X$ of self-adjoint operators and $\lambda\in\mathbb{R}$ we consider the family $p_{\lambda}$ of spectral projections over $X_{\lambda}\subset X$ defined by (5.1).
Proposition 5.2.
If $\mathcal{A}\colon X\to\mathcal{R}^{\mathrm{sa}}(H)$ is graph continuous, then $p_{\lambda}\colon X_{\lambda}\to\mathcal{P}(H)$ is strong continuous for every $\lambda\in\mathbb{R}$.
Proof. Without loss of generality we can suppose that $X={{}^{g}\mathcal{R}}^{\mathrm{sa}}(H)$ and $\mathcal{A}$ is the identity map. In other words, we need to prove that the spectral projection
$p_{\lambda}=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}\colon\mathcal{R}^{\mathrm{sa}}_{\lambda}\to\mathcal{P}(H),\quad\mathcal{R}^{\mathrm{sa}}_{\lambda}=\left\{A\in\mathcal{R}^{\mathrm{sa}}(H)\,|\,\lambda\in\operatorname{Res}(A)\right\}$
[…]
Proposition 5.3.
Let $\mathcal{A}\colon X\to\mathcal{R}^{\mathrm{sa}}(H)$ be graph continuous and $\lambda\in\mathbb{R}$. Then $\mathcal{A}$ is Riesz continuous on $X_{\lambda}$ if and only if $p_{\lambda}\colon X_{\lambda}\to\mathcal{P}(H)$ is norm continuous.
Proof. Without loss of generality we can suppose that $X_{\lambda}=X$, that is, $\lambda$ belongs to the resolvent set of each $\mathcal{A}_{x}$, $x\in X$.
Suppose that $\mathcal{A}$ is Riesz continuous. Then the family $a_{x}=\bm{\chi}(\mathcal{A}_{x})$ is norm continuous and $\mu=\bm{\chi}(\lambda)$ lies in the resolvent set of each $a_{x}$. Therefore, the map $x\mapsto p_{\lambda}(x)=\operatorname{\mathbb{1}}_{[\mu,+\infty)}(a_{x})$ is also norm continuous.
Conversely, suppose that $p_{\lambda}$ is norm continuous. Then $r_{\lambda}=2p_{\lambda}-1$ is a norm continuous family of symmetries commuting with $\mathcal{A}$. The family $\mathcal{A}^{\prime}=\mathcal{A}\cdot r_{\lambda}$ of self-adjoint operators is graph (resp Riesz) continuous if and only if $\mathcal{A}$ is graph (resp Riesz) continuous. But, in contrast with $\mathcal{A}$, the family $\mathcal{A}^{\prime}$ is uniformly bounded from below by $-|\lambda|$. The graph and Riesz topology coincide on the set of operators bounded from below by a given constant [$P_{2}$, Proposition 11.3], so $\mathcal{A}^{\prime}$ and thus also $\mathcal{A}$ are Riesz continuous. $\square$
Polarization associated with operator family.   Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators with compact resolvents acting on fibers of a Hilbert bundle $\mathcal{H}$. Then $X_{\lambda}$ is open for every $\lambda\in\mathbb{R}$ by [Ka, Theorem IV.3.1] and these sets form an open covering of $X$. By Proposition 5.2, $p_{\lambda}$ is strongly continuous in local trivializations of $\mathcal{H}$, so it determines a decomposition of $\mathcal{H}$ over $X_{\lambda}$ into the orthogonal sum $\mathcal{H}_{\lambda}^{+}\oplus\mathcal{H}_{\lambda}^{-}$ of two Hilbert subbundles. For $\lambda<\mu$ the difference (5.2) is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$ by [Ka, Theorem IV.3.16].
Suppose additionally that each $\mathcal{A}_{x}$ is neither essentially positive nor essentially negative. Then $p_{\lambda}(x)\in\mathcal{P}^{\star}(\mathcal{H}_{x})$, so the collection $(X_{\lambda},p_{\lambda})$ is an atlas of a polarization. Hence a family $\mathcal{A}$ satisfying assumptions above determines a polarization of $\mathcal{H}$ which we call the polarization associated with $\mathcal{A}$ and denote by $\Pi_{\mathcal{A}}$. (The associated polarization $\Pi_{\mathcal{A}}$ is actually finite, but we are interested only in its equivalence class with respect to compact deformations, so we will consider it as a compact polarization.)
Proposition 5.4.
Let $\mathcal{A}$ be a graph continuous family of operators $\mathcal{A}_{x}\in\mathcal{R}^{\star}_{K}(\mathcal{H}_{x})$ acting on fibers of a Hilbert bundle $\mathcal{H}$. A trivialization of $\mathcal{H}$ is adapted to $\mathcal{A}$ if and only if it is adapted to the associated polarization $\Pi_{\mathcal{A}}$.
Proof. This follows immediately from Proposition 5.3. $\square$
Remark.
More generally, consider a family $\mathcal{A}$ of Fredholm self-adjoint operators acting on fibers of $\mathcal{H}$. Suppose that $\mathcal{A}$ is graph quasi-continuous (that is, $\mathcal{A}$ admits local trivializations taking it to a graph continuous family). Then the associated polarization $\Pi_{\mathcal{A}}$ is defined in a similar manner. The only change is in the definition of the set $X_{\lambda}$: one should add the condition that the interval $[\lambda,0]$ or $[0,\lambda]$, depending on the sign of $\lambda$, is disjoint from the essential spectrum of $\mathcal{A}_{x}$.
If a trivialization is adapted to such a family $\mathcal{A}$, then it is adapted to $\Pi_{\mathcal{A}}$ as well. However, the second direction of Proposition 5.4 is no longer true: a trivialization adapted to $\Pi_{\mathcal{A}}$ is not necessarily adapted to $\mathcal{A}$, because it might turn $\mathcal{A}$ into a graph discontinuous family and then Proposition 5.3 cannot be applied. This is the place where condition of compact resolvents comes into the picture: the graph continuity of a family of operators with compact resolvents is independent of the choice of a trivialization.
Proof of Theorem E.   Let $\Pi_{\mathcal{A}}$ be the polarization of $\mathcal{H}$ associated with $\mathcal{A}$. By Proposition 5.4, a (local or global) trivialization of $\mathcal{H}$ is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$. It remains to apply Theorem 4.4. $\square$
Example 5.5.
Let us show that Theorem E does not hold without additional assumptions on the base space $X$. We will construct a graph continuous family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ of invertible self-adjoint operators with compact resolvents over a compact metrizable space $X$ such that $\mathcal{A}$ is not Riesz quasi-continuous. Moreover, for every point $x\in X$, $\mathcal{A}$ is not Riesz quasi-continuous at $x$ (that is, it is not Riesz quasi-continuous in any neighborhood of $x$).
Our example is based on a construction of a non-trivial separable Hilbert bundle $\mathcal{E}$ over a compact space $X$ by Dixmier and Douady. See [DD, Theorem 6]. Let us recall their construction. Let $H_{i}=\mathbb{C}^{2}$, $X_{i}=\mathbb{P}(H_{i})\cong\mathbb{C}\mathbb{P}^{1}$, $H=\oplus_{i=1}^{\infty}H_{i}$, and $X=\prod_{i=1}^{\infty}X_{i}$ with the product topology. The space $X$ is a countable product of metrizable spaces and thus is metrizable itself; it is a product of compact spaces and thus is compact itself. The Hilbert bundle $\mathcal{E}$ over $X$ is defined as the subbundle of the trivial Hilbert bundle $H_{X}$ with the fiber $\mathcal{E}_{x}=\oplus_{i=1}^{\infty}x_{i}$ over $x=(x_{i})$, where $x_{i}\in X_{i}$ is considered as a one-dimensional subspace of $H_{i}$. By [DD, Theorem 6], every section of $\mathcal{E}$ vanishes at some point. In particular, $\mathcal{E}$ is not trivial and thus it is not locally trivial (since all locally trivial Hilbert bundles over paracompact base spaces are trivial).
Moreover, the bundle $\mathcal{E}$ looks the same near each point of $X$ (more precisely, the group $\prod_{i=1}^{\infty}\mathcal{U}(H_{i})$ acts transitively on $X$ and this action lifts to the action on $\mathcal{E}$ by bundle automorphisms), so a local triviality of $\mathcal{E}$ near one point would imply local triviality over the whole $X$. Therefore, for every open subset $X^{\prime}\subset X$, $\mathcal{E}$ is not trivial over $X^{\prime}$.
Let $c\colon\mathbb{N}\to\mathbb{R}_{>0}$ be a sequence of positive numbers converging to $\infty$. Then
$A=\oplus_{i=1}^{\infty}c_{i}\cdot\operatorname{Id}_{H_{i}}$
is a positive regular operator with compact resolvents. Let $p(x)=\oplus_{i=1}^{\infty}p_{i}(x)$ be the orthogonal projection of $H$ onto $\mathcal{E}_{x}$, and let $r(x)=2p(x)-1$ be the corresponding symmetry. Then $r(x)$ commutes with $A$, and $\mathcal{A}(x)=A\cdot r(x)$ is an invertible self-adjoint operator with compact resolvents, which is neither essentially positive nor essentially negative.
We claim that the family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ is graph continuous. Indeed, $(1+\mathcal{A}_{x}^{2})^{-1}=(1+A^{2})^{-1}$ is a compact operator independent of $x$. The map
$x\longmapsto\mathcal{A}_{x}(1+\mathcal{A}_{x}^{2})^{-1}=r_{x}A(1+A^{2})^{-1}$
is the product of a strongly continuous family of symmetries $r_{x}$ and a fixed compact operator $A(1+A^{2})^{-1}=\bm{\chi}(A)(1+A^{2})^{-1/2}$ and thus is norm continuous. Therefore, all components of the orthogonal projection (2.6) onto the graph of $\mathcal{A}_{x}$ are norm continuous and thus $\mathcal{A}$ is graph continuous.
However, for every open subset $X^{\prime}\subset X$, the restriction of $\mathcal{A}$ to $X^{\prime}$ is not Riesz quasi-continuous. Indeed, the positive spectral projection $p(x)$ of $\mathcal{A}(x)$ is the orthogonal projection of $H$ onto $\mathcal{E}_{x}$. If $\mathcal{A}$ were Riesz continuous in some local trivialization of the trivial Hilbert bundle $X^{\prime}\times H$, then the map $x\mapsto p(x)$ would be norm continuous in the same local trivialization and $\mathcal{E}$ would be a locally trivial Hilbert bundle over $X^{\prime}$, which is not the case.
Remark.
If one multiplies the family $\mathcal{A}_{x}$ from this example by the strongly continuous family of unitary operators $r_{x}$, then the resulting family $x\mapsto\mathcal{A}_{x}\cdot r_{x}=A$ will be Riesz continuous (and even constant). However, for every strongly continuous family of unitary operators $v_{x}\in\mathcal{U}(H)$, the family $x\mapsto v_{x}\mathcal{A}_{x}v_{x}^{*}$ is Riesz discontinuous at every point $x\in X$. This demonstrates the difference between Theorems C and D, that is, between one-sided multiplication and conjugation by unitary operators, which we discussed in the Introduction.

### v3

@@
 Abstract
-We show that every graph continuous family of unbounded operators in a Hilbert space becomes Riesz continuous after multiplication by an appropriate family of unitary operators.
-This result provides a simple definition of the index for graph continuous families of Fredholm operators and leads to two corollaries for operators with compact resolvents: (1) the identity map between the spaces of such operators equipped with the Riesz and the graph topology is a homotopy equivalence; (2) every graph continuous family of such operators acting between fibers of Hilbert bundles becomes Riesz continuous in appropriate trivializations of the bundles.
-For self-adjoint operators, multiplication by unitaries should be replaced by conjugation.
+We show that every graph continuous family of unbounded operators in a Hilbert space becomes Riesz continuous after one-sided multiplication by an appropriate family of unitary operators.
+This result provides a simple definition of the index for graph continuous families of Fredholm operators, and we show that for such families this index coincides with the index defined by N.
+Ivanov in [I2].
+This result also has two corollaries for operators with compact resolvents: (1) the identity map between the space of such operators with the Riesz topology and the space of such operators with the graph topology is a homotopy equivalence; (2) every graph continuous family of such operators acting between fibers of Hilbert bundles becomes Riesz continuous in appropriate trivializations of the bundles.
+For self-adjoint operators, multiplication by unitary operators should be replaced by conjugation.
 In general, a graph continuous family of self-adjoint operators with compact resolvents cannot be made Riesz continuous by an appropriate conjugation.
-We obtain a partial analogue of the “trivialization” result above for self-adjoint operators and describe obstructions to existence of such a trivialization arising in the general case.
-This part of results is closely related to the recent work [$I_{2}$] of N.
-Ivanov.
-footnotetext: Department of Mathematics, Technion – Israel Institute of Technology.footnotetext: This work was partially supported by ISF grants no. 431/20 and 844/19.
+We obtain a partial analogue of the “trivialization” result above for self-adjoint operators and describe obstructions to existence of such a trivialization in the general case.
+This motivates the notion of a polarization of a Hilbert bundle, and we prove a similar result for polarizations.
+These results are closely related to the recent work [I2] of Ivanov and provide alternative proofs for some of his results.
+We then show that, under a minor assumption on the space of parameters and for operators which are neither essentially positive nor essentially negative, there is always a trivialization making the family Riesz continuous.
+footnotetext: Department of Mathematics, University of Haifa (Israel)
+Department of Mathematics, Technion – Israel Institute of Technology
 1 Introduction
@@
 The set $\mathcal{R}(H,H^{\prime})$ of regular operators has two important natural topologies.
-The Riesz topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
+The Riesz topology on $\mathcal{R}(H,H^{\prime})$ is induced by the so called bounded transform map
 $\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow\mathcal{B}(H,H^{\prime}),\quad\bm{\chi}(A)=A(1+A^{*}A)^{-1/2}$
-from the norm topology on the space of bounded operators. Here $\bm{\chi}$ is the so called bounded transform map. The graph topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
-$\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow\mathcal{P}(H\oplus H^{\prime})$
-from the norm topology on the space of projections, where $\mathbf{p}$ is the map taking a regular operator to the orthogonal projection onto its graph.
+from the norm topology on the space of bounded operators. The graph topology on $\mathcal{R}(H,H^{\prime})$ is induced by the inclusion
+$\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow\mathcal{P}(H\oplus H^{\prime}),$
+from the norm topology on the space $\mathcal{P}(H\oplus H^{\prime})$ of projections, where $\mathbf{p}$ is the map taking a regular operator to the orthogonal projection onto its graph.
 Let ${{}^{r}\mathcal{R}}(H,H^{\prime})$, resp.
@@
 graph topology.
-The Riesz topology is strictly stronger than the graph topology.
+The Riesz topology is strictly finer than the graph topology.
 In other words, the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is continuous, while the identity map ${{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ is not.
@@
 From graph to Riesz continuity.
-Riesz continuous families of operators are much easier to deal with than mere graph continuous families.
-This motivates the approach that we are developing in this paper.
-Suppose that one is given a graph continuous family of regular operators.
-One can try to modify it, in an appropriate way, so that the resulting family becomes Riesz continuous, then investigate the modified family and transfer results of this investigation to the original family.
-By an “appropriate way” we mean that such a transfer of results is possible, that is, some important properties of the original family may be recovered from the corresponding properties of the modified family.
-Of course, this description is quite vague; we will make it more precise below.
-Let us illustrate our approach by an example of a family consisting of invertible operators.
-Recall that an operator $A\in\mathcal{R}(H)$ is called invertible if it is bijective as a map $\operatorname{dom}(A)\to H$.
-Such operator has a unique polar decomposition $A=u\cdot|A|$, where $u\in\mathcal{U}(H)$ is a unitary operator and $|A|=\sqrt{A^{*}A}$ is a positive operator.
-If $\mathcal{A}\colon X\to\mathcal{R}(H)$ is graph continuous, then the map $x\mapsto|\mathcal{A}_{x}|$ is Riesz continuous.
-If additionally all operators $\mathcal{A}_{x}$ are invertible, then the map $x\mapsto u_{x}=\mathcal{A}_{x}|\mathcal{A}_{x}|^{-1}$ from $X$ to ${{}^{s}{\mathcal{U}}}(H)$, the unitary group equipped with the strong operator topology, is continuous.
-Taking $v_{x}=u_{x}^{*}$, we obtain a continuous map $v\colon X\to{{}^{s}{\mathcal{U}}}(H)$ such that the map $x\mapsto v_{x}\mathcal{A}_{x}=|\mathcal{A}_{x}|$ is Riesz continuous.
-The original family $\mathcal{A}$ can be recovered from the modified family $|\mathcal{A}|$ by the left multiplication by $u_{x}$.
-This construction cannot be extended continuously to general non-invertible families because the polar decomposition of a general non-invertible operator involves a partial isometry which is not unitary.
-Moreover, this partial isometry does not depend continuously on the operator.
-However, one can still modify a family to a Riesz continuous one using the multiplication by an appropriate family of unitary operators, as we show in Theorem A below (we will multiply on the right, but it really does not matter: one can always pass to adjoint operators, replacing the left multiplication with the right one).
-In contrast with the construction above, the modified family no longer consists of positive operators, nor is it determined by the original family in any canonical way.
-But the modified family still carries an important information about the original one.
-Theorem A.
-There exists a map $v\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H)$, which is both Riesz-to-norm and graph-to-strong continuous, such that the induced map
+Using the bounded transform, one can deal with Riesz continuous families of operators in essentially the same manner as with norm continuous families of bounded operators.
+Because of that, Riesz continuous families are much easier to deal with than mere graph continuous families.
+Our first result shows that one can turn a graph continuous family to a Riesz continuous one by a “change of coordinates”***The interpretation of our construction as a change of coordinates was suggested by the referee..
+Theorem 1.
+There exist a map $v\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H)$, which is both Riesz-to-norm and graph-to-strong continuous, such that the induced map
 $\Phi\colon\mathcal{R}(H,H^{\prime})\to\mathcal{R}(H,H^{\prime}),\quad A\mapsto A\cdot v(A)$
 is graph-to-Riesz continuous.
-[…]
+Passing to adjoint operators, one obtains a similar statement for the left multiplication.
+Namely, there is a map $v^{\prime}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{U}(H^{\prime})$, which is both Riesz-to-norm and graph-to-strong continuous, such that the induced map $A\mapsto v^{\prime}(A)\cdot A$ is graph-to-Riesz continuous.
+We will now present three applications of Theorem 1.
 $K^{0}$ index for graph continuous families.
-Let $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}$ be a graph continuous family and $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem A.
-Then $\bm{\chi}\circ\Phi\circ\mathcal{A}=\varphi\circ\mathcal{A}$ is a norm continuous family of bounded operators.
+Let $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}$ be a graph continuous map and $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem 1.
+Then the composition
+$\alpha=\bm{\chi}\circ\Phi\circ\mathcal{A}\colon X\to\mathcal{B}(H)$
+is norm continuous.
 Suppose that all $\mathcal{A}_{x}$ are Fredholm.
-Then the operators $\varphi(\mathcal{A}_{x})$ are also Fredholm, and one can define the index of $\mathcal{A}$ as the index of this norm continuous family of bounded Fredholm operators:
-$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\varphi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X).$
-This definition of the index is discussed in the last part of Section 3; it is justified by Theorem B below.
-Equivalently, the index of $\mathcal{A}$ may be defined as the index of the family of projections $\Gamma_{x}\to H^{\prime}$ after a trivialization of the Hilbert bundle $\Gamma$ formed by the graphs $\Gamma_{x}$ of operators $\mathcal{A}_{x}$.
-(This trivialization should be chosen in the category of bundles with the structure group $\mathcal{U}$, not ${{}^{s}{\mathcal{U}}}$.) Note that such a trivialization is not canonical; it is defined only up to a continuous map $X\to\mathcal{U}(H)$.
-However, change of a trivialization does not affect the homotopy type of the corresponding family of bounded Fredholm operators, since the unitary group $\mathcal{U}(H)$ is contractible.
-Searching for such a definition of the index was part of the motivation behind Theorem A.
-An alternative approach to the index of graph continuous families of Fredholm operators is provided by the previous paper [$P_{3}$] of the author.
+Then the operators $\alpha_{x}$ are also Fredholm, and one can define the index of $\mathcal{A}$ as the classical index of $\alpha\colon X\to\mathcal{B}_{F}$ (here $\mathcal{B}_{F}$ denotes the space of bounded Fredholm operators):
+$\operatorname{ind}\mathcal{A}=\operatorname{ind}\alpha=[\alpha]\in[X,\mathcal{B}_{F}]=K^{0}(X).$
+This definition of the index is justified by Theorem 2 below. We discuss the index of graph continuous families in the last part of Section 4.
+Searching for such a definition of the index was part of the motivation behind Theorem 1.
+An alternative approach to the index of graph continuous families of Fredholm operators is provided by the previous paper [P3] of the author.
 Another approach, for compact base spaces and in a more general framework of Hilbert modules over C*-algebras, is provided by results of M.
@@
 Recently N.
-Ivanov [$I_{1}$, $I_{2}$] developed a new approach to the index of families, which works under much weaker continuity assumptions than graph continuity.
-We show in the end of Section 3 that for graph continuous families of Fredholm operators our definition of the index given above coincides with Ivanov’s index.
+Ivanov developed a new approach to the index of families [I1, I2], which works under much weaker continuity assumptions than graph continuity.
+We show in the end of Section 4 that the definition of the index given above coincides with Ivanov’s index for graph continuous families of Fredholm operators.
 Homotopy equivalences.
-A regular operator $A$ is said to have compact resolvents***If the resolvent set $\operatorname{Res}(A)$ of $A$ is non-empty (for example, if $A$ is self-adjoint), then this definition agrees with the usual sense of the words “compact resolvent”: $(A-\lambda)^{-1}$ is compact for some (and then every) $\lambda\in\operatorname{Res}(A)$.
+A regular operator $A$ is said to have compact resolventsIf the resolvent set $\operatorname{Res}(A)$ of $A$ is non-empty (for example, if $A$ is self-adjoint), then this definition agrees with the usual sense of the words “compact resolvent”: $(A-\lambda)^{-1}$ is compact for some (and then every) $\lambda\in\operatorname{Res}(A)$.
 However, a regular operator may have an empty resolvent set; the definition above covers such operators as well.
@@
 are homotopy equivalences.
-See [$P_{3}$, Theorem A].
-However, the proof in [$P_{3}$] provides no intuition on how a homotopy inverse map might look like.
+See [P3, Theorem A].
+However, the proof in [P3] provides no intuition on how a homotopy inverse map might look like.
 The following theorem reveals a possible form of homotopy inverse maps and provides an alternative proof of homotopy equivalence of the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$.
-Theorem B.
-Let $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem A.
+Theorem 2.
+Let $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ be a map provided by Theorem 1.
 Then the restriction of $\Phi$ to the subspace of operators with compact resolvents is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$.
 Similarly, the restriction of $\Phi$ to the subspace of Fredholm operators is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$.
-Let $J$ be a subset of $\mathcal{K}(H)$ invariant under the conjugation by unitary operators. Then one can replace $\mathcal{R}_{K}$ in the statement of Theorem B with the set
-$\mathcal{R}_{J}(H)=\left\{A\in\mathcal{R}(H)\,|\,(1+A^{*}A)^{-1},(1+AA^{*})^{-1}\in J\right\}.$
-(In particular, $J$ can be the subset of trace class operators, or the subset of Hilbert–Schmidt operators.) Such a set $\mathcal{R}_{J}$ is invariant under the multiplication by unitary operators, so both $\Phi$ and homotopies used in the proof of Theorem B preserve $\mathcal{R}_{J}$.
-This provides the following generalization of the first part of Theorem B.
-Theorem B’.
-Let $J$ be a subset of $\mathcal{K}(H)$ invariant under the conjugation by unitary operators.
-Then the identity map ${{}^{r}\mathcal{R}}_{J}(H)\to{{}^{g}\mathcal{R}}_{J}(H)$ is a homotopy equivalence, with a homotopy inverse map given by $\Phi\colon{{}^{g}\mathcal{R}}_{J}\to{{}^{r}\mathcal{R}}_{J}$.
 Hilbert bundles.
 Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be locally trivial Hilbert bundles over a topological space $X$ with a fiber $H$.
-The structure group of a general Hilbert bundle is the unitary group ${{}^{s}{\mathcal{U}}}(H)$ equipped with the strong operator topology.
-As was shown by Dixmier and Douady, every such bundle over a paracompact base space $X$ is trivial, that is, isomorphic to the trivial Hilbert bundle $X\times H\to X$.
-See [DD], Theorem 1.
+(In the following we will always suppose that all Hilbert bundles are locally trivial.) The structure group of a general Hilbert bundle is ${{}^{s}{\mathcal{U}}}(H)$, the unitary group of $H$ with the strong operator topology.
+This group is contractible [DD, Lemma 3], so every such bundle over a paracompact base space $X$ is trivial, that is, isomorphic to the trivial Hilbert bundle $X\times H\to X$.
 More generally, every numerable Hilbert bundle is trivial.
 However, automorphisms of the trivial Hilbert bundle $X\times H\to X$ are given by continuous maps $X\to{{}^{s}{\mathcal{U}}}(H)$ and preserve neither norm nor Riesz continuity of operator families, since the action of ${{}^{s}{\mathcal{U}}}(H)$ on $\mathcal{B}(H)$ by conjugation, $(u,a)\mapsto uau^{-1}$, is not continuous.
-Therefore, the notions of a norm continuous family of bounded operators or of a Riesz or graph continuous family of regular operators are not defined for a general Hilbert bundle.
+Therefore, the notions of a norm continuous family of bounded operators or a Riesz/graph continuous family of regular operators are not defined for a general Hilbert bundle.
 Nevertheless, the actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on the space $\mathcal{K}(H,H^{\prime})$ of compact operators are continuous, and thus the notion of a norm continuous family of compact operators $k_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ is well defined.
-By the same reason, the notion of a graph continuous family of regular operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ with compact resolvents is well defined.
-Applying Theorem A to such a family, we obtain the following result.
-Theorem C.
+By the same reason, the notion of a graph continuous family of regular operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ with compact resolvents is well defined (Proposition 2.3).
+Applying Theorem 1 to such a family, we obtain the following result.
+Theorem 3.
 Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be numerable Hilbert bundles over a topological space $X$.
-Let $\mathcal{A}$ be a graph continuous family of regular operators $\mathcal{A}_{x}\in\mathcal{R}_{K}(\mathcal{H}_{x},\mathcal{H}^{\prime}_{x})$ with compact resolvents.
-Suppose that a global trivialization of one of these Hilbert bundles, $\mathcal{H}\cong X\times H$ or $\mathcal{H}^{\prime}\cong X\times H^{\prime}$, is fixed.
-Then there is a global trivialization of the second bundle taking $\mathcal{A}$ to a Riesz continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$.
-We prove Theorems B and C in Section 3.
-Self-adjoint operators.
+Let $\mathcal{A}$ be a graph continuous family of regular operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}^{\prime}_{x}$ with compact resolvents.
+Suppose that a trivialization of one of these Hilbert bundles, $\mathcal{H}\cong X\times H$ or $\mathcal{H}^{\prime}\cong X\times H^{\prime}$, is fixed.
+Then there is a trivialization of the second bundle taking $\mathcal{A}$ to a Riesz continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$.
+Self-adjoint case.
 For self-adjoint operators, multiplication $A\mapsto Au$ or $A\mapsto uA$ should be replaced by the conjugation $A\mapsto uAu^{-1}$.
-Instead of independent trivializations of two Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, as in Theorem C, one should look for a one trivialization of a given Hilbert bundle $\mathcal{H}$.
+Instead of independent trivializations of two Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, as in Theorem 3, one should look for a one trivialization of a given Hilbert bundle $\mathcal{H}$.
 This narrows down available options drastically.
-In particular, Theorem A has no direct analogue for self-adjoint operators, even invertible ones.
-A graph continuous family of self-adjoint operators with compact resolvents may fail to be Riesz continuous in any trivialization of $\mathcal{H}$, as we will explain below.
-The next theorem shows that the obstruction to the existence of such a trivialization has purely local nature.
-From local to global trivializations.
-Let $\mathcal{A}$ be a family of regular operators $\mathcal{A}_{x}\in\mathcal{R}(\mathcal{H}_{x})$ acting on fibers of a Hilbert bundle $\mathcal{H}$.
-Following Ivanov [$I_{2}$], we say that a trivialization $\tau\colon\left.\mathcal{H}\right|_{Y}\to Y\times H$ of $\mathcal{H}$ over a subset $Y\subset X$ is adapted to $\mathcal{A}$ if it takes $\mathcal{A}$ to a Riesz continuous map $\tau_{*}\mathcal{A}\colon Y\to\mathcal{R}(H)$.
-We say that $\mathcal{A}$ is Riesz quasi-continuous if, for every $x\in X$, it admits an adapted trivialization over a neighborhood of $x$.
-Theorem D.
-Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$, and let $\mathcal{A}$ be a family of self-adjoint operators with compact resolvents acting on fibers of $\mathcal{H}$.
-Suppose that $\mathcal{A}$ is Riesz quasi-continuous (that is, admits local adapted trivializations).
-Then there exists a global trivialization $\tau\colon\mathcal{H}\to X\times H$ adapted to $\mathcal{A}$.
+In particular, Theorem 1 has no direct analogue for self-adjoint operators, even invertible ones.
+Let $\mathcal{A}$ be a family of regular operators acting on fibers of a locally trivial Hilbert bundle $\mathcal{H}$ over $X$.
+Following Ivanov [I3], we say that a trivialization $\tau\colon\left.\mathcal{H}\right|_{Y}\to Y\times H$ of $\mathcal{H}$ over a subset $Y\subset X$ is adapted“Fully adapted” in terms of Ivanov to $\mathcal{A}$ if it takes $\mathcal{A}$ to a Riesz continuous map $\tau_{*}\mathcal{A}\colon Y\to\mathcal{R}(H)$.
+In particular, $\tau$ is adapted to a family $\mathcal{A}$ of bounded operators if it takes $\mathcal{A}$ to a norm continuous map $\tau_{*}\mathcal{A}\colon Y\to\mathcal{B}(H)$.
+Spectral projections of a graph continuous family $\mathcal{A}$ of self-adjoint operators with compact resolvents give rise to what we call a compatible family of local bundle projections.
+A trivialization is adapted to $\mathcal{A}$ if and only if it is adapted to these projections.
+Our main results in the self-adjoint case, Theorems 4 and 5, are concerned with such families of projections, and we deduce results concerning self-adjoint operators from these two theorems.
+Polarizations.
+We say that $p\colon\left.\mathcal{H}\right|_{Y}\to\left.\mathcal{H}\right|_{Y}$ is a bundle projection of $\mathcal{H}$ over $Y\subset X$ if $p$ is a continuous bundle map such that each $p_{x}$ is a projection.
+We say that two bundle projections, $p$ over $Y$ and $p^{\prime}$ over $Y^{\prime}$, are compatible (resp.
+finitely compatible) if the difference $x\mapsto p_{x}-p^{\prime}_{x}$ is a norm continuous family of compact (resp.
+finite rank) operators over $Y\cap Y^{\prime}$.
+Locally, a polarization is defined by a bundle projection $p$ over an open subset $Y\subset X$ such that each $p_{x}$ has infinite-dimensional range and kernel.
+Globally, a polarization is defined by an open covering $(X_{i})$ of $X$ together with a collection of pairwise compatible local polarizations $p_{i}$ over $X_{i}$.
+Following Ivanov [I2], we call such a collection $(X_{i},p_{i})$ a polarization atlas and define a polarization as an equivalence class of polarization atlases (as usual, two atlases are equivalent if their union is also an atlas).
+A finite polarization is defined in the same way, but with “compatible” replaced by “finitely compatible”.
+See Section 5 for details.
+Let $\Pi$ be a polarization of $\mathcal{H}$ given by an atlas $(X_{i},p_{i})$.
+We say that a (local or global) trivialization of $\mathcal{H}$ is adapted to $\Pi$ if it is adapted to each $p_{i}$ (that is, takes each $p_{i}$ to a norm continuous family).
+We say that a polarization is tame if it admits a local adapted trivialization over a neighborhood of each point $x\in X$.
+The notion of a polarization introduced by Ivanov in the appendix to [I2] is essentially equivalent to a tame finite polarization in our terminology.
+Theorem 4.
+Let $\Pi$ be a tame polarization of a locally trivial Hilbert bundle $\mathcal{H}$ over a paracompact space $X$.
+Then $\Pi$ admits a global adapted trivialization.
 Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
-Applying the last statement of the theorem to the lifting of $\mathcal{A}$ to $X\times[0,1]$, we see that every two adapted trivializations are homotopic and thus the homotopy class of a Riesz continuous map $\tau_{*}\mathcal{A}\colon X\to\mathcal{R}_{K}^{\mathrm{sa}}$ does not depend on the choice of an adapted trivialization $\tau$.
-For a family $\mathcal{A}$ satisfying assumptions of the theorem, this allows to define the $K^{1}$ index of $\mathcal{A}$ as the index of the norm continuous family $\bm{\chi}(\tau_{*}\mathcal{A})$ of bounded Fredholm self-adjoint operators.
-The main case of Theorem D is the case of families consisting of operators which are neither essentially positive nor essentially negative.
-In this case a similar trivialization theorem was recently proved by different methods and under much weaker continuity assumptions by N.
-Ivanov.
-See [$I_{2}$], Theorems 4.6 and 4.7.
-In combination with Lemma 5.1 from his paper [$I_{3}$], his trivialization theorem implies Theorem D for families consisting of operators which are neither essentially positive nor essentially negative.
-Reading Ivanov’s paper [$I_{2}$] led the author to realizing that a part of the original proof of Theorem D can be separated from the rest and has an independent interest.
-In more details, Ivanov mentioned in the introduction to the first version of [$I_{2}$] that there is a notion of a polarization of a Hilbert bundle standing behind some proofs in [$I_{2}$] (he worked out the details in the current version of [$I_{2}$]).
-In Section 4 we introduce the notion of a polarization of a Hilbert bundle related to Ivanov’s polarizations in the same way as compact operators are related to finite rank ones, and prove an analogue of his trivialization theorem.
-See Theorem 4.1.
-From graph to Riesz continuity.
-Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators with compact resolvents acting on fibers of a Hilbert bundle $\mathcal{H}$, as above.
-The first obstruction to existence of an adapted trivialization is caused by the fact that $\mathcal{R}^{\mathrm{sa}}_{K}$ is path connected in the graph topology, but has three connected components in the Riesz topology.
-Two points from different components of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$ (for example, an essentially positive and an essentially negative operator) may be connected by a graph continuous path in $\mathcal{R}_{K}^{\mathrm{sa}}(H)$, but every trivialization of the trivial Hilbert bundle $[0,1]\times H$ will leave this path Riesz discontinuous.
+In particular, every two adapted trivializations are homotopic (as trivializations adapted to $\Pi$).
+The next theorem shows that a modest assumption on the base space is enough for existing of an adapted trivialization.
+Let us say that a paracompact space $X$ has locally finite covering dimension if each point of $X$ has a neighborhood of finite covering dimension (we do not require this neighborhood to be open or small).
+In particular, every space of finite covering dimension has locally finite covering dimension.
+Theorem 5.
+Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$ of locally finite covering dimension.
+Then every polarization of $\mathcal{H}$ is tame and thus admits a global adapted trivialization.
+Self-adjoint operators and adapted trivializations.
+We want to know whether a family $\mathcal{A}$ of self-adjoint operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}_{x}$ admits a global adapted trivialization.
+Clearly, for that to be true, $\mathcal{A}$ has to admit local adapted trivializations (by that we mean that for every $x\in X$ there is an adapted trivialization over a neighborhood of $x$).
+Theorem 4 implies the following result, which shows that for essentially unitary operators this local condition is also sufficient.
+The author does not know whether this is true for arbitrary bounded self-adjoint operators, even if they are Fredholm.
+Theorem 6.
+Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$, $\mathcal{A}$ be a family of self-adjoint operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}_{x}$, and either all operators $\mathcal{A}_{x}$ are essentially unitary or they all have compact resolvents.
+Suppose that $\mathcal{A}$ admits local adapted trivializations.
+Then $\mathcal{A}$ admits a global adapted trivialization.
+Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
+In particular, every two such trivializations are homotopic (in the space of trivializations adapted to $\mathcal{A}$).
+The main case of Theorem 6 is the case of families consisting of operators which are neither essentially positive nor essentially negative.
+In this case, a similar trivialization theorem was nearly simultaneously proved by different methods and under much weaker continuity assumptions by Ivanov.
+See [I2], Theorems 4.5–4.7.
+It can be easily shown that his Theorems 4.6–4.7 imply Theorem 6 for families consisting of operators which are neither essentially positive nor essentially negative.
+Ivanov mentioned in the introduction to the first version of [I2] that there is a notion of a polarization of a Hilbert bundle standing behind some proofs in [I2].
+This led the author to introducing her version of a polarization of a Hilbert bundle and proving an analogue of Ivanov’s trivialization theorems 4.6–4.7 in terms of polarizations (Theorem 4).
+Later, Ivanov wrote down the details of his notion of a polarization in the appendix to the current version of [I2] and rephrased the proofs of his Theorems 4.5–4.7 in terms of polarizations.
+See [I2], Theorems A.2–A.4.
+First local obstruction.
+Let $\mathcal{R}^{\mathrm{sa}}_{K}(H)$ denote the set of self-adjoint operators with compact resolvents.
+The first local obstruction to existence of an adapted trivialization for a graph continuous family $\mathcal{A}$ of operators $\mathcal{A}_{x}\in\mathcal{R}^{\mathrm{sa}}_{K}(\mathcal{H}_{x})$ is caused by the fact that $\mathcal{R}^{\mathrm{sa}}_{K}$ is path connected in the graph topology, but has three connected components in the Riesz topology.
+Two points from different components of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$ (for example, an essentially positive and an essentially negative operator) may be connected by a graph continuous path in $\mathcal{R}_{K}^{\mathrm{sa}}(H)$, but no trivialization of the trivial Hilbert bundle $0,1]\times H$ makes this path Riesz continuous.
+Second local obstruction.
 When all the operators $\mathcal{A}_{x}$ are contained in a fixed connected component of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}$, there is another kind of local obstruction which we now describe.
-Let $\lambda\in\mathbb{R}$ be chosen outside of the spectrum of $\mathcal{A}_{x_{0}}$, $x_{0}\in X$.
-Then the spectral projections $p_{x}=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})$ (here $\operatorname{\mathbb{1}}_{S}$ denotes the characteristic function of a subset $S\subset\mathbb{R}$) give rise to the Hilbert bundles $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$, with fibers $\mathcal{H}^{+}_{x}=\operatorname{Ran}(p_{x})$ and $\mathcal{H}^{-}_{x}=\operatorname{Ker}(p_{x})$, over some neighborhood of $x_{0}$.
-A local adapted trivialization of $\mathcal{H}$ near $x_{0}$ exists if and only if both $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ are locally trivial near $x_{0}$.
+For a fixed $x\in X$, choose $\lambda\in\mathbb{R}$ in the resolvent set of $\mathcal{A}_{x}$.
+Then $\lambda$ also lies in the resolvent set of $\mathcal{A}_{y}$ for $y$ close to $x$, and the spectral projections $p_{y}=\operatorname{\mathds{1}}_{[\lambda,+\infty)}(\mathcal{A}_{y})$ determine a continuous map (a bundle projection) $p\colon\left.\mathcal{H}\right|_{Y}\to\left.\mathcal{H}\right|_{Y}$ over some neighborhood $Y$ of $x$ (here $\operatorname{\mathds{1}}_{S}$ denotes the characteristic function of a subset $S\subset\mathbb{R}$).
 See Proposition 7.1.
+A trivialization of $\mathcal{H}$ over $Y$ is adapted to $\mathcal{A}$ if and only if it is adapted to $p$ (Proposition 7.2).
 Suppose first that all the operators $\mathcal{A}_{x}$ are essentially negative.
-Since the norm and strong topologies coincide on the space of projections of a fixed finite rank, an obstruction to existence of local adapted trivializations near $x_{0}$ arises when the rank of $p_{x}$ is not locally constant near $x_{0}$.
-In other words, there are eigenvalues of $\mathcal{A}_{x}$ going to +$\infty$+ ∞ when $x\to x_{0}$.
-If this is not the case and the operators $\mathcal{A}_{x}$ are uniformly bounded from above in a neighborhood of each point $x_{0}\in X$, then $\mathcal{A}$ is Riesz continuous in every trivialization of $\mathcal{H}$.
-The situation with essentially positive operators is completely the same, but +$\infty$+ ∞ and bounds from above should be replaced by $\infty$- ∞ and bounds from below.
-See, e.g., Rellich’s example [Ka, Example V-4.14].
-Suppose now that all $\mathcal{A}_{x}$ are contained in the third connected component of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}(H)$, ${{}^{r}\mathcal{R}}_{K}^{\star}(H)$, which consists of operators that are neither essentially positive nor essentially negative.
-Then both Hilbert bundles $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ have infinite rank.
-By a result of Dixmier and Douady, every separable Hilbert bundle of infinite rank over a paracompact base space of finite covering dimension is trivial.
-Recall that $X$ is said to have finite covering dimension if there is an $n\in\mathbb{N}$ such that every finite open covering of $X$ has a finite open refinement of order $\leqslant n$.
-Let us say that a space has locally finite covering dimension if each point has a neighborhood of finite covering dimension.
-Using the aforementioned result of Dixmier and Douady, we prove a partial analogue of Theorem C for self-adjoint operators.
-Theorem E.
-Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a metrizable space $X$ of locally finite covering dimension.
-Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators $\mathcal{A}_{x}\in\mathcal{R}^{\star}_{K}(\mathcal{H}_{x})$, with compact resolvents, that are neither essentially positive nor essentially negative.
-Then $\mathcal{A}$ becomes Riesz continuous in some global trivialization of $\mathcal{H}$.
-Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
-Since the group ${{}^{s}{\mathcal{U}}}(H)$ is contractible, Theorem E implies that the identity map ${{}^{r}\mathcal{R}}^{\star}_{K}\to{{}^{g}\mathcal{R}}^{\star}_{K}$ is a weak homotopy equivalence.
+On the space of projections of a fixed finite rank the norm and strong topologies coincide [DD, Lemma 5], so there is an obstruction to existence of local adapted trivializations near $x$ if the rank of $p_{y}$ is not locally constant near $x$.
+In other words, there are eigenvalues of $\mathcal{A}_{y}$ going to $\infty$ when $y\to x$.
+If this is not the case and the operators $\mathcal{A}_{y}$ are uniformly bounded from above in a neighborhood of each point $x\in X$, then $\mathcal{A}$ is Riesz continuous in every trivialization of $\mathcal{H}$.
+The situation with essentially positive operators is completely the same, but $\infty$ and bounds from above should be replaced by -$\infty$ and bounds from below.
+A graph continuous family of essentially positive (resp, essentially negative) self-adjoint operators with compact resolvents may have no local adapted trivializations in a neighborhood of $x\in X$, even when the base space $X$ is an interval.
+See Rellich’s example [Kat, Example V-4.14], or [P3, Example 8.2] for more details.
+Suppose now that all $\mathcal{A}_{x}$ are contained in the third connected component ${{}^{r}\mathcal{R}}_{K}^{\star}(\mathcal{H}_{x})$ of ${{}^{r}\mathcal{R}}_{K}^{\mathrm{sa}}(\mathcal{H}_{x})$, which consists of operators that are neither essentially positive nor essentially negative.
+In general, such a family, as well as its spectral sections, may have no local adapted trivializations.
+See Example 7.5.
+However, Theorem 5 implies the following result.
+Theorem 7.
+Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$ of locally finite covering dimension.
+Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}_{x}$, with compact resolvents, which are neither essentially positive nor essentially negative.
+Then $\mathcal{A}$ admits a global adapted trivialization.
+Since the group ${{}^{s}{\mathcal{U}}}(H)$ is contractible, Theorem 7 implies that the identity map ${{}^{r}\mathcal{R}}^{\star}_{K}\to{{}^{g}\mathcal{R}}^{\star}_{K}$ is a weak homotopy equivalence.
 The stronger result, of a homotopy equivalence of this map, is proven in the previous paper of the author by a different method.
-See [$P_{3}$], Theorems B and C.
-Theorem E does not hold without additional assumptions for general metrizable base spaces, even compact, as we demonstrate by Example 5.5.
-Elliptic operators on closed manifolds.
-Let us describe an important class of situations where Theorems C and D may be useful.
-Let $\mathcal{M}\to X$ be a smooth fibration whose fibers $\mathcal{M}_{x}$ are smooth closed manifolds, and let $\mathcal{E}$, $\mathcal{E}^{\prime}$ be smooth vector bundles over $\mathcal{M}$.
-Let $\mathcal{D}_{x}$ be a continuous (in an appropriate sense) family of elliptic pseudodifferential operators on $\mathcal{M}_{x}$, of positive order, acting from fibers of $\mathcal{E}$ to fibers of $\mathcal{E}^{\prime}$.
-A local trivialization of $\mathcal{M}$ turns $\mathcal{D}$ to a family of elliptic operators acting between fibers of fixed vector bundles $E$ and $E^{\prime}$ over a fixed closed manifold $M$.
-The bounded transform takes this new family to a continuous family of zeroth order elliptic operators, which is norm continuous as a family of bounded operators $L^{2}(M;E)\to L^{2}(M;E^{\prime})$.
-Hence the family of unbounded operators $\mathcal{D}_{x}$ is Riesz continuous in this local trivialization.
-However, the fibration $\mathcal{M}\to X$ is in general non-trivial, and thus no such “geometric” global trivialization exists in the general case.
-The structure group of the corresponding Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, where $\mathcal{H}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$ and $\mathcal{H}^{\prime}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}^{\prime}_{x})$, has the topology coarser than the norm topology, so Kuiper’s theorem cannot be used in order to trivialize $\mathcal{H}$ and $\mathcal{H}^{\prime}$.
-Dixmier-Douady theorem, on the other hand, can be used, but trivializations it provides do not take $\mathcal{D}$ to a Riesz continuous family of operators.
-Theorem C closes this gap, showing that such a family $\mathcal{D}$ can be made Riesz continuous by appropriate global trivializations of the Hilbert bundles $\mathcal{H}$ and $\mathcal{H}^{\prime}$, though at the cost of losing the differential and geometric structure.
-Theorem D solves the same problem for families of self-adjoint elliptic operators acting on fibers of $\mathcal{E}$.
-The author arrived at Theorem D in an attempt to understand a step in the proof of a result of Melrose and Piazza.
-See [$MP_{1}$], Proposition 1.
-Melrose and Piazza deal with a family of self-adjoint differential operators of first order over a smooth fibration $\mathcal{M}\to X$, as above, and show that if the $K^{1}$-index of the family vanishes, then the family admits a spectral section.
-In their proof, they claim that the $L^{2}$ spaces of the fibers, $L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$, can be identified so that all operators act on a fixed Hilbert space $H$.
+See [P3], Theorems B and C.
+Example 7.5 below shows that Theorem 7 does not hold for general base spaces, even compact, without additional assumptions.
+The author does not know whether an analogue of Theorem 7 (probably with stronger assumptions on $X$ and for maps $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}^{\star}_{F}$) holds for Fredholm operators.
+See Remark 7.4.
+$K^{1}$ index for graph continuous families.
+Let $\mathcal{A}$ be a family satisfying assumptions of Theorem 7.
+Then $\mathcal{A}$ admits an adapted trivialization $\tau\colon\mathcal{H}\to H_{X}$, which leads to a norm continuous map $\alpha=\bm{\chi}(\tau_{*}\mathcal{A})\colon X\to\mathcal{B}_{F}^{\star}(H)$.
+The homotopy class of $\alpha$ does not depend on the choice of $\tau$.
+This allows to define the index of $\mathcal{A}$ as the classical index of $\alpha$, $\operatorname{ind}\mathcal{A}:=\operatorname{ind}\alpha=[\alpha]\in[X,\mathcal{B}_{F}^{\star}]=K^{1}(X)$.
+Another approach to the index of graph continuous families of self-adjoint Fredholm operators is provided by the previous paper [P3] of the author or by results of Joachim [J].
+An approach of Ivanov [I1, I2] works even under much weaker continuity assumptions than graph continuity.
+The same reasoning as for the $K^{0}$ index (Section 4) shows that, for a family admitting an adapted trivialization, the index given by the formula above coincides with Ivanov’s index.
+Applications to elliptic operators.
+The author arrived at Theorem 6 in an attempt to understand a step in the proof of a result of Melrose and Piazza.
+See [MP1], Proposition 1.
+They deal with a family $\mathcal{D}$ of self-adjoint differential operators $\mathcal{D}_{x}$ of order $1$ parametrized by $x\in X$ such that $\mathcal{D}_{x}$ acts on the fibers of a vector bundle $\mathcal{E}_{x}$ over a smooth closed manifold $\mathcal{M}_{x}$.
+More precisely, $\mathcal{M}\to X$ is a smooth fibration and $\mathcal{E}$ is a smooth vector bundle over $\mathcal{M}$.
+Melrose and Piazza showed that if the $K^{1}(X)$-index of $\mathcal{D}$ vanishes, then $\mathcal{D}$ admits a spectral section.
+In their proof, they claim that the $L^{2}$-spaces of the fibers, $\mathcal{H}_{x}=L^{2}(\mathcal{M}_{x};\mathcal{E}_{x})$, can be identified so that all operators act on a fixed Hilbert space $H$.
 Unfortunately, they did not explain how such an identification can be obtained.
 The author was not able to find a simple explanation of this step.
-Our Theorem D, as well as recent results [$I_{2}$, Theorems 4.5 or 4.6] or [$I_{3}$] of N.
-Ivanov, allows to overcome this difficulty.
-Similarly, Theorem C allows to pass to a fixed Hilbert space in the proof of [$MP_{2}$, Proposition 2], an odd analog of [$MP_{1}$, Proposition 1].
-Elliptic operators on manifolds with boundary.
-When the fibers $\mathcal{M}_{x}$ are compact manifolds with boundary, the situation becomes more complicated.
-In this case $\mathcal{D}_{x}$ is not even a closed operator usually.
-One needs to add an elliptic boundary condition $\mathcal{B}_{x}$ to $\mathcal{D}_{x}$, and then such a pair $(\mathcal{D}_{x},\mathcal{B}_{x})$ determines a regular operator with compact resolvents.
-Taking a local trivialization of $\mathcal{M}$ as above, one obtains a family of boundary value problems $(\mathcal{D}_{x},\mathcal{B}_{x})$ on a fixed manifold $M$.
-If both $\mathcal{D}_{x}$ and $\mathcal{B}_{x}$ depend continuously on $x$ in an appropriate sense, then the corresponding family of regular operators with compact resolvents is graph continuous.
-See [$P_{1}$, Appendix A.5] or [BZ, Corollary A.6.3].
-Theorem C shows that such a family $\mathcal{D}$ can be made Riesz continuous by appropriate global trivializations of the corresponding Hilbert bundles.
-Suppose now that we have a family of self-adjoint elliptic boundary value problems.
-As before, a local trivialization of $\mathcal{M}$ gives rise to a family of boundary value problems on a fixed manifold $M$.
-Unfortunately, there are no criteria of Riesz continuity for families of unbounded operators arising from boundary value problems, except for several special cases.
-Moreover, there are simple examples of such families that are not Riesz quasi-continuous, see e.g.
-Rellich’s example.
-Theorem E, however, requires only graph continuity of a family of unbounded operators, so it can be applied to families of elliptic self-adjoint boundary value problems, provided that they are neither essentially positive nor essentially negative and the space of parameters is good enough.
-Acknowledgment.   I am grateful to N. V. Ivanov for attracting my attention to the work of Dixmier and Douady [DD] and for helpful remarks.
+The structure group of the corresponding Hilbert bundle $\mathcal{H}$ has the topology coarser than the norm topology, so Kuiper’s theorem cannot be used to trivialize $\mathcal{H}$.
+Dixmier-Douady theorem, on the other hand, can be used, but trivializations it provides do not take $\mathcal{D}$ to a Riesz continuous family of operators.
+At the same time, the proof of Melrose and Piazza uses essentially Riesz continuity of $\mathcal{D}$.
+Theorem 6, as well as recent results [I2, Theorems 4.5 or 4.6] of Ivanov, allows to overcome this difficulty by turning $\mathcal{D}$ to a Riesz continuous family (though at the cost of losing the differential and geometric structure).
+Similarly, Theorem 3 allows to pass to a fixed Hilbert space in the proof of [MP2, Proposition 2], an odd analog of [MP1, Proposition 1].
+Another framework where our results may be useful is that of elliptic boundary value problems.
+A continuous, in an appropriate sense, family of elliptic operators of positive order on a compact manifold with boundary and with elliptic boundary conditions leads to a graph continuous family of regular operators with compact resolvents acting between the $L^{2}$-spaces of the corresponding vector bundles.
+See [P1, Appendix A.5] or [BZ, Corollary A.6.3].
+A fixed manifold with boundary may be replaced by a family of such manifolds, as above.
+Theorems 3 and 7 show that such a family of unbounded operators may be turned into a Riesz continuous family by appropriate global trivializations of the corresponding Hilbert bundles, both in the non-self-adjoint and self-adjoint case.
+Acknowledgments.
+This work was partially supported by ISF grants no.
+431/20 and 876/20 and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant no.
+101001677).
+It was done during my postdoctoral fellowship at the Technion – Israel Institute of Technology; the revised version was prepared during my postdoctoral fellowship at the University of Haifa.
+I am grateful to N.
+V.
+Ivanov for his interest in my work and, in particular, for attracting my attention to the work of Dixmier and Douady [DD].
+The exposition in this paper benefitted a lot from numerous suggestions of the anonymous referee, to whom I am deeply grateful.
 2 Preliminaries
 Grassmanian.
-The set $\operatorname{Gr}(H)$ of closed subspaces of $H$ is called the Grassmanian.
-An orthogonal projection may be identified with its range, which provides a bijection between $\operatorname{Gr}(H)$ and $\mathcal{P}(H)$.
-The Grassmanian is equipped with the topology (and even metric) induced by this bijection from the norm on the space $\mathcal{P}(H)$ of projections.
-We will use closed subspaces and projections by turn, depending on what is more convenient at the moment.
+The set $\mathcal{G}r(H)$ of closed subspaces of $H$ is called the Grassmanian.
+An orthogonal projection may be identified with its range, which provides a bijection between $\mathcal{G}r(H)$ and the space of projections $\mathcal{P}(H)$.
+The Grassmanian is equipped with the topology (and even metric) induced by this bijection from the norm on $\mathcal{P}(H)$.
+We will use closed subspaces and projections interchangeably, depending on what is more convenient at the moment.
 Canonical decomposition.
 Recall a standard construction that will be used throughout the paper.
-The trivial Hilbert bundle over $\operatorname{Gr}(H)$ with the fiber $H$ is canonically decomposed into the direct sum
-$\operatorname{Gr}(H)\times H=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$
-of two Hilbert bundles, whose fibers over $L\in\operatorname{Gr}$ are $\mathcal{H}^{+}_{L}=L$ and $\mathcal{H}^{-}_{L}=L^{\bot}$, the orthogonal complement of $L$ in $H$.
-This decomposition is locally trivial in the following sense: for every $L_{0}\in\operatorname{Gr}(H)$ there is a continuous map $g\colon B\to\mathcal{U}(H)$, where $B$ is the open ball of radius $1$ around $L_{0}$, such that $L=g_{L}L_{0}$ for every $L\in B$.
-In terms of projections, this is written as $p=g_{p}p_{0}g_{p}^{*}$ for $\left\|p-p_{0}\right\|<1$, where $p_{0}$ is the projection onto $L_{0}$.
-See the proof of [Ka, Theorem I.6.32] for the construction of such a map $g$.
-An important subspace of $\operatorname{Gr}(H)$ is
-$\operatorname{Gr}^{\star}(H)=\left\{L\in\operatorname{Gr}(H)\,|\,\dim L=\infty=\operatorname{codim}L\right\}.$
-Let $L_{0}\in\operatorname{Gr}^{\star}(H)$ be a fixed subspace.
-The map $\mathcal{U}(H)\to\operatorname{Gr}^{\star}(H)$, $u\mapsto uL_{0}$ is a locally trivial principal bundle with the contractible structure group $\mathcal{U}(L_{0})\times\mathcal{U}(L_{0}^{\bot})$ over a paracompact base space and therefore has a section
-(2.1) $g\colon\operatorname{Gr}^{\star}(H)\to\mathcal{U}(H),\;\;\text{so that}\;\;L=g_{L}L_{0}\;\;\text{for all}\;\;L\in\operatorname{Gr}^{\star}(H).$
-It follows that the restrictions of $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ to $\operatorname{Gr}^{\star}$ are trivial Hilbert bundles with the structure group $\mathcal{U}$.
-In addition, contractibility of the total space $\mathcal{U}(H)$ and the fiber $\mathcal{U}\times\mathcal{U}$ implies contractibility of the base space $\operatorname{Gr}^{\star}$.
+The trivial Hilbert bundle over $\mathcal{G}r(H)$ with the fiber $H$ is canonically decomposed into the direct sum
+$\mathcal{G}r(H)\times H=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$
+of two Hilbert bundles, whose fibers over $L\in\mathcal{G}r$ are $\mathcal{H}^{+}_{L}=L$ and $\mathcal{H}^{-}_{L}=L^{\bot}$, the orthogonal complement of $L$ in $H$.
+This decomposition is locally trivial in the following sense: for every $L_{0}\in\mathcal{G}r(H)$ there is a continuous map $g\colon B\to\mathcal{U}(H)$, where $B$ is the open ball of radius $1$ around $L_{0}$, such that $L=g_{L}L_{0}$ for every $L\in B$.
+In terms of projections, this is written as $p=g_{p}p_{0}g_{p}^{*}$ for $\lVert p-p_{0}\rVert<1$, where $p_{0}$ is the projection onto $L_{0}$.
+See the proof of [Kat, Theorem I.6.32] for the construction of such a map $g$.
+An important subspace of $\mathcal{G}r(H)$ is
+$\mathcal{G}r^{\star}(H)=\left\{L\in\mathcal{G}r(H)\mid\dim L=\infty=\operatorname{codim}L\right\}.$
+Let $L_{0}\in\mathcal{G}r^{\star}(H)$ be a fixed subspace.
+The map $\mathcal{U}(H)\to\mathcal{G}r^{\star}(H)$, $u\mapsto uL_{0}$ is a locally trivial principal bundle over a paracompact base space, whose structure group $\mathcal{U}(L_{0})\times\mathcal{U}(L_{0}^{\bot})$ is contractible.
+Therefore, $\mathcal{U}(H)\to\mathcal{G}r^{\star}(H)$ has a section
+(2.1) $g\colon\mathcal{G}r^{\star}(H)\to\mathcal{U}(H),\;\;\text{so that}\;\;L=g_{L}L_{0}\;\;\text{for all}\;\;L\in\mathcal{G}r^{\star}(H).$
+It follows that the restrictions of $\mathcal{H}^{+}$ and $\mathcal{H}^{-}$ to $\mathcal{G}r^{\star}$ are trivial Hilbert bundles with the structure group $\mathcal{U}$.
+In addition, contractibility of the total space $\mathcal{U}(H)$ and the fiber $\mathcal{U}\times\mathcal{U}$ implies contractibility of the base space $\mathcal{G}r^{\star}$.
 See [AS, proof of Lemma 3.6].
-Bounded transform.   Let us recall the definition and the main properties of the bounded transform. See, e.g., [Sc], Theorem 7.5 for the detail.
+Bounded transform.  Let us recall the definition and main properties of the bounded transform. See, e.g., [S], Theorem 7.5 for the detail.
 For a regular operator $A\colon H\to H^{\prime}$, the operator $1+A^{*}A$ is regular, self-adjoint, and surjective; its inverse $(1+A^{*}A)^{-1}$ is a positive bounded operator.
@@
 $\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{B}(H,H^{\prime}),\quad\bm{\chi}(A)=A(1+A^{*}A)^{-1/2},$
-defines the inclusion of the set $\mathcal{R}(H,H^{\prime})$ of regular operators to the closed unit ball
-$\mathcal{D}(H,H^{\prime})=\left\{a\in\mathcal{B}(H,H^{\prime})\,|\,\left\|a\right\|\leqslant 1\right\}.$
+defines the injective map from the set $\mathcal{R}(H,H^{\prime})$ of regular operators to the closed unit ball
+$\mathcal{D}(H,H^{\prime})=\left\{a\in\mathcal{B}(H,H^{\prime})\mid\lVert a\rVert\leqslant 1\right\}.$
 The image of $\bm{\chi}$ consists of strict contractions, that is, operators $a\in\mathcal{D}$ such that $1-a^{*}a$ is injective; this image is dense in $\mathcal{D}$.
@@
 (2.3) $\mathbf{w}_{A}=(1+A^{*}A)^{-1/2}\oplus A(1+A^{*}A)^{-1/2}=(1+A^{*}A)^{-1/2}\oplus\bm{\chi}(A).$
-The orthogonal projection onto $\Gamma_{A}$ is
+Let $\mathbf{p}_{A}$ denote the orthogonal projection onto $\Gamma_{A}$,
 (2.4) $\mathbf{p}_{A}=\mathbf{w}_{A}\mathbf{w}_{A}^{*}=\begin{pmatrix}(1+A^{*}A)^{-1}&(1+A^{*}A)^{-1}A^{*}\\ A(1+A^{*}A)^{-1}&1-(1+AA^{*})^{-1}\end{pmatrix}\in\mathcal{P}(H\oplus H^{\prime}).$
 In terms of $a=\bm{\chi}(A)$ this is written as follows:
-$\mathbf{w}_{A}=\tilde{\mathbf{w}}_{a}=\sqrt{1-a^{*}a}\oplus a,\quad\quad\mathbf{p}_{A}=\tilde{\mathbf{p}}_{a}=\tilde{\mathbf{w}}_{a}\tilde{\mathbf{w}}_{a}^{*}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix},$
-so $\mathbf{p}$ factors through a continuous map $\tilde{\mathbf{p}}\colon\mathcal{D}(H,H^{\prime})\to\mathcal{P}(H\oplus H^{\prime})$.
-Restricting $\tilde{\mathbf{p}}$ to the subspace ${{}^{r}\mathcal{R}}$ of $\mathcal{D}$, we see that the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is included into the following commutative square and thus is continuous:
-[…]
+(2.5) $\mathbf{w}_{A}=\sqrt{1-a^{*}a}\oplus a,\quad\quad\mathbf{p}_{A}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix}$,
+so $\mathbf{p}$ factors through a continuous map $\mathcal{D}(H,H^{\prime})\to\mathcal{P}(H\oplus H^{\prime})$.
+Restricting this map to the subspace $^{r}\mathcal{R}$} of $\mathcal{D}$, we see that the identity map ${{}^{r}\mathcal{R}}\to{{}^{g}\mathcal{R}}$ is continuous.
 Fredholm operators.
@@
 By the definition of Riesz topology, the bounded transform $\bm{\chi}\colon\mathcal{R}(H,H^{\prime})\to\mathcal{D}(H,H^{\prime})$ is Riesz-to-norm continuous.
-In the proof of Theorem A we also need another type of continuity of $\bm{\chi}$.
-Recall that a family of bounded operators $a_{x}\in\mathcal{B}(H,H^{\prime})$ is said to be continuous in the strong operator topology (or strongly continuous) if the family $a_{x}\xi$ of vectors in $H^{\prime}$ is continuous for every $\xi\in H$.
-A family $a_{x}$ is said to be continuous in the strong* operator topology if both families $a_{x}$ and $a_{x}^{*}$ are strongly continuous.
+In the proof of Theorem 1 we also need another type of continuity of $\bm{\chi}$.
+Let ${{}^{s}{\mathcal{B}}}(H,H^{\prime})$ denote the space of bounded operators $H\to H^{\prime}$ equipped with the strong operator topology.
+Recall that this topology is defined as the weakest topology for which the map ${{}^{s}{\mathcal{B}}}(H,H^{\prime})\to H^{\prime}$, $a\mapsto a\xi$ is continuous for every $\xi\in H$.
+The strong* operator topology on the same set is the weakest topology for which both $a\mapsto a\xi$ and $a\mapsto a^{*}\xi$ are continuous for every $\xi\in H$.
 Proposition 2.1.
@@
 (2.6) $a=\bm{\chi}(A),\quad q=(1+A^{*}A)^{-1/2}=(1-a^{*}a)^{1/2},\quad b=aq.$
-The projection $\mathbf{p}_{A}$ is given by the formula
-(2.6) $\mathbf{p}_{A}=\tilde{\mathbf{p}}_{a}=\begin{pmatrix}1-a^{*}a&\sqrt{1-a^{*}a}\,a^{*}\\ a\sqrt{1-a^{*}a}&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&qa^{*}\\ aq&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&b^{*}\\ b&aa^{*}\end{pmatrix}.$
+In these terms, the second equality of (2.5) is written as
+(2.7) $\mathbf{p}_{A}=\begin{pmatrix}q^{2}&qa^{*}\\ aq&aa^{*}\end{pmatrix}=\begin{pmatrix}q^{2}&b^{*}\\ b&aa^{*}\end{pmatrix}$.
 Since the map $A\mapsto b$ is given by the $(2,1)$-component of $\mathbf{p}_{A}$, it is graph-to-norm continuous.
@@
 Let $\varepsilon>0$.
-The operator $q_{0}$ is self-adjoint and injective and thus has dense range, so one can find $\eta\in H$ such that $\left\|q_{0}\eta-\xi\right\|<\varepsilon$.
-Let $V$ be a neighborhood of $A_{0}$ in ${{}^{g}\mathcal{R}}(H,H^{\prime})$ such that $\left\|(q-q_{0})\eta\right\|<\varepsilon$ and $\left\|(b-b_{0})\eta\right\|<\varepsilon$ for every $A\in V$.
+The operator $q_{0}$ is self-adjoint and injective and thus has dense range, so one can find $\eta\in H$ such that $\lVert q_{0}\eta-\xi\rVert<\varepsilon$.
+Let $V$ be a neighborhood of $A_{0}$ in ${{}^{g}\mathcal{R}}(H,H^{\prime})$ such that $\lVert(q-q_{0})\eta\rVert<\varepsilon$ and $\lVert(b-b_{0})\eta\rVert<\varepsilon$ for every $A\in V$.
 Then
-$\left\|a\xi-a_{0}\xi\right\|\;<\;\left\|\left(a-a_{0}\right)q_{0}\eta\right\|+2\varepsilon\;\leqslant\;\left\|a\left(q_{0}-q\right)\eta\right\|+\left\|\left(aq-a_{0}q_{0}\right)\eta\right\|+2\varepsilon$
-$\leqslant\;\left\|\left(q-q_{0}\right)\eta\right\|+\left\|\left(b-b_{0}\right)\eta\right\|+2\varepsilon<4\varepsilon$
+$\lVert a\xi-a_{0}\xi\rVert\;<\;\lVert\left(a-a_{0}\right)q_{0}\eta\rVert+2\varepsilon\;\leqslant\;\lVert a\left(q_{0}-q\right)\eta\rVert+\lVert\left(aq-a_{0}q_{0}\right)\eta\rVert+2\varepsilon$
+$\leqslant\;\lVert\left(q-q_{0}\right)\eta\rVert+\lVert\left(b-b_{0}\right)\eta\rVert+2\varepsilon<4\varepsilon$
 for every $A\in V$. It follows that $\bm{\chi}$ is graph-to-strong continuous at an arbitrary point $A_{0}\in\mathcal{R}$ and thus on the whole $\mathcal{R}$. $\square$
@@
 In fact, this reasoning provides a stronger result.
-It concerns another natural topology on $\mathcal{R}(H,H^{\prime})$, which is is strictly coarser then the graph topology, induced by the inclusion $\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow{{}^{s}{\mathcal{P}}}(H\oplus H^{\prime})$ from the strong operator topology on the space of projections.
-On the subset of self-adjoint operators it coincides with the strong resolvent topology, which is induced by the map $A\mapsto(A+i)^{-1}$, or equivalently by the map $A\mapsto(A-i)^{-1}$, from the strong operator topology on bounded operators.
-The reasoning in the proof of Proposition 2.1 shows that the bounded transform $\bm{\chi}\colon\mathcal{R}\to\mathcal{D}$ is continuous with respect to this coarser topology on $\mathcal{R}$ and the strong* topology on $\mathcal{D}$.
-Proposition 2.1 follows from a similar result for self-adjoint operators and the restriction of $\bm{\chi}$ to $\mathcal{R}^{\mathrm{sa}}$, $A\mapsto A(1+A^{2})^{-1/2}$, by the standard trick of passing from $A$ to the self-adjoint operator $\left(\begin{smallmatrix}0&A^{*}\\ A&0\end{smallmatrix}\right)$.
-For self-adjoint operators, a similar reasoning provides continuity of an arbitrary function $f\in C_{b}(\mathbb{R})$ with respect to the strong resolvent topology on $\mathcal{R}^{\mathrm{sa}}$ and the strong topology on $\mathcal{D}$.
-A different proof of this fact based on the spectral theorem is given by [RS], Theorem VIII.20(b).
-A similar result for compact spaces of parameters is given by [W], Proposition 1.1 of Ch.
-Wahl.
+It concerns another natural topology on $\mathcal{R}(H,H^{\prime})$ induced by the injection $\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow{{}^{s}{\mathcal{P}}}(H\oplus H^{\prime})$ from the strong operator topology on the space of projections.
+On $\mathcal{R}^{\mathrm{sa}}(H)$ it coincides with the strong resolvent topology.
+The reasoning in the proof of Proposition 2.1 shows that the bounded transform $\bm{\chi}\colon\mathcal{R}\to\mathcal{D}$ is continuous with respect to this coarser topology on $\mathcal{R}$ and the strong* operator topology on $\mathcal{D}$.
+For an arbitrary function $f\in C_{b}(\mathbb{R})$, a similar reasoning provides continuity of the map $f\colon\mathcal{R}^{\mathrm{sa}}(H)\to\mathcal{B}(H)$ with respect to the strong resolvent topology on $\mathcal{R}^{\mathrm{sa}}$ and the strong topology on $\mathcal{B}$.
+A different proof of such continuity of $f$ based on the spectral theorem is given by [RS], Theorem VIII.20(b).
+Conversely, one can deduce Proposition 2.1 from that theorem taking $f(x)=x(1+x^{2})^{-1/2}$ and using the standard trick of passing from $A$ to the self-adjoint operator $\left(\begin{smallmatrix}0&A^{*}\\ A&0\end{smallmatrix}\right)$.
+However, the direct proof of the proposition given above and not based on the spectral theorem seems to be more transparent.
 Action of the unitary groups.
@@
 These actions preserve the subsets $\mathcal{R}_{F}$ and $\mathcal{R}_{K}$.
-For $u\in\mathcal{U}(H^{\prime})$ and $v\in\mathcal{U}(H)$ we have
+Proposition 2.2.
+The actions of the groups $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ (equipped with the norm topology) on ${{}^{r}\mathcal{R}}(H,H^{\prime})$ and ${{}^{g}\mathcal{R}}(H,H^{\prime})$ are continuous.
+Proof. For $u\in\mathcal{U}(H^{\prime})$ and $v\in\mathcal{U}(H)$ we have
 $\bm{\chi}(uA)=u\bm{\chi}(A),\quad\bm{\chi}(Av)=\bm{\chi}(A)v,\quad\Gamma_{uA}=(1\oplus u)\Gamma_{A},\quad\Gamma_{Av}=(v^{-1}\oplus 1)\Gamma_{A},$
-so the actions of $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ (equipped with the norm topology) are continuous on both ${{}^{r}\mathcal{R}}$ and ${{}^{g}\mathcal{R}}$.
-Recall that ${{}^{s}{\mathcal{U}}}(H)$, the unitary group equipped with the strong operator topology, is a topological group.
-Indeed, the multiplication map $\mathcal{B}(H)\times\mathcal{B}(H)\to\mathcal{B}(H)$, $(a,b)\mapsto ab$, is strongly continuous on $S\times\mathcal{B}(H)$ for every bounded subset $S$ of $\mathcal{B}(H)$.
-In particular, it is strongly continuous on $\mathcal{U}(H)\times\mathcal{U}(H)$.
-For every $u,u_{0}\in\mathcal{U}(H)$ and $\xi\in H$,
-$\left\|(u^{-1}-u_{0}^{-1})\xi\right\|=\left\|u(u^{-1}-u_{0}^{-1})\xi\right\|=\left\|(u-u_{0})u_{0}^{-1}\xi\right\|,$
-so the inverse map $\mathcal{U}(H)\to\mathcal{U}(H)$, $u\mapsto u^{-1}=u^{*}$, is also strongly continuous.
-The actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{r}\mathcal{R}}$ and ${{}^{g}\mathcal{R}}$ are not continuous.
-However, continuity of their actions on $\mathcal{K}(H\oplus H^{\prime})$ provides the following result.
+which proves the proposition. $\square$
+Let ${{}^{s}{\mathcal{U}}}(H)$ denote the unitary group equipped with the strong operator topology.
+It is a topological group, see [RW, Lemma 1.3].
+The actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on $^{r}\mathcal{R}$} and $^{g}\mathcal{R}$} are not continuous, but continuity of the action of $^{s}{\mathcal{U}}$} on $\mathcal{K}$ provides the following result.
 Proposition 2.3.
-The actions of the groups ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$ are continuous
+The actions of the groups ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$ are continuous.
 Proof. The graph of an operator with compact resolvents belongs to the restricted Grassmanian
-(2.7) $\operatorname{Gr}_{K}(H\oplus H^{\prime})\cong\mathcal{P}_{K}(H\oplus H^{\prime})=\left\{p\in\mathcal{P}(H\oplus H^{\prime})\,|\,p-p_{\infty}\text{ is compact}\right\},$
+(2.8) $\mathcal{G}r_{K}(H\oplus H^{\prime})\cong\mathcal{P}_{K}(H\oplus H^{\prime})=\left\{p\in\mathcal{P}(H\oplus H^{\prime})\mid p-p_{\infty}\text{ is compact}\right\},$
 where $p_{\infty}\in\mathcal{P}(H\oplus H^{\prime})$ is the projection onto $0\oplus H^{\prime}$.
@@
 $\square$
-3 Non-self-adjoint operators
-[…]
-[…]
-we obtain a desired map $v\colon{{}^{g}\mathcal{R}}\to{{}^{s}{\mathcal{U}}}(H)$.
-If $\Gamma$ is considered as a Hilbert bundle over ${{}^{r}\mathcal{R}}$, then $\mathbf{w}$ becomes a trivialization in the category of bundles with the structure group $\mathcal{U}(H)$, so $v$ is continuous as a map ${{}^{r}\mathcal{R}}\to\mathcal{U}(H)$ as well.
+3 Proof of Theorem 1
+The idea of the proof can be breifly described as follows.
+The graphs $\Gamma_{A}$ of regular operators $A\colon H\to H^{\prime}$ form fibers of a locally trivial Hilbert bundle $\Gamma$ over ${{}^{g}\mathcal{R}}(H,H^{\prime})$ with the structure group $\mathcal{U}(H)$.
+Since $\mathcal{U}(H)$ is contractible, this bundle admits a trivialization $u\colon{{}^{g}\mathcal{R}}\times H\stackrel{{\scriptstyle{\approx}\hskip 1.63885pt}}{{\longrightarrow}}\Gamma$.
+The composition of the unitary operator $u_{A}\colon H\stackrel{{\scriptstyle{\approx}\hskip 1.63885pt}}{{\longrightarrow}}\Gamma_{A}\subset H\oplus H^{\prime}$ with the projection $H\oplus H^{\prime}\to H^{\prime}$ provides the contraction $\varphi_{A}\colon H\to H^{\prime}$.
+We define $\Phi$ by the formula $\bm{\chi}(\Phi_{A})=\varphi_{A}$ and show that this map satisfies all conclusions of the theorem.
 Isometries.  Let
-=$\mathcal{I}=\left\{u\in\mathcal{B}(H,H\oplus H^{\prime})\,|\,u^{*}u=\operatorname{Id}_{H}\right\}$
+$\mathcal{I}=\left\{u\in\mathcal{B}(H,H\oplus H^{\prime})\mid u^{*}u=\operatorname{Id}_{H}\right\}$
 be the space of isometries (i.e., isometric embeddings) $H\to H\oplus H^{\prime}$ equipped with the usual norm topology. The canonical map
-$\bm{\pi}\colon\mathcal{I}\to\operatorname{Gr}(H\oplus H^{\prime}),\quad u\mapsto\operatorname{Ran}(u).$
+$\bm{\pi}\colon\mathcal{I}\to\mathcal{G}r(H\oplus H^{\prime}),\quad u\mapsto\operatorname{Ran}(u).$
 takes an isometry $u$ to its range; the corresponding projection is $uu^{*}$.
 The image $\bm{\pi}(\mathcal{I})$ of this map consists of infinite-dimensional subspaces of $H\oplus H^{\prime}$.
-We will need the following simple corollary of Kuiper’s theorem [Ku] about contractibility of $\mathcal{U}(H)$.
-Lemma 3.1.
-The map $\bm{\pi}\colon\mathcal{I}\to\bm{\pi}(\mathcal{I})$ is a principal $\mathcal{U}(H)$-bundle admitting a continuous section (and thus is trivial).
-Proof.
 The group $\mathcal{U}(H)$ acts on $\mathcal{I}$ by the right multiplication.
-Every two isometries with the same range differ by an uniquely defined unitary operator $H\to H$, so this action turns $\bm{\pi}$ into a principal $\mathcal{U}(H)$-bundle.
-This bundle has a section by the same standard arguments as were used in the beginning of Section 2.
-Indeed, for every $u_{0}\in\mathcal{I}$ and $L_{0}=\bm{\pi}(u_{0})$ there is a continuous map $g$ from a neighborhood $B$ of $L_{0}$ to $\mathcal{U}(H\oplus H^{\prime})$ such that $L=g_{L}L_{0}$ for $L\in B$.
-The rule $L\mapsto g_{L}\cdot u_{0}$ determines a local section of $\bm{\pi}$ over $B$.
-Hence $\bm{\pi}$ is a locally trivial principal bundle over $\bm{\pi}(\mathcal{I})$ with the contractible structure group $\mathcal{U}(H)$.
-The base space $\bm{\pi}(\mathcal{I})$ is metric and thus paracompact.
-Therefore, $\bm{\pi}$ has a global section.
-$\square$
-Proof of Theorem A.
-The graphs of regular operators have infinite dimension, so ${{}^{g}\mathcal{R}}\subset\bm{\pi}(\mathcal{I})\subset\operatorname{Gr}(H\oplus H^{\prime})$.
-By Lemma 3.1, the restriction of $\bm{\pi}$ to ${{}^{g}\mathcal{R}}$ has a section.
-For the rest of the proof we fix such a section $u\colon{{}^{g}\mathcal{R}}\to\mathcal{I}$.
-In other words, $u$ is a graph-to-norm continuous map taking a regular operator $A$ to an isometry $u_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$.
+Kuiper’s theorem [Ku] together with the same standard arguments as were used in the beginning of Section 2 show that $\bm{\pi}\colon\mathcal{I}\to\bm{\pi}(\mathcal{I})$ is a principal $\mathcal{U}(H)$-bundle admitting a continuous section (and thus is trivial).
+Proof of Theorem 1.
+The graphs of regular operators have infinite dimension, so ${{}^{g}\mathcal{R}}\subset\bm{\pi}(\mathcal{I})\subset\mathcal{G}r(H\oplus H^{\prime})$.
+For the rest of the proof we fix a section of $\bm{\pi}\colon\mathcal{I}\to\bm{\pi}(\mathcal{I})$ and denote by $u\colon{{}^{g}\mathcal{R}}\to\mathcal{I}$ its restriction to $^{g}\mathcal{R}$}.
+Then $u$ is a graph-to-norm continuous map taking a regular operator $A$ to an isometry $u_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$.
 Let $p^{\prime}\colon H\oplus H^{\prime}\to H^{\prime}$ be the projection onto the second summand. The composition
@@
 $\square$
-Remark.
-Our construction of $v$ depends on the choice of a section $u$.
-Since $\bm{\pi}$ is a principal $\mathcal{U}(H)$-bundle, any other choice of a section differs from $u$ by the multiplication by a graph-to-norm continuous map ${{}^{g}\mathcal{R}}(H,H^{\prime})\to\mathcal{U}(H)$.
-Therefore, our construction determines $v$ up to the right multiplication by such a map.
-Proof of Theorem B.
-In order to prove the first part of the theorem we need to show that $\Phi$ is homotopic to the identity as a self-map of both ${{}^{r}\mathcal{R}}_{K}$ and ${{}^{g}\mathcal{R}}_{K}$.
+4 Applications of Theorem 1
+Proof of Theorem 2.
+In order to prove the first part of the theorem, we need to show that $\Phi$ is homotopic to the identity as a self-map of both ${{}^{r}\mathcal{R}}_{K}$ and ${{}^{g}\mathcal{R}}_{K}$.
 For the second part we need to prove a similar result for ${{}^{r}\mathcal{R}}_{F}$ and ${{}^{g}\mathcal{R}}_{F}$.
@@
 This part is more delicate because the action of $^{s}{\mathcal{U}}$} on ${{}^{g}\mathcal{R}}_{F}$ is not continuous, so we cannot directly apply a contraction of $^{s}{\mathcal{U}}$} for this purpose.
-However, we can use a part of [$P_{3}$, Theorem A] which states that the embedding ${{}^{g}\mathcal{R}}_{K}\hookrightarrow{{}^{g}\mathcal{R}}_{F}$ is a homotopy equivalence.
+However, we can use a part of [P3, Theorem A] which states that the embedding ${{}^{g}\mathcal{R}}_{K}\hookrightarrow{{}^{g}\mathcal{R}}_{F}$ is a homotopy equivalence.
 We already proved that the maps $\Phi\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ are homotopic; it follows that $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ are also homotopic.
@@
 $\square$
-Proof of Theorem C.
+Proof of Theorem 3.
 Suppose that a global trivialization of $\mathcal{H}^{\prime}$ is fixed.
@@
 In these trivializations $\mathcal{A}$ becomes a graph continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$.
-Let $v\colon{{}^{g}\mathcal{R}}(H,H^{\prime})\to{{}^{s}{\mathcal{U}}}(H)$ be a map provided by Theorem A.
+Let $v\colon{{}^{g}\mathcal{R}}(H,H^{\prime})\to{{}^{s}{\mathcal{U}}}(H)$ be a map provided by Theorem 1.
 Then the map $X\to{{}^{s}{\mathcal{U}}}(H)$, $x\mapsto v(\mathcal{A}_{x})$ is continuous, so
@@
 It follows that the composition $\tau_{v}^{-1}\circ\tau$ satisfies conclusion of the theorem.
-If instead a global trivialization of $\mathcal{H}$ is fixed, then one can interchange $\mathcal{H}$ and $\mathcal{H}^{\prime}$ by passing to the adjoint operators.
+If instead a global trivialization of $\mathcal{H}$ is fixed, then one can pass to the adjoint operators, swapping $\mathcal{H}$ and $\mathcal{H}^{\prime}$.
 Since the map $A\mapsto A^{*}$ is both graph and Riesz continuous, it follows from the first part of the proof that there is a global trivialization of $\mathcal{H}^{\prime}$ taking $\mathcal{A}$ to a Riesz continuous map.
@@
 See [A], Theorem A1.
-For an arbitrary parameter space $X$, $K^{0}(X)$ might be defined as the set $[X,\mathcal{B}_{F}]$ of homotopy classes of maps $X\to\mathcal{B}_{F}(H)$, with the group structure on $[X,\mathcal{B}_{F}]$ introduced in the standard way.
+For an arbitrary parameter space $X$, $K^{0}(X)$ might be defined as the set $X,\mathcal{B}_{F}]$ of homotopy classes of maps $X\to\mathcal{B}_{F}(H)$, with the group structure on $X,\mathcal{B}_{F}]$ introduced in the standard way.
 The class of a continuous map $\alpha\colon X\to\mathcal{B}_{F}$ in $K^{0}(X)$ is called the index of $\alpha$:
 $\operatorname{ind}(\alpha)=[\alpha]\in[X,\mathcal{B}_{F}]=K^{0}(X).$
-It can be easily seen, using linear homotopies, that the bounded transform map $\bm{\chi}\colon{{}^{r}\mathcal{R}}_{F}\to\mathcal{B}_{F}$ is homotopy inverse to the embedding $\mathcal{B}_{F}\hookrightarrow{{}^{r}\mathcal{R}}_{F}$, so both these maps are homotopy equivalencies.
-Therefore, the index of a Riesz continuous family of Fredholm operators is naturally defined as the index of its bounded transform:
-$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{r}\mathcal{R}}_{F}.$
-Let now $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ be a graph continuous family of Fredholm operators.
-The space ${{}^{g}\mathcal{R}}_{F}$ is also a classifying space for the functor $K^{0}$, see [J], Theorem 3.5(i) or [$P_{3}$], Theorem A.
+It can be easily seen, using linear homotopies, that the bounded transform $\bm{\chi}\colon{{}^{r}\mathcal{R}}_{F}\to\mathcal{B}_{F}$, from regular Fredholm operators to bounded Fredholm operators, is homotopy inverse to the embedding $\mathcal{B}_{F}\hookrightarrow{{}^{r}\mathcal{R}}_{F}$, so both these maps are homotopy equivalencies.
+Therefore, the index of a Riesz continuous map $\mathcal{A}\colon X\to{{}^{r}\mathcal{R}}_{F}$ is naturally defined as the index of its bounded transform:
+(4.1) $\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{r}\mathcal{R}}_{F}.$
+Let now $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ be a graph continuous map.
+The space ${{}^{g}\mathcal{R}}_{F}$ is also a classifying space for the functor $K^{0}$, see [J], Theorem 3.5(i) or [P3], Theorem A.
 Hence the homotopy class of $\mathcal{A}$ determines an element of $K^{0}(X)$ which can be called the index of $\mathcal{A}$:
 (4.2) $\operatorname{ind}(\mathcal{A})=[\mathcal{A}]\in[X,{{}^{g}\mathcal{R}}_{F}]=K^{0}(X).$
-By the definition, two graph continuous families of Fredholm operators are homotopic if and only if their indices coincide.
-Theorems A and B of the present paper allow to relate the index of $\mathcal{A}$ given by (3.1) to the Fredholm index of bounded operators as follows.
-The composition of $\mathcal{A}$ with a map $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ provided by Theorem A is Riesz continuous, so $\bm{\chi}\circ\Phi\circ\mathcal{A}$ is a norm continuous family of bounded Fredholm operators.
-Theorem B provides the equality
-$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\Phi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F},$
+Theorems 1 and 2 of the present paper allow to relate the index of $\mathcal{A}$ given by (4.2) to the Fredholm index of bounded operators as follows.
+The composition of $\mathcal{A}$ with a map $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ provided by Theorem 1 is Riesz continuous, so $\bm{\chi}\circ\Phi\circ\mathcal{A}$ is a norm continuous map $X\to\mathcal{B}_{F}$.
+Theorem 2 implies
+(4.3) $\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\Phi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F},$
 which can be used as the definition of the index of $\mathcal{A}$.
-Note that Theorem B implies that the homotopy class of $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{r}\mathcal{R}}_{F}$ does not depend on the choice of $v$, so $\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$ is also independent of this choice.
-Equivalently, the index can be defined in terms of graphs of operators as described in the Introduction.
-For a Riesz continuous family $\mathcal{A}\colon X\to\mathcal{R}_{F}$ (in particular, for a norm continuous family of bounded Fredholm operators) the two definitions of the index coincide.
-Indeed, $\Phi(\mathcal{A}_{x})=\mathcal{A}_{x}\cdot v(\mathcal{A}_{x})$, so $\Phi\circ\mathcal{A}$ differs from $\mathcal{A}$ by the multiplication by a map $v\circ\mathcal{A}\colon X\to\mathcal{U}(H)$.
-If $\mathcal{A}$ is Riesz continuous, then $v\circ\mathcal{A}$ is norm continuous.
-Since $\mathcal{U}(H)$ is contractible, $\Phi\circ\mathcal{A}$ and $\mathcal{A}$ are homotopic as maps from $X$ to ${{}^{r}\mathcal{R}}_{F}$ and thus their indices coincide.
+It follows from Theorem 2 that the homotopy class of $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{r}\mathcal{R}}_{F}$ is independent of the choice of $v$, so $\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$ is also independent of this choice.
+Similarly, Theorem 2 implies that for a Riesz continuous map $\mathcal{A}\colon X\to\mathcal{R}_{F}$ (in particular, for a norm continuous map $X\to\mathcal{B}_{F}$) the two definitions of the index given by (4.1) and (4.2) coincide.
 Recently N.
-Ivanov [$I_{1}$, $I_{2}$] developed a new approach to the analytical index of families.
-He introduced the notion of a “Fredholm family” and defined the $K^{0}(X)$-valued index, which we will denote by $\operatorname{Ind}$, for Fredholm families parametrized by points of a paracompact space $X$.
-See [$I_{2}$], Section 3.
-A Fredholm family is a family of Fredholm operators $\mathcal{A}_{x}$ such that the finite-dimensional subspaces $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$ and $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}^{*}_{x}|)$, as well as the restriction of $\mathcal{A}_{x}$ to $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$, depend continuously on $x$ for an appropriate locally chosen $\varepsilon>0$, without any additional continuity assumptions.
+Ivanov [I1, I2] developed a new approach to the family index.
+He introduced the notion of a “Fredholm family” and defined a $K^{0}(X)$-valued index, which we will denote by $\operatorname{Ind}$, for Fredholm families parametrized by points of a paracompact space $X$.
+See [I2], Section 3.
+A Fredholm family is a family of Fredholm operators $\mathcal{A}_{x}$ such that the finite-dimensional subspaces $\operatorname{\mathds{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$ and $\operatorname{\mathds{1}}_{[0,\varepsilon]}(|\mathcal{A}^{*}_{x}|)$, as well as the restriction of $\mathcal{A}_{x}$ to $\operatorname{\mathds{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$, depend continuously on $x$ for an appropriate locally chosen $\varepsilon>0$, without any additional continuity assumptions.
 Ivanov’s index $\operatorname{Ind}$ is defined both for operators acting on a fixed Hilbert space and on fibers of a Hilbert bundle; it is invariant under isomorphisms of Hilbert bundles and under passing to the bounded transform.
-Let $\mathcal{A}$ be a graph continuous family of Fredholm operators in $H$ parametrized by points of a paracompact space $X$.
-Then $\mathcal{A}$ is a Fredholm family in Ivanov’s sense, and we claim that our definition of the index of $\mathcal{A}$ provides the same element of $K^{0}(X)$ as Ivanov’s definition.
-Indeed, the properties of $\operatorname{Ind}$ mentioned above imply the equality $\operatorname{Ind}(\mathcal{A})=\operatorname{Ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$.
-Consider the norm continuous family $\alpha=\bm{\chi}\circ\Phi\circ\mathcal{A}$ of bounded Fredholm operators in $H$.
-By Theorem 7.2 of [$I_{2}$], $\operatorname{Ind}\alpha$ coincides with the index of $\alpha$ defined by Segal in [Se].
-The last index is equal to the homotopy class $[\alpha]$ of $\alpha$ in $[X,\mathcal{B}_{F}(H)]=K^{0}(X)$ by [Se], Theorem 5.1(b).
-Taking all this together, we obtain
-$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\alpha)=[\alpha]=\operatorname{Ind}(\alpha)=\operatorname{Ind}(\mathcal{A})\in K^{0}(X).$
-4 Polarizations
+By the construction, for a norm continuous map $\mathcal{A}\colon X\to\mathcal{B}_{F}$ Ivanov’s index coincides with the usual index $\operatorname{ind}(\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)$.
+Let now $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ be a graph continuous map from a paracompact space $X$.
+Then $\mathcal{A}$ is a Fredholm family in Ivanov’s sense, and the properties of $\operatorname{Ind}$ mentioned above imply the equality
+$\operatorname{Ind}(\mathcal{A})=\operatorname{Ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A}).$
+Therefore, our definition of the index of $\mathcal{A}$ provides the same element of $K^{0}(X)$ as Ivanov’s definition.
+Remark.
+Our proof of Theorem 1 shows that the definition of the index of $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ given by the formula $\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$ can be equivalently stated in terms of the graphs $\Gamma_{x}$ of operators $\mathcal{A}_{x}$, as the index of the family of projections $\Gamma_{x}\to H^{\prime}$.
+For this, one needs to trivialize the Hilbert bundle $\Gamma$, formed by subspaces $\Gamma_{x}\subset H\oplus H^{\prime}$, in the category of bundles with the structure group $\mathcal{U}$.
+Note that such a trivialization is not canonical; it is defined only up to a continuous map $X\to\mathcal{U}(H)$.
+However, change of a trivialization does not affect the homotopy type of the corresponding family of bounded Fredholm operators, since the unitary group $\mathcal{U}(H)$ is contractible.
+5 Polarizations and proof of Theorem 4
 Polarizations of a Hilbert space.
 A decomposition of a Hilbert space $H$ into the orthogonal sum $H=H^{+}\oplus H^{-}$ of two closed infinite-dimensional subspaces is called a polarization of $H$.
-It is determined by the subspace $H^{+}$ or equivalently by the orthogonal projection $p\in\mathcal{P}^{\star}(H)$ onto $H^{+}$; we will identify polarizations with the corresponding subspaces/projections.
+It is determined by the subspace $H^{+}\in\mathcal{G}r^{\star}(H)$ or equivalently by the orthogonal projection onto $H^{+}$.
+We denote by $\mathcal{P}^{\star}(H)\cong\mathcal{G}r^{\star}(H)$ the subspace of $\mathcal{P}(H)$ consisting of projections of infinite rank and corank and will identify polarizations with the corresponding subspaces/projections.
 There are two natural equivalence relations for polarizations of $H$, which we call “finite” and “compact”.
-Two polarizations $p$ and $q$ of $H$ are said to be compactly compatible if $p-q$ is a compact operator; they are said to be finitely compatible if $p-q$ is a finite rank operator (that is, the corresponding subspaces are commensurable).
-The compact restricted Grassmanian corresponding to the polarization given by a projection $p$ is defined as the space of all polarizations compatible with $p$:
-$\operatorname{Gr}_{K}=\operatorname{Gr}_{K}(H;p)=\left\{q\in\mathcal{P}^{\star}(H)\,|\,q-p\in\mathcal{K}(H)\right\}.$
-It is equipped with the norm topology induced from the natural inclusion $\operatorname{Gr}_{K}(H;p)\hookrightarrow\mathcal{P}(H)$, or equivalently the inclusion $\operatorname{Gr}_{K}(H;p)\hookrightarrow\mathcal{K}(H)$, $q\mapsto q-p$.
-The finite restricted Grassmanian $\operatorname{Gr}_{f}(H;p)$ is defined in a similar manner; it is a subspace of $\operatorname{Gr}_{K}(H;p)$.
+Two polarizations $p$ and $q$ of $H$ are said to be compactly compatible if pqp-q is a compact operator; they are said to be finitely compatible if pqp-q is a finite rank operator (that is, the corresponding subspaces are commensurable).
+The compact restricted Grassmanian corresponding to a polarization $p$ is defined as the space of all polarizations compatible with $p$:
+$\mathcal{G}r_{K}(H;p)=\left\{q\in(H)\mid q-p\in\mathcal{K}(H)\right\}.$
+It is equipped with the norm topology induced by the natural inclusion $\mathcal{G}r_{K}(H;p)\hookrightarrow\mathcal{G}r(H)\cong\mathcal{P}(H)$, or equivalently by the inclusion $\mathcal{G}r_{K}(H;p)\hookrightarrow\mathcal{K}(H)$, $q\mapsto q-p$.
+The finite restricted Grassmanian $\mathcal{G}r_{f}(H;p)$ is defined in a similar manner; it is a subspace of $\mathcal{G}r_{K}(H;p)$.
 The restricted Grassmanian depends only on the corresponding equivalence class of the projection $p$.
@@
 Such subbundles arise, for example, as spectral sections of graph continuous families of Fredholm operators.
-In general, a subbundle $\mathcal{E}$ of $\mathcal{H}$ is given by a family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ which is continuous in an appropriate sense.
+In general, a subbundle $\mathcal{E}$ of $\mathcal{H}$ is given by a family of closed subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ which is continuous in an appropriate sense.
 We will use the following notion of continuity: we say that $\mathcal{E}$ is a subbundle of $\mathcal{H}$ if $\mathcal{E}$ is the range of a bundle projection, that is, a continuous bundle map $p\colon\mathcal{H}\to\mathcal{H}$ such that $p_{x}\in\mathcal{P}(\mathcal{H}_{x})$ for every $x\in X$.
-In this case, $\mathcal{H}\ominus\mathcal{E}=\operatorname{Ran}(1-p)$ is also a subbundle.
-If both $\mathcal{E}$ and $\mathcal{H}\ominus\mathcal{E}$ are subbundles of $\mathcal{H}$ in the sense of [DD], then they are subbundles in our sense.
+In this case $\mathcal{H}\ominus\mathcal{E}=\operatorname{Ran}(1-p)$ is also a subbundle.
+Equivalently, for every $x\in X$, some (and then every) local trivialization of $\mathcal{H}$ over a neighborhood $Y$ of $x$ takes $p$ to a continuous map $Y\to{{}^{s}{\mathcal{P}}}(H)$, where ${{}^{s}{\mathcal{P}}}(H)$ denotes the space of projections equipped with the strong operator topology.
+In particular, subbundles of the trivial Hilbert bundle $H_{X}=X\times H$ are in one-to-one correspondence with continuous maps $X\to{{}^{s}{\mathcal{P}}}(H)$.
+A family $\mathcal{E}$ of closed subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ is a subbundle of $\mathcal{H}$ in our sense if and only if both $(\mathcal{E}_{x})$ and $(\mathcal{H}_{x}\ominus\mathcal{E}_{x})$ are continuous subfields of the field $(\mathcal{H}_{x})$ of Hilbert spaces in the sense of Dixmier and Douady [DD].
 See [DD], Proposition 16.
-Equivalently, for every $x\in X$, some (and then every) local trivialization of $\mathcal{H}$ over a neighborhood $X^{\prime}$ of $x$ takes $p$ to a continuous map $X^{\prime}\to{{}^{s}{\mathcal{P}}}(H)$, where ${{}^{s}{\mathcal{P}}}(H)$ denotes the space $\mathcal{P}(H)$ equipped with the strong operator topology.
-In particular, subbundles of the trivial Hilbert bundle $H_{X}=X\times H$ are in one-to-one correspondence with continuous maps $X\to{{}^{s}{\mathcal{P}}}(H)$.
-Two subbundles of $\mathcal{H}$ given by bundle projections $p$ and $q$ are said to be (compactly) compatible if $p-q$ is a norm continuous family of compact operators.
-They are said to be finitely compatible if additionally all operators $p_{x}-q_{x}$ have finite rank.
+We say that two subbundles of $\mathcal{H}$ given by bundle projections $p$ and $q$ are (compactly) compatible if pqp-q is a norm continuous family of compact operators; they are finitely compatible if additionally all operators $p_{x}-q_{x}$ have finite rank.
 Polarizations of a Hilbert bundle.
-We need only compact polarizations in this paper, so we will mostly use the term “polarization” without the adjective “compact”.
-The changes required to obtain the finite version are obvious.
-Locally, a polarization of $\mathcal{H}$ is determined by a decomposition of $\mathcal{H}$, over some open subset $X^{\prime}\subset X$, into the orthogonal sum $\mathcal{H}=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$ of two subbundles of infinite rank.
-Equivalently, it is determined by the subbundle $\mathcal{H}^{+}$ of $\left.\mathcal{H}\right|_{X^{\prime}}$, or by the bundle projection with the range $\mathcal{H}^{+}$.
-By definition, two compatible subbundles of $\mathcal{H}$ over $X^{\prime}$ determine the same polarization.
-Globally, a polarization of $\mathcal{H}$ is determined by an open covering $(X_{i})$ of $X$ and subbundles $\mathcal{H}^{+}_{i}$ of $\left.\mathcal{H}\right|_{X_{i}}$ such that $\mathcal{H}^{+}_{i}$ has infinite rank and corank and is compatible with $\mathcal{H}^{+}_{j}$ over $X_{i}\cap X_{j}$ for every $i,j$.
-Following Ivanov [$I_{2}$], we call such a collection $(X_{i},\mathcal{H}^{+}_{i})$ a polarization atlas.
-Two polarization atlases are said to be equivalent if their union is also a polarization atlas.
-A polarization of $\mathcal{H}$ is an equivalence class of polarization atlases.
-A polarization $\Pi$ of $\mathcal{H}$ determines the restricted Grassmanian bundle $\operatorname{Gr}_{K}(\mathcal{H};\Pi)$ in an obvious way.
+A natural forgetful functor from finite to compact polarizations preserves the notion of an adapted trivialization.
+Therefore, the “finite” analogues of Theorems 4 and 5 follow immediately from these theorems.
+Because of that, we will deal only with compact polarizations in the rest of the paper and will mostly omit the adjective “compact” before “polarization”.
+The changes required to obtain the finite version of our statements are obvious.
+Locally, a polarization of $\mathcal{H}$ is determined by a decomposition of $\mathcal{H}$, over some open subset $Y\subset X$, into the orthogonal sum $\mathcal{H}=\mathcal{H}^{+}\oplus\mathcal{H}^{-}$ of two subbundles of infinite rank.
+Equivalently, it is determined by the subbundle $\mathcal{H}^{+}$ of $\left.\mathcal{H}\right|_{Y}$ or by the bundle projection with the range $\mathcal{H}^{+}$.
+By the definition, two compatible subbundles of $\mathcal{H}$ over $Y$ determine the same local polarization.
+[…]
+A polarization $\Pi$ of $\mathcal{H}$ determines the restricted Grassmanian bundle $\mathcal{G}r_{K}(\mathcal{H};\Pi)$ in an obvious way.
 This is one of the motivations behind the notion of a polarization.
-Two polarizations of $\mathcal{H}$ are said to be homotopic if they are restrictions to $X\times\left\{0\right\}$ and $X\times\left\{1\right\}$ of some polarization of the Hilbert bundle ×$\mathcal{H}\times[0,1]\to X\times[0,1]$.
+Two polarizations of $\mathcal{H}$ are said to be homotopic if they are restrictions to $X\times\left\{0\right\}$ and $X\times\left\{1\right\}$ of some polarization of the Hilbert bundle $\mathcal{H}\times[0,1]\to X\times[0,1]$.
 Adapted trivializations.
-Following [$I_{2}$], we say that a trivialization $\mathcal{H}\to X\times H$ is adapted to a subbundle $\mathcal{E}$ of $\mathcal{H}$ if it takes the family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ to a continuous map $X\to\operatorname{Gr}(H)$.
-If a trivialization of $\mathcal{H}$ is adapted to $\mathcal{E}$, then it is adapted to every subbundle compatible with $\mathcal{E}$.
+A trivialization $\mathcal{H}\to X\times H$ is said to be adapted to a subbundle $\mathcal{E}$ of $\mathcal{H}$ if it takes the family of subspaces $\mathcal{E}_{x}\subset\mathcal{H}_{x}$ to a continuous map $X\to\mathcal{G}r(H)$, or equivalently takes the corresponding family of projections to a norm continuous map $X\to\mathcal{P}(H)$.
+If a trivialization of $\mathcal{H}$ is adapted to $\mathcal{E}$, then it is adapted to every subbundle compatible with $\mathcal{E}$, since the difference pqp-q of corresponding bundle projections is norm continuous in every trivialization.
 Let $\Pi$ be a polarization given by an atlas $(X_{i},\mathcal{E}_{i})$.
-We say that a trivialization of $\mathcal{H}$ is adapted to $\Pi$ if its restrictions to $X_{i}$ are adapted to $\mathcal{E}_{i}$.
+We say that a trivialization of $\mathcal{H}$ is adapted to $\Pi$ if its restriction to $X_{i}$ is adapted to $\mathcal{E}_{i}$ for every $i$.
 The notion of an adapted local trivialization is defined in an obvious way.
-We call a polarization $\Pi$ of $\mathcal{H}$ tame if it admits local adapted trivializations (that is, for every point $x\in X$ there is a local trivialization of $\mathcal{H}$ adapted to $\Pi$ over some neighborhood of $x$).
-The notion of a polarization introduced by Ivanov in the appendix to the third version of his preprint [$I_{2}$] is a particular case of a tame finite polarization in our terminology.
-Theorem 4.1.
-Let $\Pi$ be a tame polarization of a locally trivial Hilbert bundle $\mathcal{H}$ over a paracompact space $X$.
-Then $\Pi$ admits a global adapted trivialization.
-Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
-In particular, every two adapted trivializations are homotopic (as trivializations adapted to $\Pi$).
+We call a subbundle $\mathcal{E}$, resp.
+a polarization $\Pi$ of $\mathcal{H}$ tame if it admits local adapted trivializations (that is, for every point $x\in X$ there is a local trivialization of $\mathcal{H}$ adapted to $\mathcal{E}$, resp.
+$\Pi$ over some neighborhood of $x$).
+Trivializations of $H_{X}$ adapted to $E_{X}$.
+As a first step to the proof of Theorem 4, we describe trivializations of the trivial Hilbert bundle $H_{X}=X\times H$ adapted to its trivial subbundle $E_{X}=X\times E$, where $E$ is a fixed closed subspace of $H$.
+Trivializations of $H_{X}$ are given by continuous maps $f\colon X\to{{}^{s}{\mathcal{U}}}(H)$.
+Such a trivialization is adapted to $E_{X}$ if and only if the map $X\to\mathcal{G}r(H)$ taking $x$ to $f(x)E$ is continuous.
+Therefore, trivializations of $H_{X}$ adapted to $E_{X}$ are in one-to-one correspondence with continuous maps $X\to T_{E}$, where the space $T_{E}$ is obtained from ${{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that the map
+(5.1) $\iota_{E}\colon T_{E}\to\mathcal{G}r(H),\quad u\mapsto uE$
+becomes continuous. In other words, $T_{E}$ has the topology induced by the injection
+$T_{E}\hookrightarrow{{}^{s}{\mathcal{U}}}(H)\times\mathcal{G}r(H),\quad u\mapsto(u,uE).$
+Two maps $f,g\colon X\to T_{E}$ are homotopic if and only if the corresponding trivializations are homotopic in the class of trivializations adapted to $E_{X}$.
+Lemma 5.1.
+If $E\in\mathcal{G}r^{\star}(H)$, then $T_{E}$ is contractible.
 Proof.
-Let us describe the plan of the proof.
-For an arbitrary polarization $\Pi$ we define a fiber bundle $\mathcal{T}=\mathcal{T}_{\Pi}$ over $X$ whose (local or global) sections are exactly (local or global) trivializations of $\mathcal{H}$ adapted to $\Pi$.
-In particular, $\mathcal{T}$ admits local sections if and only if $\Pi$ is tame.
-We will show that in this case the bundle $\mathcal{T}$ is locally trivial and has a contractible fiber.
-Such a bundle, over a paracompact base space $X$, admits global sections, and every section over a closed subset $Y\subset X$ can be extended to a global section [DD, Lemma 4].
-Therefore, $\Pi$ admits global adapted trivializations, and every trivialization of $\mathcal{H}$ over $Y$ adapted to $\Pi$ can be extended to a global adapted trivialization.
-Let $H_{X}$ be the trivial Hilbert bundle over $X$ with the fiber $H$.
-Consider the fiber bundle $\mathcal{U}(\mathcal{H},H_{X})$ associated with $\mathcal{H}$ whose sections are Hilbert bundle isomorphisms $\mathcal{H}\to H_{X}$ (that is, trivializations of $\mathcal{H})$.
-Every two such isomorphisms differ by an automorphism of $H_{X}$, which is defined by a continuous map $X\to{{}^{s}{\mathcal{U}}}(H)$.
-Therefore, $\mathcal{U}(\mathcal{H},H_{X})$ is a locally trivial principal bundle over $X$ with the structure group ${{}^{s}{\mathcal{U}}}(H)$.
-Its fiber over $x$ is ${{}^{s}{\mathcal{U}}}(\mathcal{H}_{x},H)$, and ${{}^{s}{\mathcal{U}}}(H)$ acts on the fiber by the left multiplication.
-A trivialization $\tau\colon\mathcal{H}\to H_{X}$ takes $\mathcal{U}(\mathcal{H},H_{X})$ to the trivial bundle $\mathcal{U}(H_{X},H_{X})=X\times{{}^{s}{\mathcal{U}}}(H)$.
-The key ingredient of our proof is the bundle $\mathcal{T}\to X$ which we will now define.
-The total space $\mathcal{T}$ coincides with $\mathcal{U}(\mathcal{H},H_{X})$ as a set, and the projection to $X$ is the same, but we equip $\mathcal{T}$ with a finer topology determined by $\Pi$.
-Hence the identity map
-(4.1) $\iota\colon\mathcal{T}\to\mathcal{U}(\mathcal{H},H_{X})$
-will be continuous and, in particular, the projection $\mathcal{T}\to X$ will be continuous.
-Suppose first that the polarization $\Pi$ is defined by a subbundle $\mathcal{E}$ of $\mathcal{H}$, that is, the atlas of $\Pi$ has only one chart $(X,\mathcal{E})$.
-Then $\mathcal{E}$ determines the map
-(4.2) $\iota_{\mathcal{E}}\colon\mathcal{T}\to\operatorname{Gr}(H)\cong\mathcal{P}(H),\quad(x,u)\mapsto u(\mathcal{E}_{x})\;\;\text{for}\;\;x\in X\;\;\text{and}\;\;u\in\mathcal{U}(\mathcal{H}_{x},H).$
-We equip $\mathcal{T}=\mathcal{T}_{\mathcal{E}}$ with the weakest topology for which both maps (4.1) and (4.2) are continuous.
-If $\mathcal{E}$ is replaced by a compatible subbundle $\mathcal{F}$, then $\mathcal{T}_{\mathcal{F}}=\mathcal{T}_{\mathcal{E}}$.
-Indeed, in this case the difference $k_{x}=p_{x}-q_{x}$ of projections onto $\mathcal{F}_{x}$ and $\mathcal{E}_{x}$ is a norm continuous family of compact operators, so $\iota_{\mathcal{F}}$ differs from $\iota_{\mathcal{E}}$ by a continuous map
-$\mathcal{U}(\mathcal{H},H_{X})\to\mathcal{K}(H),\quad(x,u)\mapsto uk_{x}u^{*}.$
-Therefore, $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ induce the same refinement of the original topology on $\mathcal{U}(\mathcal{H},H_{X})$ and thus $\mathcal{T}_{\mathcal{E}}$ depends only on the equivalence class of compatible subbundles $\mathcal{E}$.
-This allows to define the bundle $\mathcal{T}_{\Pi}$ for an arbitrary polarization $\Pi$ of $\mathcal{H}$.
-More precisely, let $\Pi$ be given by an atlas $(X_{i},\mathcal{E}_{i})$, $i\in I$.
-Then the restrictions of $\mathcal{T}_{\mathcal{E}_{i}}$ and $\mathcal{T}_{\mathcal{E}_{j}}$ to $X_{i}\cap X_{j}$ coincide.
-We equip $\mathcal{T}_{\Pi}$ with the topology defined by the requirement that its restriction to $X_{i}$ coincides with $\mathcal{T}_{\mathcal{E}_{i}}$ for every $i\in I$.
-By the definition of $\mathcal{T}_{\Pi}$, its (local) sections are exactly (local) trivializations of $\mathcal{H}$ adapted to $\Pi$.
-In particular, $\mathcal{T}_{\Pi}$ admits local sections if $\Pi$ is tame.
-The following lemma provides a stronger result.
-Lemma 4.2.
-If $\Pi$ is tame, then the bundle $\mathcal{T}_{\Pi}$ is locally trivial.
-Proof.
-Since $\Pi$ is tame, each point $x_{0}\in X$ has a neighborhood $X^{\prime}$ such that some trivialization of $\mathcal{H}$ over $X^{\prime}$ is adapted to $\Pi$.
-We can suppose that $X^{\prime}$ is contained in some $X_{i}$; then this trivialization takes $\mathcal{E}_{i}$ to the subbundle $\mathcal{E}^{\prime}$ given by a continuous map $\mathcal{E}^{\prime}\colon X^{\prime}\to\operatorname{Gr}^{\star}(H)$.
-Let $g\colon\operatorname{Gr}^{\star}\to\mathcal{U}(H)$ be a map given by (2.1) and $\bar{g}_{L}=g_{L}^{-1}$.
-Then the composition of $\mathcal{E}^{\prime}$ with $\bar{g}\colon\operatorname{Gr}^{\star}\to\mathcal{U}(H)$ provides the automorphism $G$ of the trivial Hilbert bundle $H_{X^{\prime}}$ which takes $\mathcal{E}^{\prime}_{x}$ to the subspace $E$ of $H$ independent of $x$.
-We can then replace the original adapted trivialization by its composition with $G$.
-The new trivialization takes $\mathcal{E}_{i}$ to the subbundle $E_{X^{\prime}}$ of $H_{X^{\prime}}$ and thus provides a bundle isomorphism $\left.\mathcal{T}\right|_{X^{\prime}}\to\mathcal{T}^{\prime}$, where the bundle $\mathcal{T}^{\prime}$ is build (in the way described above) from the trivial bundle $H_{X^{\prime}}$ and its subbundle $E_{X^{\prime}}$.
-We will prove that the bundle $\mathcal{T}^{\prime}$ is trivial and find its fiber.
-By the definition, the total space $\mathcal{T}^{\prime}$ is obtained from the product $X^{\prime}\times{{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that the map $\mathcal{T}^{\prime}\to\operatorname{Gr}^{\star}(H)$, $(x,u)\mapsto uE$ becomes continuous.
-This map does not actually depend on $x$.
-Therefore, $\mathcal{T}^{\prime}$ coincides with the trivial bundle $X^{\prime}\times T\to X^{\prime}$, where the fiber $T$ is obtained from ${{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that
-(4.3) $\iota_{E}\colon T\to\operatorname{Gr}^{\star}(H),\quad u\mapsto uE$
-becomes continuous. In other words, $T$ has the topology induced by the inclusion
-$T\hookrightarrow{{}^{s}{\mathcal{U}}}(H)\times\operatorname{Gr}^{\star}(H),\quad u\mapsto(u,uE).$
-Hence the restriction of $\mathcal{T}_{\Pi}$ to a neighborhood of an arbitrary point $x_{0}$ is a trivial bundle with the fiber $T$. $\square$
-Lemma 4.3.
-The space $T$ defined above is contractible.
-Proof.
-The fiber of $\iota_{E}$ over $E$ is the subspace of ${{}^{s}{\mathcal{U}}}(H)$ consisting of unitary operators preserving $E$.
+For $E\in\mathcal{G}r^{\star}$ the map (5.1) takes $T_{E}$ to $\mathcal{G}r^{\star}$.
+We will show that $\iota_{E}\colon T_{E}\to\mathcal{G}r^{\star}$ is a trivial bundle and construct its trivialization.
+The base space $\mathcal{G}r^{\star}$ has a distinguished point $E\in\mathcal{G}r^{\star}$.
+The fiber of $\iota_{E}$ over this point is the subspace of ${{}^{s}{\mathcal{U}}}(H)$ consisting of unitary operators $u\colon H\to H$ preserving $E\subset H$.
 Denote this subspace by ${{}^{s}{\mathcal{U}}}_{E}$.
-We will show that (4.3) is a trivial bundle over $\operatorname{Gr}^{\star}$ and construct its trivialization $\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}\to T$.
-Let $g\colon\operatorname{Gr}^{\star}(H)\to\mathcal{U}(H)$ be a continuous map such that $L=g_{L}E$ for all $L\in\operatorname{Gr}^{\star}(H)$, as in (2.1). We define $\varphi$ by the formula
-$\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}(H)\;\to\;T,\quad\varphi(v,L)=g_{L}\cdot v.$
-Since
-(4.4) $(g_{L}v)E=g_{L}E=L\;\;\text{for every}\;\;v\in{{}^{s}{\mathcal{U}}}_{E},$
-the composition $\iota_{E}\circ\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}\to\operatorname{Gr}^{\star}$ is the projection of the product on the second factor; in particular, it is continuous.
-The composition of $\varphi$ with the identity map $T\to{{}^{s}{\mathcal{U}}}$ is given by $(v,L)\mapsto g_{L}v$ and is also continuous.
+Let $g\colon\mathcal{G}r^{\star}\to\mathcal{U}(H)$ be a continuous map such that $L=g_{L}E$ for all $L\in\mathcal{G}r^{\star}$, as in (2.1).
+Consider the map
+$\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star}\;\to\;T_{E},\quad\varphi(v,L)=g_{L}\cdot v.$
+The composition $\iota_{E}\circ\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star}\to\mathcal{G}r^{\star}$ is the projection of the product on the second factor, as can be seen from
+(5.2) $(g_{L}v)E=g_{L}E=L\;\;\text{for every}\;\;v\in{{}^{s}{\mathcal{U}}}_{E}.$
+In particular, $\iota_{E}\circ\varphi$ is continuous.
+The composition of $\varphi$ with the identity map $T_{E}\to{{}^{s}{\mathcal{U}}}$ is given by $(v,L)\mapsto g_{L}v$ and is also continuous.
 Therefore, $\varphi$ itself is continuous.
@@
 Since $vE=g_{L}^{-1}uE=g_{L}^{-1}L=E$, the unitary operator $v=g_{L}^{-1}u$ belongs to ${{}^{s}{\mathcal{U}}}_{E}$, and we obtain the continuous map
-$\psi\colon T\to{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}(H),\quad\psi(u)=\left(g_{uE}^{-1}\cdot u,\,uE\right).$
-By the construction, $\psi\circ\varphi(v,L)=(v,L)$ for every $(v,L)\in{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}$. Conversely,
-$\varphi\circ\psi(u)=\varphi\left(g_{uE}^{-1}\cdot u,\,uE\right)=g_{uE}\cdot g_{uE}^{-1}\cdot u=u\;\;\text{for every}\;\;u\in T.$
+$\psi\colon T_{E}\to{{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star},\quad\psi(u)=\left(g_{uE}^{-1}\cdot u,\,uE\right).$
+By the construction, $\psi\circ\varphi(v,L)=(v,L)$ for every $(v,L)\in{{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star}$. Conversely,
+$\varphi\circ\psi(u)=\varphi\left(g_{uE}^{-1}\cdot u,\,uE\right)=g_{uE}\cdot g_{uE}^{-1}\cdot u=u\;\;\text{for every}\;\;u\in T_{E}.$
 Therefore, $\psi$ is inverse to $\varphi$ and thus $\varphi$ is a homeomorphism.
 The domain of $\varphi$ is the product of two factors.
-The second factor $\operatorname{Gr}^{\star}$ is contractible, see [AS, proof of Lemma 3.6].
+The second factor $\mathcal{G}r^{\star}$ is contractible, see [AS, proof of Lemma 3.6].
 The first factor ${{}^{s}{\mathcal{U}}}_{E}={{}^{s}{\mathcal{U}}}(E)\times{{}^{s}{\mathcal{U}}}(E^{\bot})$ is the product of two copies of $^{s}{\mathcal{U}}$}.
 By [DD, Lemma 3], $^{s}{\mathcal{U}}$} is contractible, so ${{}^{s}{\mathcal{U}}}_{E}$ is also contractible.
-It follows that the product ${{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}$ is contractible.
-This completes the proof of the lemma.
+Therefore, the product ${{}^{s}{\mathcal{U}}}_{E}\times\mathcal{G}r^{\star}\cong T_{E}$ is contractible.
 $\square$
-End of the proof of Theorem 4.1.
-Lemmas 4.2 and 4.3 together with the reasoning in the beginning of the proof show that every adapted trivialization of $\mathcal{H}$ over a closed subspace $Y\subset X$ can be extended to a global adapted trivialization.
-If $\tau_{0}$ and $\tau_{1}$ are trivializations of $\mathcal{H}$ adapted to $\Pi$, then $\tau_{0}\sqcup\tau_{1}$ is a trivialization of $\tilde{\mathcal{H}}$ over $X\times\left\{0,1\right\}$ adapted to $\tilde{\Pi}$, where $\tilde{\mathcal{H}}$ and $\tilde{\Pi}$ are the liftings of $\mathcal{H}$ and $\Pi$ to $X\times[0,1]$.
-The polarization $\tilde{\Pi}$ is tame, so $\tau_{0}\sqcup\tau_{1}$ can be extended from $X\times\left\{0,1\right\}$ to a trivialization of $\tilde{\mathcal{H}}$ adapted to $\tilde{\Pi}$.
-This provides a homotopy between $\tau_{0}$ and $\tau_{1}$ in the class of trivializations adapted to $\Pi$ and completes the proof of the theorem.
+Trivializations of $\mathcal{H}$.
+Let $\mathcal{H}$ be a locally trivial Hilbert bundle and $H_{X}$ be the trivial Hilbert bundle with the fiber $H$ over a topological space $X$.
+Let $\mathcal{U}(\mathcal{H},H_{X})$ be the fiber bundle associated with $\mathcal{H}$ whose fiber over $x$ is ${{}^{s}{\mathcal{U}}}(\mathcal{H}_{x},H)$.
+This is a locally trivial principal bundle over $X$ with the structure group ${{}^{s}{\mathcal{U}}}(H)$ acting on fibers by the left multiplication.
+Sections of $\mathcal{U}(\mathcal{H},H_{X})$ are Hilbert bundle isomorphisms $\mathcal{H}\to H_{X}$, that is, trivializations of $\mathcal{H}$.
+A trivialization $\mathcal{H}\to H_{X}$ takes $\mathcal{U}(\mathcal{H},H_{X})$ to the trivial bundle $\mathcal{U}(H_{X},H_{X})=X\times{{}^{s}{\mathcal{U}}}(H)$.
+Trivializations adapted to $\mathcal{E}$.
+Let $\mathcal{E}$ be a subbundle of $\mathcal{H}$.
+Consider the space $\mathcal{T}=\mathcal{T}_{\mathcal{E}}$ which coincides with $\mathcal{U}(\mathcal{H},H_{X})$ as a set, but has finer topology determined by $\mathcal{E}$.
+Namely, we equip $\mathcal{T}$ with the weakest topology such that both the identity map $\iota\colon\mathcal{T}\to\mathcal{U}(\mathcal{H},H_{X})$ and the map
+(5.3) $\iota_{\mathcal{E}}\colon\mathcal{T}\to\mathcal{G}r(H)\cong\mathcal{P}(H),\quad\iota_{\mathcal{E}}(x,u)\ =u(\mathcal{E}_{x})\;\;\text{for}\;\;x\in X\;\;\text{and}\;\;u\in\mathcal{U}(\mathcal{H}_{x},H)$
+are continuous.
+Then the composition $\mathcal{T}\to\mathcal{U}(\mathcal{H},H_{X})\to X$ is also continuous, and we consider $\mathcal{T}$ as a space over $X$.
+By the construction, sections of $\mathcal{T}\to X$ are in one-to-one correspondence with trivializations of $\mathcal{H}$ adapted to $\mathcal{E}$.
+Lemma 5.2.
+If $\mathcal{E}$ and $\mathcal{F}$ are compatible subbundles of $\mathcal{H}$, then $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ induce the same refinement of the original topology on $\mathcal{U}(\mathcal{H},H_{X})$.
+In other words, $\mathcal{T}_{\mathcal{E}}$ depends only on the equivalence class of compatible subbundles $\mathcal{E}$.
+Proof.
+Let $p_{x}$ and $q_{x}$ be projections onto $\mathcal{E}_{x}$ and $\mathcal{F}_{x}$ respectively.
+By the definition of compatible subbundles, the difference $k_{x}=q_{x}-p_{x}$ is a norm continuous family of compact operators, so $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ considered as maps from $\mathcal{U}(\mathcal{H},H_{X})$ to $\mathcal{P}(H)$ differ by a continuous map
+$\mathcal{U}(\mathcal{H},H_{X})\to\mathcal{K}(H),\quad(x,u)\mapsto\iota_{\mathcal{F}}(x,u)-\iota_{\mathcal{E}}(x,u)=uq_{x}u^{*}-up_{x}u^{*}=uk_{x}u^{*}.$
+Hence $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ induce the same refinement of the original topology on $\mathcal{U}(\mathcal{H},H_{X})$. $\square$
+Lemma 5.3.
+If $\mathcal{E}$ is tame, then $\mathcal{T}_{\mathcal{E}}\to X$ is a locally trivial bundle, whose fiber $T=T_{E}$ is the space defined before Lemma 5.1.
+Proof.
+Let $x_{0}\in X$.
+Since $\mathcal{E}$ is tame, there is a trivialization $\tau^{\prime}$ of $\mathcal{H}$ adapted to $\mathcal{E}$ over a neighborhood $X^{\prime}$ of $x_{0}$.
+This trivialization takes $\mathcal{E}$ to the subbundle $\mathcal{E}^{\prime}$ given by a continuous map $X^{\prime}\to\mathcal{G}r(H)$.
+Replacing $X^{\prime}$ by a smaller neighborhood if necessary, we can find a map $g\colon X^{\prime}\to\mathcal{U}(H)$ such that $g_{x}\mathcal{E}^{\prime}_{x}=E$, where $E=\mathcal{E}^{\prime}_{x_{0}}$.
+The composition of $\tau^{\prime}$ with the automorphism of the trivial Hilbert bundle $H_{X^{\prime}}$ provided by $g$ determines the new trivialization $\tau$ which takes $\mathcal{E}_{x}$ to the subspace $E$ of $H$ independent of $x$ and thus takes $\mathcal{E}$ to the constant subbundle $E_{X^{\prime}}$ of $H_{X^{\prime}}$.
+Such a trivialization $\tau$ provides a bundle isomorphism $\left.\mathcal{T}_{\mathcal{E}}\right|_{X^{\prime}}\to\mathcal{T}^{\prime}$, where the bundle $\mathcal{T}^{\prime}$ is build (in the way described above) from the trivial bundle $H_{X^{\prime}}$ and its subbundle $E_{X^{\prime}}$.
+By the definition, the total space $\mathcal{T}^{\prime}$ is obtained from the product $X^{\prime}\times{{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that the map $\mathcal{T}^{\prime}\to\mathcal{G}r(H)$, $(x,u)\mapsto uE$ becomes continuous.
+But this map does not actually depend on $x$, so $\mathcal{T}^{\prime}$ coincides with the trivial bundle $X^{\prime}\times T_{E}\to X^{\prime}$.
 $\square$
-Finite covering dimension.   The proof of Theorem E in the next section is based on the following result.
+Trivializations adapted to $\Pi$.
+Let $\Pi$ be a polarization of $\mathcal{H}$ given by an atlas $(X_{i},\mathcal{E}_{i})$, $i\in I$.
+Consider the space $\mathcal{T}_{\Pi}$ which coincides with $\mathcal{U}(\mathcal{H},H_{X})$ as a set and is equipped with the weakest topology such that the identity map
+(5.4) $\left.\mathcal{T}_{\Pi}\right|_{X_{i}}\to\mathcal{T}_{\mathcal{E}_{i}}$
+is continuous for every $i\in I$.
+By Lemma 5.2 the restrictions of $\mathcal{T}_{\mathcal{E}_{i}}$ and $\mathcal{T}_{\mathcal{E}_{j}}$ to $X_{i}\cap X_{j}$ coincide, so (5.4) is actually a homeomorphism.
+Moreover, this definition does not depend on the choice of an atlas of $\Pi$.
+The composition $\mathcal{T}_{\Pi}\to\mathcal{U}(\mathcal{H},H_{X})\to X$ is continuous, and we consider $\mathcal{T}_{\Pi}$ as a space over $X$.
+By the construction, (local or global) sections of $\mathcal{T}_{\Pi}\to X$ are exactly (local or global) trivializations of $\mathcal{H}$ adapted to $\Pi$.
+In particular, $\mathcal{T}_{\Pi}\to X$ admits local sections if and only if $\Pi$ is tame.
+Proof of Theorem 4.
+Let $\Pi$ be a tame polarization of a locally trivial Hilbert bundle $\mathcal{H}$ over $X$.
+Then, by Lemma 5.3, $\mathcal{T}_{\Pi}\to X$ is a locally trivial bundle with the fiber $T_{E}$, $E\in\mathcal{G}r^{\star}(H)$.
+This fiber is contractible by Lemma 5.1.
+If additionally $X$ is paracompact, then $\mathcal{T}_{\Pi}\to X$ admits global sections, every two sections are homotopic, and a section over a closed subspace $Y\subset X$ can be extended to a global section [DD, Lemma 4].
+Since sections of $\mathcal{T}_{\Pi}$ correspond to adapted trivializations, this proves the theorem.
+$\square$
+6 Proof of Theorem 5
+Let $\dim Z$ denote the covering dimension of a paracompact space $Z$.
+Recall that a closed subset $Y$ of a paracompact space $Z$ is itself paracompact [E, Corollary 5.1.29], normal [E, Theorem 5.1.5], and $\dim Y\leqslant\dim Z$ [E, Theorem 7.1.8].
+Finite covering dimension.
+Our proof of Theorem 5 is based on Theorem 6.1 below.
+Theorem 6.1 follows from a theorem of Dixmier and Douady [DD, Theorem 5].
+For the convenience of the reader, we include here a proof of Theorem 6.1; it is based on the same ideas as the proof of [DD, Theorem 5], but is stated in terms of operators in a fixed Hilbert space instead of “continuous fields of Hilbert spaces” of Dixmier–Douady.
+Let $E$ be a subspace of $H$ of infinite dimension and codimension and $p_{E}\in\mathcal{P}^{\star}(H)$ be the orthogonal projection onto $E$.
+In terms of projections, the map $\mathcal{U}(H)\to\mathcal{G}r^{\star}(H)$, $u\mapsto uE$ is written as $\mathcal{U}(H)\to\mathcal{P}^{\star}(H)$, $u\mapsto up_{E}u^{-1}$.
+This map has a global section, as explained in the beginning of Section 2.
+In contrast with this, the continuous map
+${{}^{s}{\mathcal{U}}}(H)\to{{}^{s}{\mathcal{P}}}^{\star}(H)$
+given by the same formula $u\mapsto up_{E}u^{-1}$ has no sections, even locally. However, the following result holds.
 Theorem 6.1.
-Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a metrizable space $X$ of locally finite covering dimension.
-Then every polarization $\Pi$ of $\mathcal{H}$ is tame and thus satisfies all the conclusions of Theorem 4.1.
-Proof.
-Every subspace $X^{\prime}$ of $X$ is metrizable and thus paracompact, and has locally finite covering dimension by [E, Theorem 7.3.1].
-By [DD, Theorem 5], every separable Hilbert bundle over $X^{\prime}$ of infinite rank is locally trivial.
-The bundle $\mathcal{H}$ is trivial by [DD, Theorem 1] and has separable fiber, so it is separable in the sense of [DD], and so are its subbundles over $X^{\prime}$.
-Therefore, every infinite rank subbundle of $\mathcal{H}$ over $X^{\prime}$ is locally trivial.
-Let $(X_{i},\mathcal{E}_{i})$ be an atlas of $\Pi$.
-By the reasoning above, both $\mathcal{E}_{i}$ and $\mathcal{E}_{i}^{\bot}$ are locally trivial.
-The direct sum of their local trivializations provides a local trivialization of $\mathcal{H}$ which takes $\mathcal{E}_{i}$ to a constant subspace and thus is adapted to $\mathcal{E}_{i}$.
-Therefore, $\Pi$ is tame.
-Since $X$ is paracompact, $\Pi$ satisfies the assumptions of Theorem 4.1.
+Let $X$ be a paracompact space of finite covering dimension.
+Then every continuous map $p\colon X\to{{}^{s}{\mathcal{P}}}^{\star}(H)$ can be lifted to a map $u\colon X\to{{}^{s}{\mathcal{U}}}(H)$.
+In other words, every subbundle of $H_{X}$ with infinite rank and corank is trivial.
+Proof. Let $S(H)$ be the unit sphere in $H$. For $q\in\mathcal{P}(H)$ and $F=\operatorname{Ran}q$ we denote the unit sphere in $F$ by $S(q)=S(F)$.
+The map $p\colon X\to{{}^{s}{\mathcal{P}}}^{\star}(H)$ determines a bundle projection $H_{X}\to H_{X}$ of the trivial bundle $H_{X}$, which we also denote by $p$.
+Let $\mathcal{E}$ be the range of $p$ (it is a subbundle of $H_{X}$) and
+$S(p)=S(\mathcal{E}):=(X\times S(H))\cap\mathcal{E}=\left\{(x,e)\mid x\in X,\;e\in S(\mathcal{E}_{x})\right\}$
+be the bundle of unit spheres in $\mathcal{E}$.
+Lemma 6.2.
+Let $\eta$ be a continuous section of $\mathcal{E}\to X$ and $\varepsilon>0$.
+Then there is a continuous section $\xi$ of $S(\mathcal{E})\to X$ such that the distance from $\eta(x)$ to the linear span $\left\langle\xi(x)\right\rangle$ of $\xi(x)$ is less than $\varepsilon$ for every $x\in X$.
+Proof. For a non-zero vector $v\in H$, we will denote by $\mathbf{s}(v)=\lVert v\rVert^{-1}\cdot v\in S(H)$ the corresponding unit vector.
+Consider the family $\mathcal{S}=\left\{S(q)\mid q\in\mathcal{P}^{\star}(H)\right\}$ of closed subsets of $H$.
+Let us check that the composition $X\to{{}^{s}{\mathcal{P}}}^{\star}\to\mathcal{S}$ taking $x\in X$ to $S(p_{x})\in\mathcal{S}$ satisfies all four assumptions of Michael’s Selection Theorem [M, Theorem 1.2].
+(1) $X$ is paracompact and has finite covering dimension.
+(2) Every element of the family $\mathcal{S}$ is contractible. Indeed, the unit spheres $S(q)$ are contractible by Kakutani’s theorem [Kak, Theorem 3].
+(3) Let us show that the family $\mathcal{S}$ is equi-$LC^{n}$ in terms of [M] for every $n$.
+In fact, the stronger condition holds: for every open ball $V$ of radius <1<1 with the center in $v_{0}\in S(H)$ and for every $q\in\mathcal{P}^{\star}$, the intersection $V\cap S(q)$ is either empty or contractible.
+Indeed, if $V\cap S(q)$ is non-empty, then it coincides with $V^{\prime}\cap S(q)$, where $V^{\prime}$ is an open ball of the same or smaller radius with the center in $v^{\prime}_{0}=\mathbf{s}(qv_{0})\in S(q)$.
+The formula $(v,t)\mapsto\mathbf{s}((1-t)v+tv^{\prime}_{0})$ provides a contraction of $V^{\prime}\cap S(q)$.
+Therefore, $\mathcal{S}$ is equi-$LC^{n}$.
+(4) Let us show that the map ${{}^{s}{\mathcal{P}}}^{\star}(H)\to\mathcal{S}$, $q\mapsto S(q)$ is lower semi-continuous, that is, for every open $V\subset H$ the set $U=\left\{q\in\mathcal{P}^{\star}\mid S(q)\cap V\neq\emptyset\right\}$ is open in $^{s}{\mathcal{P}}$$^{\star}$.
+Let $q_{0}\in U$ and $v_{0}\in S(q_{0})\cap V$.
+Then $W=\left\{q\in\mathcal{P}^{\star}\mid\lVert(q-q_{0})v_{0}\rVert<1\right\}$ is a neighborhood of $q_{0}$ in $^{s}{\mathcal{P}}$$^{\star}$ and the map $f\colon W\to H$, $f(q)=\mathbf{s}(qv_{0})\in S(q)$ is continuous and has the prescribed value $v_{0}$ at $q_{0}$.
+Thus the inverse image $f^{-1}(V)\subset W$ is open in $^{s}{\mathcal{P}}$$^{\star}$, contains $q_{0}$, and is contained in $U$.
+Therefore, $U$ is open in $^{s}{\mathcal{P}}$$^{\star}$.
+We see that the map $X\to\mathcal{S}$, $x\mapsto S(p_{x})\in\mathcal{S}$ satisfies assumptions of [M, Theorem 1.2] and thus every section of $S(\mathcal{E})$ over a closed subset $Y\subset X$ can be extended to a global section of $S(\mathcal{E})$.
+Let $Y=\left\{x\in X\mid\lVert\eta(x)\rVert\geqslant\varepsilon\right\}$.
+Then the vector field $\xi\colon Y\to S(H)$, $\xi(x)=\mathbf{s}(\eta(x))$ is a section of $S(\mathcal{E})$ over $Y$.
+Let $\xi$ be an extension of this section to the whole $X$.
+By the construction, the distance from $\eta(x)$ to $\left\langle\xi(x)\right\rangle$ is zero for $x\in Y$ and less than $\varepsilon$ for $x\notin Y$, so $\xi$ is a required section.
 $\square$
-Remark.
+Continuation of the proof of Theorem 6.1.
+We aim to construct continuous maps $\xi_{n}\colon X\to S(H)$, $n\in\mathbb{N}$ such that $\xi_{1}(x),\xi_{2}(x),\ldots$ is an orthonormal basis of $\mathcal{E}_{x}$ for every $x\in X$.
+To this end, fix a countable dense subset $\bar{\eta}_{1},\bar{\eta}_{2},\ldots$ of $H$.
+Then $p\bar{\eta}_{n}$ are sections of $\mathcal{E}\to X$ and the set $\left\{p(x)\bar{\eta}_{n}\right\}_{n\in\mathbb{N}}$ is dense in $\mathcal{E}_{x}$ for every $x$.
+We will now construct bundle projections $p_{1}\geqslant p_{2}\geqslant\ldots$ of $H_{X}$ and sections $\xi_{n}$ of $S(p_{n})$ by induction as follows.
+We start with $p_{1}=p$.
+Suppose that $p_{n}$ is already constructed.
+Applying the lemma above to $\varepsilon_{n}=1/n$, $p_{n}$, and $\eta_{n}=p_{n}\bar{\eta}_{n}$, we obtain a continuous section $\xi_{n}\colon X\to S(p_{n})$ such that the distance from $\eta_{n}(x)$ to $\left\langle\xi_{n}(x)\right\rangle$ is less than $\varepsilon_{n}$ for every $x\in X$.
+Let $q_{n}(x)$ be the orthogonal projection onto $\left\langle\xi_{n}(x)\right\rangle$.
+Then the map $q_{n}\colon X\to\mathcal{P}(H)$ is norm continuous, and we proceed with the next bundle projection $p_{n+1}=p_{n}-q_{n}$.
+In such a way, we obtain continuous sections $\xi_{1},\xi_{2},\ldots$ of $S(\mathcal{E})$ such that the vectors $\xi_{n}(x)$ are pairwise orthogonal and their finite linear combinations are dense in $\mathcal{E}_{x}$ for every $x\in X$.
+The same reasoning applied to 1p1-p instead of $p$ provides continuous sections $\xi^{\prime}_{1},\xi^{\prime}_{2},\ldots$ of $S(H_{X}\ominus\mathcal{E})$ such that for every $x$ the vectors $\xi^{\prime}_{n}(x)$ form an orthonormal basis of $H\ominus\mathcal{E}_{x}$.
+Fix an orthonormal basis $e_{1},e_{2},\ldots$ of $E$ and an orthonormal basis $e^{\prime}_{1},e^{\prime}_{2},\ldots$ of $H\ominus E$.
+Let $u(x)\colon H\to H$ be the isometry given by the rule $u(x)e_{n}=\xi_{n}(x)$ and $u(x)e^{\prime}_{n}=\xi^{\prime}_{n}(x)$.
+Then the range of $u(x)$ is dense in $H$, so $u(x)$ is unitary.
+By the construction, $x\mapsto u(x)$ is strongly continuous and $u(x)E=\mathcal{E}_{x}$, so the map $u\colon X\to{{}^{s}{\mathcal{U}}}(H)$ lifts $p\colon X\to{{}^{s}{\mathcal{P}}}(H)$.
+The trivialization of the trivial Hilbert bundle $H_{X}$ given by $u^{-1}$ takes $\mathcal{E}$ to the constant subbundle $E_{X}$ of $H_{X}$ and thus is adapted to $\mathcal{E}$. $\square$
+Proof of Theorem 5.
+The structure group $^{s}{\mathcal{U}}$} of a locally trivial Hilbert bundle $\mathcal{H}$ is contractible [DD, Lemma 3].
+Since the base space $X$ is paracompact, $\mathcal{H}$ is trivial.
+Let $(X_{i},\mathcal{E}_{i})$ be an atlas of a polarization $\Pi$.
+Let $x\in X_{i}$ be an arbitrary point.
+By the assumption of the theorem, $x$ has a neighborhood $X^{\prime}$ of finite covering dimension.
+Since $X$ is normal, there is a closed neighborhood $Y$ of $x$ which is contained in $X^{\prime}\cap X_{i}$.
+Such an $Y$ is a closed subset of a paracompact space $X$, so it is also paracompact; it is a closed subset of $X^{\prime}$, so $\dim Y\leqslant\dim X^{\prime}<\infty$.
+By Theorem 6.1, the restriction of $\mathcal{E}_{i}$ to $Y$ is trivial.
+Since $x$ was chosen arbitrarily, each $\mathcal{E}_{i}$ is tame.
+Therefore, $\Pi$ is also tame, so it satisfies the assumptions of Theorem 4 and thus admits a global adapted trivialization.
+$\square$
+Remark 6.3.
 For a general space $X$, such a result is no longer true.
-Indeed, Dixmier and Douady constructed a subbundle $\mathcal{E}$ of the trivial Hilbert bundle $\mathcal{H}$ over a compact space $X$, which is not locally trivial [DD, Theorem 6].
+Indeed, Dixmier and Douady constructed a subbundle $\mathcal{E}$ of the trivial Hilbert bundle $\mathcal{H}$ over a compact space $X$, which is not locally trivial.
 We recall their construction in Example 7.5 below.
@@
 7 Self-adjoint operators
-Adapted trivializations.
-We say that a (local or global) trivialization of $\mathcal{H}$ is adapted to a family $\mathcal{A}$ of regular operators $\mathcal{A}_{x}\in\mathcal{R}(\mathcal{H}_{x})$ if it takes $\mathcal{A}$ to a Riesz continuous family of operators in $H$.
-If operators $\mathcal{A}_{x}$ are bounded (say, they are projections), then an adapted trivialization takes $\mathcal{A}$ to a norm continuous family.
-So the definition from the previous section, of a trivialization adapted to a subbundle, agrees with the present definition.
-Bounded operators.
-Let $\mathcal{A}$ be a family of self-adjoint bounded operators acting on fibers of a Hilbert bundle $\mathcal{H}$.
-We want to know whether $\mathcal{A}$ admits a global adapted trivialization.
-Clearly, for that to be true, $\mathcal{A}$ has to be norm quasi-continuous (that is, admit local adapted trivializations).
-The following theorem shows that for essentially unitary operators this local condition is also sufficient.
-The author does not know whether this is true for arbitrary bounded self-adjoint operators.
-Theorem 5.1.
-Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$ and $\mathcal{A}$ be a family of self-adjoint essentially unitary operators.
-Suppose that $\mathcal{A}$ is norm quasi-continuous.
-Then there exists a global trivialization of $\mathcal{H}$ adapted to $\mathcal{A}$.
-Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization.
-In particular, every two such trivializations are homotopic (in the space of trivializations adapted to $\mathcal{A}$).
-Proof.
+Proof of Theorem 6.
+A regular operator has compact resolvents if and only if its bounded transform is essentially unitary.
+A trivialization is adapted to a family $\mathcal{A}$ of regular operators if and only if it is adapted to the bounded transform $\bm{\chi}(\mathcal{A})$.
+Therefore, we only need to prove Theorem 6 for essentially unitary operators $\mathcal{A}_{x}$; the case of operators with compact resolvents will immediately follow.
 Let $X^{+}$, resp.
@@
 negative, and let $X^{\star}=X\setminus(X^{+}\cup X^{-})$.
-Since $\mathcal{A}$ is norm quasi-continuous, each of the three subspaces $X^{+}$, $X^{-}$, and $X^{\star}$ is open and thus closed in $X$, so it is enough to prove the theorem for each of these subspaces separately.
+In a local adapted trivialization $\mathcal{A}$ is norm continuous, so each of the three subspaces $^{+}, XX^{-}$, and $X^{\star}$ is open in $X$.
+Therefore, $X$ is the disjoint union of $^{+}, XX^{-}$, and $X^{\star}$, so it is enough to prove the theorem for each of these subspaces separately.
 For $x\in X^{+}$, the operator $\mathcal{A}_{x}$ is essentially positive and essentially unitary, so $\mathcal{A}_{x}-1$ is compact.
-Over $X^{+}$ the family $x\mapsto\mathcal{A}_{x}-1$ is a norm quasi-continuous family of compact operators, so both $\mathcal{A}-1$ and $\mathcal{A}$ are norm continuous in every trivialization of $\mathcal{H}$ over $X^{+}$.
+Over $X^{+}$ the family $x\mapsto\mathcal{A}_{x}-1$ is a norm continuous family of compact operators, so both $\mathcal{A}-1$ and $\mathcal{A}$ are norm continuous in every trivialization of $\mathcal{H}$ over $X^{+}$.
 The similar reasoning shows that $\mathcal{A}$ is norm continuous in every trivialization of $\mathcal{H}$ over $X^{-}$.
@@
 It remains to prove the theorem over $X^{\star}$. Without loss of generality we can suppose that $X^{\star}=X$. Consider the family of spectral projections
-(5.1) $p_{\lambda}(x)=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})\;\;\text{for}\;\;x\in X_{\lambda}=\left\{x\in X\,|\,\lambda\in\operatorname{Res}(\mathcal{A}_{x})\right\}.$
-The sets $X_{\lambda}$ with $\lambda$ running the interval $(-1,1)$ form an open covering of $X$.
+(7.1) $p_{\lambda}(x)=\operatorname{\mathds{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})\;\;\text{for}\;\;x\in X_{\lambda}=\left\{x\in X\mid\lambda\in\operatorname{Res}(\mathcal{A}_{x})\right\}.$
+The sets $X_{\lambda}$ with $\lambda$ running the interval (1,1)(-1,1) form an open covering of $X$.
 A local trivialization of $\mathcal{H}$ adapted to $\mathcal{A}$ takes $p_{\lambda}$ to a norm continuous family over $X_{\lambda}$.
 For $\lambda<\mu$ the difference
-(5.2) $p_{\lambda}(x)-p_{\mu}(x)=\operatorname{\mathbb{1}}_{[\lambda,\mu)}(\mathcal{A}_{x})=\operatorname{\mathbb{1}}_{[\lambda,\mu]}(\mathcal{A}_{x})$
+(7.2) $p_{\lambda}(x)-p_{\mu}(x)=\operatorname{\mathds{1}}_{[\lambda,\mu)}(\mathcal{A}_{x})=\operatorname{\mathds{1}}_{[\lambda,\mu]}(\mathcal{A}_{x})$
 is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$.
@@
 Hence a trivialization is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$.
-It remains to apply Theorem 4.1.
+It remains to apply Theorem 4.
 $\square$
-Remark.
-In fact, the main part of Theorem 5.1, about operators which are neither essentially positive nor essentially negative, is equivalent to Theorem 4.1.
-Indeed, let $(X_{i},p_{i})$ be an atlas of a polarization $\Pi$.
-We associate with it the family $\alpha$ of self-adjoint essentially unitary operators $\alpha(x)=\sum_{i}\rho_{i}(x)r_{i}(x)$, where $(\rho_{i})$ is a partition of unity subordinated to the covering $(X_{i})$ and $r_{i}=2p_{i}-1$.
-Then $\alpha-r_{i}=\sum_{j}\rho_{j}(r_{j}-r_{i})$ is a norm continuous family of compact operators over $X_{i}$.
-Therefore, a (local or global) trivialization of $\mathcal{H}$ is adapted to $\Pi$ if and only if it is adapted to $\alpha$.
-Proof of Theorem D.
-The bounded transform takes self-adjoint operators with compact resolvents to self-adjoint essentially unitary operators and Riesz (quasi-)continuous families to norm (quasi-)continuous families.
-Therefore, Theorem D follows immediately from Theorem 5.1.
 Spectral projections.
-For a family $\mathcal{A}_{x}$, $x\in X$ of self-adjoint operators and $\lambda\in\mathbb{R}$ we consider the family $p_{\lambda}$ of spectral projections over $X_{\lambda}\subset X$ defined by (5.1).
+For a family $\mathcal{A}_{x}$, $x\in X$ of self-adjoint operators and $\lambda\in\mathbb{R}$ we consider the family $p_{\lambda}$ of spectral projections over the subset $X_{\lambda}\subset X$ defined by (7.1).
 Proposition 7.1.
-If $\mathcal{A}\colon X\to\mathcal{R}^{\mathrm{sa}}(H)$ is graph continuous, then $p_{\lambda}\colon X_{\lambda}\to\mathcal{P}(H)$ is strong continuous for every $\lambda\in\mathbb{R}$.
+If $\mathcal{A}\colon X\to\mathcal{R}^{\mathrm{sa}}(H)$ is graph continuous, then $p_{\lambda}\colon X_{\lambda}\to\mathcal{P}(H)$ is strongly continuous for every $\lambda\in\mathbb{R}$.
 Proof. Without loss of generality we can suppose that $X={{}^{g}\mathcal{R}}^{\mathrm{sa}}(H)$ and $\mathcal{A}$ is the identity map. In other words, we need to prove that the spectral projection
-$p_{\lambda}=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}\colon\mathcal{R}^{\mathrm{sa}}_{\lambda}\to\mathcal{P}(H),\quad\mathcal{R}^{\mathrm{sa}}_{\lambda}=\left\{A\in\mathcal{R}^{\mathrm{sa}}(H)\,|\,\lambda\in\operatorname{Res}(A)\right\}$
-[…]
+$p_{\lambda}=\operatorname{\mathds{1}}_{[\lambda,+\infty)}\colon\mathcal{R}^{\mathrm{sa}}_{\lambda}\to\mathcal{P}(H),\quad\mathcal{R}^{\mathrm{sa}}_{\lambda}=\left\{A\in\mathcal{R}^{\mathrm{sa}}(H)\mid\lambda\in\operatorname{Res}(A)\right\}$
+is graph-to-strong continuous.
+Let $A_{0}\in\mathcal{R}^{\mathrm{sa}}_{\lambda}$.
+Then $\operatorname{Res}(A_{0})$ contains the interval $\lambda-\varepsilon,\lambda+\varepsilon]$ for some $\varepsilon>0$.
+By [Kat, Theorem IV.3.1], there is a neighborhood $V$ of $A_{0}$ in ${{}^{g}\mathcal{R}}^{\mathrm{sa}}$ consisting of operators whose resolvent set contains this interval.
+Let $f\colon\mathbb{R}\to[0,1]$ be a continuous function equal to 0 on $(-\infty,\lambda-\varepsilon]$ and to 1 on $\lambda+\varepsilon,\infty)$.
+The function $g=\bm{\chi}-(2f-1)$ is continuous on $\mathbb{R}$ and vanishes at $\pm\infty$, so it can be approximated by polynomials in $(x+i)^{-1}$ and $(x-i)^{-1}$ and thus induces a graph-to-norm continuous map $g\colon\mathcal{R}^{\mathrm{sa}}\to\mathcal{B}$.
+See [RS], Theorem VIII.20(a).
+By Proposition 2.1 $\bm{\chi}$ is graph-to-strong continuous, so $f=(\bm{\chi}-g+1)/2$ is also graph-to-strong continuous.
+Since $p_{\lambda}$ coincides with $f$ on $V$ and $A_{0}$ was chosen arbitrary, $p_{\lambda}$ is graph-to-strong continuous on the whole $\mathcal{R}^{\mathrm{sa}}_{\lambda}$.
+$\square$
 Proposition 7.2.
@@
 Then the family $a_{x}=\bm{\chi}(\mathcal{A}_{x})$ is norm continuous and $\mu=\bm{\chi}(\lambda)$ lies in the resolvent set of each $a_{x}$.
-Therefore, the map $x\mapsto p_{\lambda}(x)=\operatorname{\mathbb{1}}_{[\mu,+\infty)}(a_{x})$ is also norm continuous.
+Therefore, the map $x\mapsto p_{\lambda}(x)=\operatorname{\mathds{1}}_{[\mu,+\infty)}(a_{x})$ is also norm continuous.
 Conversely, suppose that $p_{\lambda}$ is norm continuous.
-Then $r_{\lambda}=2p_{\lambda}-1$ is a norm continuous family of symmetries commuting with $\mathcal{A}$.
-The family $\mathcal{A}^{\prime}=\mathcal{A}\cdot r_{\lambda}$ of self-adjoint operators is graph (resp Riesz) continuous if and only if $\mathcal{A}$ is graph (resp Riesz) continuous.
-But, in contrast with $\mathcal{A}$, the family $\mathcal{A}^{\prime}$ is uniformly bounded from below by $-|\lambda|$.
-The graph and Riesz topology coincide on the set of operators bounded from below by a given constant [$P_{2}$, Proposition 11.3], so $\mathcal{A}^{\prime}$ and thus also $\mathcal{A}$ are Riesz continuous.
+Then $r_{\lambda}=2p_{\lambda}-1$ is a norm continuous family of symmetries (that is, self-adjoint unitary operators) commuting with $\mathcal{A}$.
+By Proposition 2.2, the family $\mathcal{A}^{\prime}=\mathcal{A}\cdot r_{\lambda}$ of self-adjoint operators is graph (resp Riesz) continuous if and only if $\mathcal{A}$ is graph (resp Riesz) continuous.
+But, in contrast with $\mathcal{A}$, the family $\mathcal{A}^{\prime}$ is uniformly bounded from below by $|\lambda|$.
+The graph and Riesz topology coincide on the set of operators bounded from below by a given constant [P2, Proposition 11.2], so $\mathcal{A}^{\prime}$ and thus also $\mathcal{A}$ are Riesz continuous.
 $\square$
 Polarization associated with operator family.
-Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators with compact resolvents acting on fibers of a Hilbert bundle $\mathcal{H}$.
-Then $X_{\lambda}$ is open for every $\lambda\in\mathbb{R}$ by [Ka, Theorem IV.3.1] and these sets form an open covering of $X$.
-By Proposition 5.2, $p_{\lambda}$ is strongly continuous in local trivializations of $\mathcal{H}$, so it determines a decomposition of $\mathcal{H}$ over $X_{\lambda}$ into the orthogonal sum $\mathcal{H}_{\lambda}^{+}\oplus\mathcal{H}_{\lambda}^{-}$ of two Hilbert subbundles.
-For $\lambda<\mu$ the difference (5.2) is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$ by [Ka, Theorem IV.3.16].
+Let $\mathcal{A}$ be a graph continuous family of self-adjoint operators $\mathcal{A}_{x}\colon\mathcal{H}_{x}\to\mathcal{H}_{x}$ with compact resolvents.
+Then $X_{\lambda}$ is open for every $\lambda\in\mathbb{R}$ by [Kat, Theorem IV.3.1] and these sets form an open covering of $X$.
+By Proposition 7.1, $p_{\lambda}$ is a bundle projection of $\mathcal{H}$ over $X_{\lambda}$.
+For $\lambda<\mu$ the difference (7.2) is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$ by [Kat, Theorem IV.3.16].
 Suppose additionally that each $\mathcal{A}_{x}$ is neither essentially positive nor essentially negative.
-Then $p_{\lambda}(x)\in\mathcal{P}^{\star}(\mathcal{H}_{x})$, so the collection $(X_{\lambda},p_{\lambda})$ is an atlas of a polarization.
+Then the collection $(X_{\lambda},p_{\lambda})$ is an atlas of a (finite) polarization.
 Hence a family $\mathcal{A}$ satisfying assumptions above determines a polarization of $\mathcal{H}$ which we call the polarization associated with $\mathcal{A}$ and denote by $\Pi_{\mathcal{A}}$.
-(The associated polarization $\Pi_{\mathcal{A}}$ is actually finite, but we are interested only in its equivalence class with respect to compact deformations, so we will consider it as a compact polarization.)
 Proposition 7.3.
@@
 Proof. This follows immediately from Proposition 7.2. $\square$
-Remark.
-More generally, consider a family $\mathcal{A}$ of Fredholm self-adjoint operators acting on fibers of $\mathcal{H}$.
-Suppose that $\mathcal{A}$ is graph quasi-continuous (that is, $\mathcal{A}$ admits local trivializations taking it to a graph continuous family).
-Then the associated polarization $\Pi_{\mathcal{A}}$ is defined in a similar manner.
-The only change is in the definition of the set $X_{\lambda}$: one should add the condition that the interval $[\lambda,0]$ or $[0,\lambda]$, depending on the sign of $\lambda$, is disjoint from the essential spectrum of $\mathcal{A}_{x}$.
-If a trivialization is adapted to such a family $\mathcal{A}$, then it is adapted to $\Pi_{\mathcal{A}}$ as well.
-However, the second direction of Proposition 5.4 is no longer true: a trivialization adapted to $\Pi_{\mathcal{A}}$ is not necessarily adapted to $\mathcal{A}$, because it might turn $\mathcal{A}$ into a graph discontinuous family and then Proposition 5.3 cannot be applied.
-This is the place where condition of compact resolvents comes into the picture: the graph continuity of a family of operators with compact resolvents is independent of the choice of a trivialization.
-Proof of Theorem E.
+Proof of Theorem 7.
 Let $\Pi_{\mathcal{A}}$ be the polarization of $\mathcal{H}$ associated with $\mathcal{A}$.
 By Proposition 7.3, a (local or global) trivialization of $\mathcal{H}$ is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$.
-It remains to apply Theorem 4.4.
+It remains to apply Theorem 5.
 $\square$
+Remark 7.4.
+More generally, one can consider a family $\mathcal{A}$ of Fredholm self-adjoint operators acting on fibers of $\mathcal{H}$.
+Suppose that $\mathcal{A}$ admits local trivializations taking it to a graph continuous family.
+Then the associated polarization $\Pi_{\mathcal{A}}$ is defined in a similar manner.
+The only change is in the definition of the subset $X_{\lambda}$: one should add the condition that the interval $\lambda,0]$ or $0,\lambda]$, depending on the sign of $\lambda$, is disjoint from the essential spectrum of $\mathcal{A}_{x}$.
+If a trivialization is adapted to such a family $\mathcal{A}$, then it is adapted to $\Pi_{\mathcal{A}}$ as well.
+However, the converse statement is no longer true: a trivialization adapted to $\Pi_{\mathcal{A}}$ is not necessarily adapted to $\mathcal{A}$, because it might turn $\mathcal{A}$ into a graph discontinuous family and then Proposition 7.2 cannot be applied.
+This problem arises already for continuous maps $X\to{{}^{g}\mathcal{R}}^{\star}_{F}(H)$.
+This is the place where condition of having compact resolvents comes into the picture: the graph continuity of a family of operators with compact resolvents is independent of the choice of a trivialization.
 Example 7.5.
-Let us show that Theorem E does not hold without additional assumptions on the base space $X$.
-We will construct a graph continuous family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ of invertible self-adjoint operators with compact resolvents over a compact metrizable space $X$ such that $\mathcal{A}$ is not Riesz quasi-continuous.
-Moreover, for every point $x\in X$, $\mathcal{A}$ is not Riesz quasi-continuous at $x$ (that is, it is not Riesz quasi-continuous in any neighborhood of $x$).
-Our example is based on a construction of a non-trivial separable Hilbert bundle $\mathcal{E}$ over a compact space $X$ by Dixmier and Douady.
+Let us show that Theorem 7 does not hold without additional assumptions on the base space $X$.
+We will construct a graph continuous family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ of invertible self-adjoint operators with compact resolvents over a compact metrizable space $X$ such that, for every point $x\in X$, $\mathcal{A}$ has no adapted trivialization in any neighborhood of $x$.
+Our example is based on a Dixmier–Douady’s construction of a non-trivial separable continuous field $\mathcal{E}$ of Hilbert spaces over a compact space $X$.
 See [DD, Theorem 6].
@@
 The space $X$ is a countable product of metrizable spaces and thus is metrizable itself; it is a product of compact spaces and thus is compact itself.
-The Hilbert bundle $\mathcal{E}$ over $X$ is defined as the subbundle of the trivial Hilbert bundle $H_{X}$ with the fiber $\mathcal{E}_{x}=\oplus_{i=1}^{\infty}x_{i}$ over $x=(x_{i})$, where $x_{i}\in X_{i}$ is considered as a one-dimensional subspace of $H_{i}$.
+For $x=(x_{i})\in X$, let $\mathcal{E}_{x}\subset H$ be the closed span of $\oplus_{i=1}^{\infty}x_{i}$, where $x_{i}\in X_{i}$ is considered as a one-dimensional subspace of $H_{i}\subset H$.
+Let $p_{x}\in\mathcal{P}(H)$ be the projection with the range $\mathcal{E}_{x}$.
+The corresponding map $p\colon X\to\mathcal{P}(H)$ is strongly (but not uniformly) continuous, so the subspaces $\mathcal{E}_{x}\subset H$ give rise to the subbundle (in our sense) $\mathcal{E}$ of $H_{X}$.
 By [DD, Theorem 6], every section of $\mathcal{E}$ vanishes at some point.
-In particular, $\mathcal{E}$ is not trivial and thus it is not locally trivial (since all locally trivial Hilbert bundles over paracompact base spaces are trivial).
-Moreover, the bundle $\mathcal{E}$ looks the same near each point of $X$ (more precisely, the group $\prod_{i=1}^{\infty}\mathcal{U}(H_{i})$ acts transitively on $X$ and this action lifts to the action on $\mathcal{E}$ by bundle automorphisms), so a local triviality of $\mathcal{E}$ near one point would imply local triviality over the whole $X$.
-Therefore, for every open subset $X^{\prime}\subset X$, $\mathcal{E}$ is not trivial over $X^{\prime}$.
-Let $c\colon\mathbb{N}\to\mathbb{R}_{>0}$ be a sequence of positive numbers converging to $\infty$. Then
+Therefore, $\mathcal{E}$ is not a locally trivial Hilbert bundle (since every locally trivial Hilbert bundle over a paracompact base space is trivial and thus have non-vanishing sections).
+Moreover, $\mathcal{E}$ looks the same near each point of $X$ (more precisely, the group $\prod_{i=1}^{\infty}\mathcal{U}(H_{i})$ acts transitively on $X$ and this action lifts to the action on $\mathcal{H}$ by bundle automorphisms preserving $\mathcal{E}$), so a local triviality of $\mathcal{E}$ near one point would imply local triviality over the whole $X$.
+Therefore, for every open subset $X^{\prime}\subset X$, $\mathcal{E}$ is not trivial over $X^{\prime}$ and thus has no adapted trivializations over $X^{\prime}$.
+Let $(c_{i})_{i\in\mathbb{N}}$ be a sequence of positive numbers converging to $\infty$. Then
 $A=\oplus_{i=1}^{\infty}c_{i}\cdot\operatorname{Id}_{H_{i}}$
-is a positive regular operator with compact resolvents.
-Let $p(x)=\oplus_{i=1}^{\infty}p_{i}(x)$ be the orthogonal projection of $H$ onto $\mathcal{E}_{x}$, and let $r(x)=2p(x)-1$ be the corresponding symmetry.
-Then $r(x)$ commutes with $A$, and $\mathcal{A}(x)=A\cdot r(x)$ is an invertible self-adjoint operator with compact resolvents, which is neither essentially positive nor essentially negative.
-We claim that the family $\mathcal{A}\colon X\to\mathcal{R}^{\star}_{K}(H)$ is graph continuous. Indeed, $(1+\mathcal{A}_{x}^{2})^{-1}=(1+A^{2})^{-1}$ is a compact operator independent of $x$. The map
-$x\longmapsto\mathcal{A}_{x}(1+\mathcal{A}_{x}^{2})^{-1}=r_{x}A(1+A^{2})^{-1}$
-is the product of a strongly continuous family of symmetries $r_{x}$ and a fixed compact operator $A(1+A^{2})^{-1}=\bm{\chi}(A)(1+A^{2})^{-1/2}$ and thus is norm continuous.
-Therefore, all components of the orthogonal projection (2.6) onto the graph of $\mathcal{A}_{x}$ are norm continuous and thus $\mathcal{A}$ is graph continuous.
-However, for every open subset $X^{\prime}\subset X$, the restriction of $\mathcal{A}$ to $X^{\prime}$ is not Riesz quasi-continuous.
-Indeed, the positive spectral projection $p(x)$ of $\mathcal{A}(x)$ is the orthogonal projection of $H$ onto $\mathcal{E}_{x}$.
-If $\mathcal{A}$ were Riesz continuous in some local trivialization of the trivial Hilbert bundle $X^{\prime}\times H$, then the map $x\mapsto p(x)$ would be norm continuous in the same local trivialization and $\mathcal{E}$ would be a locally trivial Hilbert bundle over $X^{\prime}$, which is not the case.
+is a positive regular operator on $H$ with compact resolvents.
+The symmetry $r_{x}=2p_{x}-1$ commutes with $A$ and $\mathcal{A}_{x}=A\cdot r_{x}$ is an invertible self-adjoint operator with compact resolvents, which is neither essentially positive nor essentially negative.
+Since $r_{x}$ is a strongly continuous family of unitary operators, the family $\mathcal{A}_{x}$ is graph continuous by Proposition 2.2.
+The positive spectral projection of $\mathcal{A}_{x}$ is $p_{x}$.
+By Proposition 7.2, each local trivialization adapted to $p$ is also adapted to $\mathcal{A}$.
+Since $p$ has no adapted trivializations, the family $\mathcal{A}$ has no adapted trivializations as well.
+Moreover, for every open subset $X^{\prime}\subset X$, the restriction of $\mathcal{A}$ to $X^{\prime}$ has no adapted trivialization.
 Remark.
-If one multiplies the family $\mathcal{A}_{x}$ from this example by the strongly continuous family of unitary operators $r_{x}$, then the resulting family $x\mapsto\mathcal{A}_{x}\cdot r_{x}=A$ will be Riesz continuous (and even constant).
+If one multiplies the family $\mathcal{A}_{x}$ from this example by the strongly continuous family of unitary operators $r_{x}$, then the resulting family $x\mapsto\mathcal{A}_{x}r_{x}=r_{x}\mathcal{A}_{x}=A$ will be Riesz continuous (and even constant).
 However, for every strongly continuous family of unitary operators $v_{x}\in\mathcal{U}(H)$, the family $x\mapsto v_{x}\mathcal{A}_{x}v_{x}^{*}$ is Riesz discontinuous at every point $x\in X$.
-This demonstrates the difference between Theorems C and D, that is, between one-sided multiplication and conjugation by unitary operators, which we discussed in the Introduction.
+This demonstrates the difference between the self-adjoint and non-self-adjoint cases, that is, between one-sided multiplication and conjugation by unitary operators, which we discussed in the Introduction.